### v2

# On the numerically and cohomologically trivial automorphisms of elliptic surfaces II: $\chi(S)>0$

Abstract.
In this second part we study first the group $\mathrm{Aut}_{\mathbb{Q}}(S)$ of numerically trivial automorphisms of an algebraic properly elliptic surface $S$, that is, of a minimal algebraic surface with Kodaira dimension $\kappa(S)=1$, in the case $\chi(S)\geq 1$. Our first surprising result is that, against what has been believed for over 40 years, there exist nontrivial such groups for $p_{g}(S)>0$. Indeed, we show even that there is no absolute upper bound for their cardinalities $|\mathrm{Aut}_{\mathbb{Q}}(S)|$. At any rate, we give explicit and nearly optimal upper bounds for $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ in terms of the numerical invariants of $S$, as $\chi(S)$, or the irregularity $q(S)$, or the bigenus $P_{2}(S)$. Moreover, we come quite close to a complete description of the possible groups $\mathrm{Aut}_{\mathbb{Q}}(S)$ as 2-generated finite abelian groups and we give effective criteria for surfaces to have trivial $\mathrm{Aut}_{\mathbb{Q}}(S)$.
Our second surprising results concern the group $\mathrm{Aut}_{\mathbb{Z}}(S)$ of cohomologically trivial automorphisms; we are able to give the explicit upper bounds for $|\mathrm{Aut}_{\mathbb{Z}}(S)|$ in special cases: $9$ when $p_{g}(S)=0$, and the sharp upper bound $3$ when $S$ (i.e., the pluricanonical elliptic fibration) is isotrivial. We produce also non isotrivial examples where $\mathrm{Aut}_{\mathbb{Z}}(S)$ is a cyclic group of order $2$ or $3$.
Key words and phrases:
Compact Kähler manifolds, algebraic surfaces, elliptic surfaces, automorphisms, cohomologically trivial automorphisms, numerically trivial automorphisms, Enriques–Kodaira classfication
2010 Mathematics Subject Classification:
14J50, 14J80, 14J27, 14H30, 14F99, 32L05, 32M99, 32Q15, 32Q55
1. Introduction
Let $X$ be a compact connected complex manifold. By [BM46, BM47] the automorphism group $\mathrm{Aut}(X)$ is a finite dimensional complex Lie Group, whose connected component of the identity is denoted $\mathrm{Aut}^{0}(X)$.
We continue our investigation of two special subgroups:
(i) $\mathrm{Aut}_{\mathbb{Q}}(X)$, called the group of numerically trivial automorphisms, is the subgroup of the automorphisms acting as the identity on $H^{*}(X,\mathbb{Q})$,
(ii) respectively $\mathrm{Aut}_{\mathbb{Z}}(X)$, the group of cohomologically trivial automorphisms, consists of those automorphisms acting as the identity on $H^{*}(X,\mathbb{Z})$.
Of course they coincide, $\mathrm{Aut}_{\mathbb{Z}}(X)=\mathrm{Aut}_{\mathbb{Q}}(X)$, if $H^{*}(X,\mathbb{Z})$ has no torsion.
Recall moreover, as already mentioned in Part I [CFGLS24], that in the case where $X$ is a compact Kähler manifold, $\mathrm{Aut}_{\mathbb{Q}}(X)/\mathrm{Aut}^{0}(X)$ is a finite group ([Lie78], Fujiki [Fuj78]).
Hence the major problems to be considered are: to determine when this quotient group or $\mathrm{Aut}_{\mathbb{Z}}(X)/\mathrm{Aut}^{0}(X)$ may be nontrivial, to give explicit upper bounds for their cardinality in terms of the invariants of $X$, and to try to establish sharpness of these upper bounds by exhibiting appropriate examples.
Since for curves (complex dimension $n=1$) $\mathrm{Aut}_{\mathbb{Q}}(X)=\mathrm{Aut}^{0}(X)$, the first important case to consider is the case $n=2$ (hence $X$, being a surface, shall be denoted by $S$), which we already treated in [CL21] and in Part I [CFGLS24].
This paper is dedicated to the remaining case of minimal surfaces of Kodaira dimension $1$ with $\chi(S)>0$: they are automatically Kähler, and for them $\mathrm{Aut}^{0}(S)$ is trivial; we shall deal here with the case of algebraic surfaces, leaving the case of non algebraic surfaces for a future paper.
The main purpose of this article is to revive the interest for this very subtle topic of research, which, like the Arab Phoenix, has a rebirth from its own ashes. Indeed, for over 40 years it was believed that these groups are trivial for $p_{g}(S)>0$ (Theorem 4.5 of [Pe80], confirmed by Cor. 0.3 of [Cai09]).
Recent results of Dolgachev–Martin [DM22] for the case where $S$ is a rational surface of any characteristic (which we shall use) go also in the direction of investigating numerically trivial automorphisms (these are automatically cohomologically trivial if the surfaces are rational).
Our results, which will be more amply illustrated later on in 1.3, are rather complete for the case of the group $\mathrm{Aut}_{\mathbb{Q}}(S)$ of numerically trivial automorphisms.
For the intriguing case of the group $\mathrm{Aut}_{\mathbb{Z}}(S)$ of cohomologically trivial automorphisms we establish again, with very hard work, the existence (as in part I) of nontrivial groups of order $2,3$: but we cannot establish an absolute upper bound for $|\mathrm{Aut}_{\mathbb{Z}}(S)|$ except under special assumptions, for instance if $p_{g}(S)=0$, or if $S$ is isotrivially fibred.
Hence the main open question remaining is whether there is an absolute upper bound for $|\mathrm{Aut}_{\mathbb{Z}}(S)|$ in the non isotrivial case with $p_{g}(S)>0$.
To explain the difficulty of the question, the problem is that, even if one shows that an automorphism has trivial action on $H_{1}(S,\mathbb{Z})=H^{3}(S,\mathbb{Z})$, there remains to handle the case of $H^{2}(S,\mathbb{Z})$, which has a non canonical splitting as a free part plus a torsion subgroup, canonically isomorphic to the torsion of $H_{1}(S,\mathbb{Z})$. Hence the difficulty is to find a system of explicit generators for $H^{2}(S,\mathbb{Z})$ in terms of which we can then describe the action of automorphisms; this is quite hard in general, and almost impossible in the case where there is nontrivial torsion and $p_{g}(S)>0$.
1.1. Main ideas and tools: two allied points of view
Throughout the paper, $f\colon S\to B$ denotes the pluricanonical fibration of a properly elliptic surface with $\chi(S):=\chi(\mathcal{O}_{S})>0$.
As the reader shall see, there are two allied points of view to take.
The most algebraic point of view is to look at the function field extension $\mathbb{C}(S)\supset\mathbb{C}(B)$ as providing a curve ${\mathcal{C}}$ of genus $1$ (the fibre of $f$ over the generic point of $B$) defined over the non algebraically closed field ${\mathcal{K}}:=\mathbb{C}(B)$. This point of view is essential to have a simple algebraic picture of $\mathrm{Aut}(S)$: since in our case the group $\mathrm{Aut}(S)$ equals the group $\mathrm{Bir}(S)$ of birational self maps of $S$, and $\mathrm{Bir}(S)=Aut(\mathbb{C}(S))=Aut({\mathcal{C}})$.
On the other hand the study of the topological properties of automorphisms is based on the relative minimal model $f\colon S\rightarrow B$, that is, on the action on the fibration and on its singular fibres; here there is a crucial difference for checking when automorphisms are numerically trivial, respectively cohomologically trivial.
The algebraic approach tells us immediately that, if we look at the subgroup $\mathrm{Aut}_{{\mathcal{K}}}({\mathcal{C}})$ of the automorphisms which act as the identity on the base field ${\mathcal{K}}$, these embed in $\mathrm{Aut}_{\bar{{\mathcal{K}}}}({\mathcal{C}}\otimes\bar{{\mathcal{K}}})$, where ${\mathcal{E}}:={\mathcal{C}}\otimes\bar{{\mathcal{K}}}$ is an elliptic curve over an algebraically closed field of characteristic zero,*** Indeed, it suffices to take a finite field extension ${\mathcal{K}}^{\prime}\supset{\mathcal{K}}$ to get an elliptic curve. hence
(1.1) $\mathrm{Aut}_{\bar{{\mathcal{K}}}}({\mathcal{E}})={\mathcal{E}}\rtimes\mu_{r},\quad r\in\{2,4,6\},$
($r=4$ for the Gaussian = harmonic elliptic curve, $r=6$ for the Fermat = equianharmonic elliptic curve). Here $\mu_{r}$ is the group of $r$th roots of $1$ in $\mathbb{C}$, and ${\mathcal{E}}$ acts by translations via the isomorphism ${\mathcal{E}}\cong\operatorname{Pic}^{1}({\mathcal{E}})\cong\operatorname{Pic}^{0}({\mathcal{E}})$ provided by the choice of a point $0\in{\mathcal{E}}$ ($\operatorname{Pic}^{d}({\mathcal{E}})$ is as usual the set of linear equivalence classes of divisors of degree $d$).
If we go back to ${\mathcal{C}}$, there is no longer (as we were taught by Kodaira [Kod63] and Shafarevich [Shaf65, Chapter VII]) an isomorphism ${\mathcal{C}}\cong\operatorname{Pic}^{0}({\mathcal{C}})$ as curves over the field ${\mathcal{K}}$ if there is no ${\mathcal{K}}$-rational point on ${\mathcal{C}}$, but the isomorphism ${\mathcal{C}}\cong\operatorname{Pic}^{1}({\mathcal{C}})$ holds true (by Riemann Roch and by descent).
Then we denote as usual $\mathrm{Jac}({\mathcal{C}}):=\operatorname{Pic}^{0}({\mathcal{C}})$, and ${\mathcal{C}}$ is a principal homogeneous space over $\mathrm{Jac}({\mathcal{C}})$ ([Shaf65], Chapter 2 of [Dolg81]), in particular, the Mordell–Weil group, the group of ${\mathcal{K}}$-rational points of $\mathrm{Jac}({\mathcal{C}})$ acts on ${\mathcal{C}}$.
The torsion subgroup of the Mordell–Weil group will be a major source for the construction of numerically trivial automorphisms (see especially Section 7).
The geometrical counterpart of $\mathrm{Jac}({\mathcal{C}})$ is the Jacobian surface $J(S)$, a suitable compactification of the sheaf of groups ${\mathcal{R}}^{1}f_{*}\mathcal{O}_{S}/{\mathcal{R}}^{1}f_{*}\mathbb{Z}$ (whose stalks have a connected component of the identity which is either an elliptic curve, or $\mathbb{C}^{*}$, or $\mathbb{C}$: in the last case one says that the fibre is of additive type).
$J(S)$ has the property that the singular fibres of $J(S)$ are the same as the singular fibres $F$ of $S$ when $F$ is not multiple, while they are isogenous to $F_{\mathrm{red}}$ when $F$ is multiple. In this way, even if the global topological behaviour of $J(S)$ differs from that of $S$, at least the topological nature of the fibres is the same.
Now, the curve ${\mathcal{C}}$ admits also automorphisms which do not act as the identity on the field ${\mathcal{K}}$, and we have an exact sequence
$1\rightarrow\mathrm{Aut}_{{\mathcal{K}}}({\mathcal{C}})\rightarrow\mathrm{Aut}({\mathcal{C}})\rightarrow\mathrm{Aut}({\mathcal{K}})$
where the left hand side group by (1.1) contains the Mordell–Weil group as a subgroup of index $r\leq 6$.
The geometrical counterpart is the exact sequence
(1.2) $1\rightarrow\mathrm{Aut}_{B}(S)\rightarrow\mathrm{Aut}(S)\xrightarrow{r_{B}}\mathrm{Aut}(B)$
and we denote by $\mathrm{Aut}(S)|_{B}$ the image of the right hand homomorphism.
1.2. Main strategy: a dichotomy of cases
Based on (1.2), the strategy is of course to give upper bounds for $|\mathrm{Aut}(S)|$ via upper bounds for the cardinalities $|\mathrm{Aut}(S)|_{B}|$ of the image group (here an essential role is given by the analysis of the singular fibres) and $|\mathrm{Aut}_{B}(S)|$ of the group of automorphisms acting trivially on the base curve $B$. Obviously this approach can also be applied to compute $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ and $|\mathrm{Aut}_{\mathbb{Z}}(S)|$.
Here, the analysis bifurcates: if the fibration is isotrivial, that is, birational to a quotient $(C\times E)/G$ (where the action of $G$ is no longer assumed to be free), then we use many of the tools developed in Part I, but now enhanced (as we shall see) by the presence of reducible fibres, which have to be of additive type.
Whereas, if the fibration is not isotrivial (that is, the smooth fibres are not all isomorphic), then the Mordell–Weil group plays a dominant role: since then, in view of the finiteness of $\mathrm{Aut}_{\mathbb{Q}}(S)$, we have to study when elements of the torsion subgroup $\mathrm{MW}(J(S))_{\mathrm{tor}}$ yield numerically trivial automorphisms.
One of the main advantages is that there is a huge literature on these groups (see [SS19]), especially in the case where $J(S)$ is a rational surface (this is the case for $p_{g}(S)=0$) there are explicit lists by Miranda–Persson [MP86], [Mir89], Oguiso–Shioda [OS90] and others describing these groups.
And when $J(S)$ is not rational, the method of viewing it as a pull back from an elliptic modular surface paves the way to exhibiting large groups of numerically trivial automorphisms (in Section 7).
Some existence results and arguments also apply to Enriques surfaces, see Theorem 7.6 and Remark 9.7: the case of Enriques surfaces indeed suggests the choices of candidates for an integral basis of the second cohomology (as we shall exploit in Section 10) and motivates our basic construction of numerically trivial automorphisms, Construction 7.5, which is an important special case of constructing $S$ as a principal homogeneous space over a jacobian elliptic surface.
1.3. Main results for numerically trivial automorphisms
Let $f\colon S\to B$ denote a properly elliptic algebraic surface with $\chi(S)>0$. Our main theorems show that the group $\mathrm{Aut}_{\mathbb{Q}}(S)$ of numerically trivial automorphisms can be arbitrarily large, but in a controlled way. We do this by controlling the induced action of $\mathrm{Aut}_{\mathbb{Q}}(S)$ on the base $B$ of the fibration and then by relating $\mathrm{Aut}_{\mathbb{Q}}(S)$ to the torsion part $\mathrm{MW}(J(S))_{\mathrm{tor}}$ of the Mordell–Weil group of the relative Jacobian $J(S)$.
The behaviour of $\mathrm{Aut}_{\mathbb{Q}}(S)$ differs a lot according to whether $p_{g}(S)$ is positive or not. In the case where $p_{g}(S)>0$, or, equivalently, where $J(S)$ is not rational, $\mathrm{Aut}_{\mathbb{Q}}(S)$ preserves each fibre and induces translations on smooth fibres, thus comes from a subgroup of $\mathrm{MW}(J(S))_{\mathrm{tor}}$, as is illustrated in Lemma 4.4. The upshot is then that, via our main construction of examples (Construction 7.5), we give a flexible sufficient condition for a subgroup $G\subset\mathrm{MW}(J(S))_{\mathrm{tor}}$ to be realized as a subgroup of $\mathrm{Aut}_{\mathbb{Q}}(S)$.
Theorem 1.1 (Groups for positive geometric genus).
Suppose that $p_{g}(S)>0$. Then the following hold:
(i) $\mathrm{Aut}_{\mathbb{Q}}(S)$ is isomorphic to a subgroup of $\mathrm{MW}(J(S))_{\mathrm{tor}}$, and as such it is a finite 2-generated abelian group.
(ii) Conversely, for any finite 2-generated abelian group $G$, there is a properly elliptic surface $S$ such that $\mathrm{Aut}_{\mathbb{Q}}(S)\supseteq G$. Equality $\mathrm{Aut}_{\mathbb{Q}}(S)=G$ can be ensured if we have an isomorphism $G\cong\mathbb{Z}/N\mathbb{Z}\times\mathbb{Z}/N^{\prime}\mathbb{Z}$ such that, at every prime $p$, the $p$-adic valuations satisfy $|v_{p}(N)-v_{p}(N^{\prime})|\leq 1$.
(iii) $\mathrm{Aut}_{\mathbb{Q}}(S)$ is trivial if there is a fibre of additive type or if all multiple fibres of $S\to B$ have smooth support; in particular, this holds true if there are no multiple fibres, e.g. if the fibration admits a section, that is, $S\cong J(S)$.
(iv) $\mathrm{Aut}_{\mathbb{Q}}(S)$ is trivial if the fibration is isotrivial.
By bounding the Mordell–Weil group of the elliptic modular surfaces $X(N)$ and $X_{1}(N)$ associated to the congruence modular groups $\Gamma(N)$ and $\Gamma_{1}(N)$, we obtain an upper bound for $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ in the case $p_{g}(S)>0$. There are even some arithmetic relations between the prime exponents of $\mathrm{Aut}_{\mathbb{Q}}(S)$ and the Euler characteristic $\chi(S)$.
Theorem 1.2 (Bounds for $\mathrm{Aut}_{\mathbb{Q}}$ in positive geometric genus).
Suppose that $p_{g}(S)>0$. Then the following hold:
  1. (i)
    In terms of the irregularity $q(S)$, there is the global bound
    (1.3) $|\mathrm{Aut}_{\mathbb{Q}}(S)|\leq 12\pi^{2}(q(S)+2).$
  2. (ii)
    If $p\mid|\mathrm{Aut}_{\mathbb{Q}}(S)|$ for some prime $p>5$, then
    $p_{g}\geq\frac{p^{2}-1}{12}-\frac{p-1}{2}\;\;\;\text{ and }\;\;\;\frac{p^{2}-1}{24}\mid\chi(S).$
    For $p\in\{2,3,5\}$, there are properly elliptic surfaces $S$ with any given $\chi>1$ or $p_{g}>0$ such that $p\mid|\mathrm{Aut}_{\mathbb{Q}}(S)|$.
  3. (iii)
    If $(\mathbb{Z}/p\mathbb{Z})^{2}\subset\mathrm{Aut}_{\mathbb{Q}}(S)$ for some prime $p\geq 3$, then
    $p_{g}\geq\frac{1}{12}(p-3)(p^{2}-1)\;\;\;\text{ and }\;\;\;\frac{p(p^{2}-1)}{24}\mid\chi(S).$
Details about the bounds in Theorem 1.2 can be found in Subsections 8.5, 8.6. We emphasize that the bound in (1.3) also holds if $p_{g}(S)=0$, with exception of a few cases, see Remark 6.14. This can also be deduced from (and improved by) our main theorem in the case $p_{g}=0$, see also Theorem 3.16 for part (ii).
Theorem 1.3 (Bounds for $\mathrm{Aut}_{\mathbb{Q}}$ for geometric genus zero).
Suppose that $p_{g}(S)=0$.
  1. (i)
    If $f$ is not isotrivial, then $|\mathrm{Aut}_{\mathbb{Q}}(S)|\leq 9$ and equality can hold. Moreover, in the equality case, the Jacobian $J(S)$ must be the extremal rational elliptic surface $X_{3333}$ in the list of [MP86].
  2. (ii)
    If $f$ is isotrivial, then, letting $s$ be the number of multiple fibres,
    $|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|\leq s\leq P_{2}(S)+1\;\;\;\text{ and }\;\;\;|\mathrm{Aut}_{\mathbb{Q}}(S)|\leq 4\cdot|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|;$
    conversely, for any $s\in\mathbb{N}$, there is an elliptic surface $S$ with $P_{2}(S)=s-1$, $|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|=s$ and $|\mathrm{Aut}_{\mathbb{Q}}(S)|=4s$.
    Moreover, if $|\mathrm{Aut}_{\mathbb{Q}}(S)|=4s$, then $J(S)$ must be the extremal rational elliptic surface $X_{33}$ in the list of [MP86].
We should point out that, for ease of presentation, the above theorem is largely simplified, in the sense that for many isotrivial fibrations the bounds from (i) hold, cf. Proposition 6.1.
1.4. Main results for cohomologically trivial automorphisms
Cohomologically trivial automorphisms are much rarer than numerically trivial ones. In contrast to Theorem 1.3, we have a uniform bound for $\mathrm{Aut}_{\mathbb{Z}}(S)$ in the case $p_{g}=0.$ The actual existence of properly elliptic surfaces with nontrivial $\mathrm{Aut}_{\mathbb{Z}}(S)$ is also surprising in view of the previous statements in the literature (cf. [Pe80, Theorem 4.5]).
Theorem 1.4 (Bounds for $\mathrm{Aut}_{\mathbb{Z}}$ for genus zero).
Suppose that $p_{g}(S)=0$. Then the following hold for $\mathrm{Aut}_{\mathbb{Z}}(S)$:
  1. (i)
    $\mathrm{Aut}_{\mathbb{Z}}(S)$ preserves every fibre of $f$.
  2. (ii)
    There is a uniform bound
    $|\mathrm{Aut}_{\mathbb{Z}}(S)|\leq 9.$
  3. (iii)
    There are non-isotrivial properly elliptic surfaces with $|\mathrm{Aut}_{\mathbb{Z}}(S)|=2$ and $3$. More precisely, there is a 1-dimensional family with $|\mathrm{Aut}_{\mathbb{Z}}(S)|=3$, and there are families of arbitrary dimension with $|\mathrm{Aut}_{\mathbb{Z}}(S)|=2$.
This can be further improved in the isotrivial case, in fact covering uniformly all $p_{g}$ (and also the case $\chi(S)=0$ by inspection of Part I). A more precise version of the following result is given in Theorem 3.6.
Theorem 1.5 (The isotrivial case).
Assume that $f:S\to B$ is isotrivial. Then
$|\mathrm{Aut}_{\mathbb{Z}}(S)|\leq 3,\;\;\;{\rm and\ the\ bound\ is\ sharp}.$
There are arbitrary dimensional families of isotrivial properly elliptic surfaces $S$ with $p_{g}(S)=0$ and with $|\mathrm{Aut}_{\mathbb{Z}}(S)|=2$, respectively with $|\mathrm{Aut}_{\mathbb{Z}}(S)|=3$.
A posteriori, we can infer the following from our results in this paper combined with those from [CFGLS24]:
Theorem 1.6.
For any algebraic properly elliptic surface $S$, $\mathrm{Aut}_{\mathbb{Z}}(S)$ is abelian.
1.5. Organization of the paper
Section 2 reviews preliminaries about automorphism groups, especially numerically trivial automorphisms, and elliptic fibrations, such as singular fibres and Mordell–Weil groups.
Section 3 focusses on the isotrivial case where we can make detailed use of the arguments of [CFGLS24], but also take into account the reducible fibres (by methods which will also be used in later parts of the paper); this gives a complete proof of Theorem 1.5, and partially of Theorem 1.3.
From Section 4 on, we treat elliptic fibrations in full generality, but with a certain focus on jacobian fibrations, i.e., those admitting a section. In particular, we study the interplay of $\mathrm{Aut}(S)$ and $\mathrm{MW}(J(S))$ for an elliptic surface $S$ and its Jacobian surface $J(S)$.
Section 5 bounds the size of the torsion subgroup of $\mathrm{MW}(X)$ for a jacobian elliptic surface $X$ in terms of the genus of the base curve.
This is used in Section 6 to prove Theorem 1.2 (i) (and eventually (ii), (iii)) as well as most of Theorem 1.4.
Section 7 introduces our main technique to construct explicit properly elliptic surfaces with given groups of numerically trivial automorphisms.
In particular, this paves the way to the proof of Theorem 1.1 in Section 8, and of Theorem 1.3 in Section 9. We turn to cohomologically trivial automorphisms in Sections 10, 11 by explicitly constructing the surfaces required to complete the proof of Theorem 1.4.
Convention 1.7.
Throughout the paper, we work with algebraic surfaces over the complex numbers, although our constructions work over arbitrary fields as long as the characteristic does not divide the group order.
Root lattices are assumed to be negative definite, and $\sim$ is meant to indicate numerical equivalence of divisors.
2. Notation and preliminaries
2.1. Special subgroups of the automorphism groups
Let $X$ be a compact complex manifold, and $R$ a commutative ring, which may be $\mathbb{Z}$, or $\mathbb{Q}$.
We denote by $\mathrm{Aut}(X)$ the Lie group of biholomorphic automorphisms of $X$, and consider its action on the cohomology ring $H^{*}(X,R)$.
The main concern of this paper is the study of the kernel of such an action:
$\mathrm{Aut}_{R}(X):=\left\{\sigma\in\mathrm{Aut}(X)\mid\sigma^{*}\xi=\xi\text{ for any $\xi\in H^{*}(X,R)$}\right\}.$
The elements of $\mathrm{Aut}_{\mathbb{Z}}(X)$ is said to be cohomologically trivial, while those of $\mathrm{Aut}_{\mathbb{Q}}(X)$ are said to be numerically trivial. It is clear that,
$\mathrm{Aut}_{\mathbb{Z}}(X)\subset\mathrm{Aut}_{\mathbb{Q}}(X).$
And these coincide if $H^{*}(X,\mathbb{Z})$ has no torsion.
If $X$ is a compact Kähler manifold, then $\mathrm{Aut}_{\mathbb{Q}}(X)$ has finitely many components by [Lie78] or [Fuj78], hence $\mathrm{Aut}_{\mathbb{Q}}(X)$ is finite unless there are vector fields on $X$.
For an algebraic properly elliptic surface $S$, the existence of vector fields is equivalent to being pseudo-elliptic, which means that $S=(C\times E)/G$ is isogenous to a product, and with $G$ acting by translations on the elliptic curve $E$: in particular $\chi(S)=0$, and we have the first of the two cases studied in Part I, [CFGLS24].
In this paper, we will thus focus on the case $\chi(S)>0$.
Occasionally, we use the group of ${\mathcal{O}}$-cohomologically trivial automorphisms:
$\mathrm{Aut}_{\mathcal{O}}(X):=\left\{\sigma\in\mathrm{Aut}(X)\mid\sigma^{*}\xi=\xi\text{ for any $\xi\in H^{*}(X,{\mathcal{O}}_{X})$}\right\};$
key examples, also for our considerations, are provided by translations by sections on elliptic surfaces. Via the Hodge decomposition, one has the inclusion $\mathrm{Aut}_{\mathbb{Q}}(X)\subset\mathrm{Aut}_{\mathcal{O}}(X)$.
Suppose that there is a fibration $f\colon X\rightarrow B$, that is, a surjective morphism between complex spaces with connected fibres. Then we may consider the group of fibration preserving automorphisms
$\mathrm{Aut}_{f}(X):=\left\{\sigma\in\mathrm{Aut}(X)\mid\text{$\exists\,\sigma_{B}\in\mathrm{Aut}(B)$ such that $\sigma_{B}\circ f=f\circ\sigma$ }\right\}.$
There is a natural homomorphism $r_{B}\colon\mathrm{Aut}_{f}(X)\rightarrow\mathrm{Aut}(B)$ such that, for $\sigma\in\mathrm{Aut}_{f}(X)$, its image $\sigma_{B}:=r_{B}(\sigma)$ satisfies $\sigma_{B}\circ f=f\circ\sigma$. The kernel
$\mathrm{Aut}_{B}(X):=\ker(r_{B})$
is the group of fibre preserving automorphisms.
Now let $S$ be a smooth projective surface, and $f\colon S\rightarrow B$ a relative minimal elliptic fibration over a smooth projective curve. Suppose that the Kodaira dimension of $S$ is nonzero.
Then the fibration $f$ is induced by $|mK_{S}|$ for some nonzero integer $m$ and hence preserved by the whole $\mathrm{Aut}(S)$, that is, $\mathrm{Aut}(S)=\mathrm{Aut}_{f}(S)$. Therefore, we have an exact sequence
$1\rightarrow\mathrm{Aut}_{B}(S)\rightarrow\mathrm{Aut}(S)\xrightarrow{r_{B}}\mathrm{Aut}(B)$
For ease of notation, denote
$\mathrm{Aut}_{B,R}(S):=\mathrm{Aut}_{B}(S)\cap\mathrm{Aut}_{R}(S),\quad\mathrm{Aut}_{R}(S)|_{B}=\mathrm{im}(\mathrm{Aut}_{R}(S)\xrightarrow{r_{B}}\mathrm{Aut}(B)).$
Then we have an exact sequence
(2.1) $1\rightarrow\mathrm{Aut}_{B,R}(S)\rightarrow\mathrm{Aut}_{R}(S)\rightarrow\mathrm{Aut}_{R}(S)|_{B}\rightarrow 1$
In order to give an upper bound for $|\mathrm{Aut}_{R}(S)|$, it suffices to bound $|\mathrm{Aut}_{B,R}(S)|$ and $|\mathrm{Aut}_{R}(S)|_{B}|$ from above.
Similarly, we introduce the notation:
(2.2) $\mathrm{Aut}_{B,{\mathcal{O}}}(S):=\mathrm{Aut}_{B}(S)\cap\mathrm{Aut}_{\mathcal{O}}(S).$
2.2. Notation and basic formulae
Establishing some further notation, we let
(2.3) $f\colon S\rightarrow B$
be a relatively minimal elliptic surface $S$, and we assume that $\chi(S)>0$, so not all the fibres can be multiple of a smooth elliptic curve (this is the case when $\chi(S)=0$, as treated in Part I). In this case we have that the genus of $B$ equals the irregularity $q(S)=h^{1}(\mathcal{O}_{S})$.
We denote by $s$ the number of multiple fibres, and by $m_{1},\dots,m_{s}$ their multiplicities: for $i=1,\dots,s$, we let $F_{i}$ be the corresponding multiple fibre, which can be written as $F_{i}=m_{i}F^{\prime}_{i}$ where $F^{\prime}_{i}$ cannot be written as a multiple of an effective divisor.
We have the Kodaira canonical bundle formula
(2.4) $K_{S}=f^{*}(K_{B}+L)+\sum_{1\leq i\leq s}(m_{i}-1)F^{\prime}_{i}$
where the $m_{i}F^{\prime}_{i}$ are as above the multiple fibres of $f$, and $\deg L=\chi(S)>0$.
If $S$ has Kodaira dimension $1$, we have
(2.5) $\deg(K_{B}+L)+\sum_{1\leq i\leq r}\left(1-\frac{1}{m_{i}}\right)=2g(B)-2+\chi(S)+\sum_{1\leq i\leq r}\left(1-\frac{1}{m_{i}}\right)>0.$
This formula implies that, in the case where $B$ is rational ($q(S)=0$ in our case), and $\chi(S)=1$, then there are at least two multiple fibres.
We denote further by $s_{\mathrm{red}}$ the number of reducible fibres which are not multiple, and by $s_{\mathrm{irr}}$ the number of irreducible and not multiple singular fibres.
Hence the total number of singular fibres equals $s^{\prime}:=s+s_{\mathrm{red}}+s_{\mathrm{irr}}$.
By the Noether formula, and since $K^{2}_{S}=0$, we have
(2.6) $12\chi(S)=e(S)=\sum_{i}^{s^{\prime}}e(f^{*}b_{i})$
where $f^{*}b_{i}$ runs through all the (finitely many) singular fibres.
2.3. Properly elliptic surfaces with $\chi(S)>0$ are Kähler, but not necessarily algebraic
Here we borrow first from the survey paper [Cat85], especially the classification table on page 21, remarks 44, 45 on page 22, and the discussion on pages 25-28, giving a very succinct presentation of the results obtained by Kodaira in [Kod60], [Kod63], [Kod64].
If a compact complex surface $S$ has $Kod(S)=1$, then we are in one of the two classes $4_{1}$ or $6_{1}$.
The latter class $6_{1}$ consists of elliptic quasi-bundles, and for these $b_{1}(S)$ is odd, hence they are not Kähler, and obviously for them $\chi(S)=0$.
In the class $4_{1}$, $b_{1}(S)$ is even, the algebraic dimension $a(S)\in\{1,2\}$, $S$ deforms to an algebraic surface and is Kähler.
Hence
I) a surface with $Kod(S)=1$ and $\chi(S)>0$ is necessarily Kähler.
Moreover
II) If $p_{g}(S)=0$ and $b_{1}(S)$ is even, then $S$ is algebraic (since $b^{+}(S)=2p_{g}(S)+1$).
If $S$ is not algebraic, then it has algebraic dimension $a(S)=1$ and the pluricanonical elliptic fibration $f:S\rightarrow C$ does not admit any multisection.
These surfaces $S$ are obtained from a fibration without multiple fibres as quotient by an abelian group $G$ operating equivariantly on $f^{\prime}:S^{\prime}\rightarrow B$: meaning that $S=S^{\prime}/G$, $C=B/G$.
In particular we have that $\chi(S^{\prime})>0\Leftrightarrow\chi(S)>0$.
If $f:S\rightarrow C$ has no multiple fibres and all fibres are smooth, then $S$ is a deformation of a product $E\times C$, where $E$ is an elliptic curve.
Assume now that $p_{g}(S)>0$ and $a(S)=1$.
As proven by Kodaira in [Kod60], $S^{\prime}$ is a deformation of an elliptic surface $J\rightarrow B$ admitting a section, which is called the Jacobian elliptic surface of $S^{\prime}$.
For a fixed Jacobian elliptic surface $J$, the parameter space for the surfaces $S^{\prime}$ is the cohomology group $H^{1}(C,{\mathcal{B}})$ , ${\mathcal{B}}$ being the sheaf of local holomorphic sections of the base curve $B$ to the open set of $J$ where the fibre is smooth. And to $S^{\prime}$ corresponds the invariant $\eta\in H^{1}(C,{\mathcal{B}})$.
We borrow now from [FM94] Proposition 5.23:
III) $S^{\prime}$, hence $S$, is algebraic if and only if $\eta$ is a torsion element.
IV) If $S^{\prime}$, equivalently $J$, has some singular fibres, then
$H^{1}(B,{\mathcal{B}})=H^{2}(\mathcal{O}_{J})/H^{2}(J,{\mathbb{Z}}).$
The above criterion shows that, if $p_{g}(S)>0$, and there are no multiple fibres, then we get a dense set of algebraic surfaces which deform to $S$. And also that there exist surfaces which are not algebraic and which have $Kod(S)=1$ and $\chi(S)>0$.
We shall consider in the sequel to this article the non algebraic case, the main difficulty being that the Jacobian of an elliptic surface with multiple fibres has been defined only for algebraic elliptic surfaces.
2.4. Invariance of the fibres by a numerically trivial automorphism
Let $\Psi\in Aut_{\mathbb{Q}}(S)$ be a numerically trivial automorphism. Then (see Part I, [CFGLS24], Section 2)
  1. (1)
    If $C$ is an irreducible curve with $C^{2}<0$, then $\Psi(C)=C$,
  2. (2)
    if there is a fibration $f:S\rightarrow D$ over a curve, then $\Psi$ preserves it, inducing an action on $D$;
  3. (3)
    any fibre $F$ such that $F_{red}$ is reducible is left invariant ($\Psi(F)=F$);
  4. (4)
    if a fibre is multiple of a smooth elliptic curve $F=mF^{\prime}$, and $\Psi\in Aut_{\mathbb{Z}}(S)$, then $\Psi(F)=F$ if $q(S)={\rm genus}(D)$: this holds unless all the fibres are smooth elliptic curves;
  5. (5)
    $\Psi$ acts as the identity on $D$ if $D$ has genus $\geq 2$, or if $D$ has genus $1$ and there is an invariant fibre, or if $D$ has genus 00 and there are at least three invariant fibres.
  6. (6)
    If $D$ has genus 00 and there are two invariant fibres, then $\Psi$ belongs to a cyclic group of automorphisms of $\mathbb{P}^{1}$, fixing two points.
A first consequence is:
Lemma 2.1.
Let $f:S\rightarrow B$ be a relatively minimal elliptic fibration, admitting a singular fibre which is not the multiple of a smooth elliptic curve (that is, $e(S),\chi(S)>0$): then
$H_{1}(S,\mathbb{Z})=H_{1}(B,\mathbb{Z})\bigoplus(\oplus_{1}^{s}(\mathbb{Z}/m_{i})\gamma_{i})/\mathbb{Z}(\sum_{1}^{s}\gamma_{j}),$
and $\Psi\in Aut_{\mathbb{Z}}(S)$ acts trivially on $H_{1}(S,\mathbb{Z})$.
Proof.
This is an immediate consequence of Lemma 2.5 of Part I and of (4), (5) above.
The next table collects the possible types for $F^{\prime}$ in Kodaira’s notation, together with the corresponding Dynkin types obtained by omitting a single simple fibre component:
Kodaira $\begin{array}[]{c||c|c|c|c|c|c|c|c}\text{Kodaira type}&\mathrm{I}_{n}\,(n>0)&\mathrm{I}_{n}^{*}\,(n\geq 0)&\mathrm{II}&\mathrm{III}&\mathrm{IV}&\mathrm{IV}^{*}&\mathrm{III}^{*}&\mathrm{II}^{*}\\ \hline\cr\text{Dynkin type}&A_{n-1}&D_{n+4}&A_{0}&A_{1}&A_{2}&E_{6}&E_{7}&E_{8}\end{array}$
Recall that the multiple fibres can only be of type ${}_{m}I_{k}$, with $m\geq 2$, and $k\geq 0$; while the fibres of type $I_{k}$ are called semistable (or multiplicative), and all the others are said to be of additive type; this means that the connected component of the identity of the algebraic group of which $F^{\#}$ (the smooth locus of the fibre $F$) is a principal homogeneous space equals $\mathbb{C}^{*}$.
The fibration is said to be semistable if all the fibres are semistable.
The only fibres above with $F_{red}$ irreducible are for $F=\phantom{}_{m}\mathrm{I}_{0},m\geq 2$, $F=\phantom{}_{m}\mathrm{I}_{1}$ or $F=\mathrm{II}$. Summing up,
Corollary 2.2.
The action of $\Psi\in Aut_{\mathbb{Q}}(S)$ on the base $B$ is trivial if $B$ has genus $\geq 2$, and also if $B$ has genus $1$, unless all fibres are of type ${}_{m}\mathrm{I}_{0}$, ${}_{m}\mathrm{I}_{1}$ or $\mathrm{II}$.
If $\Psi\in Aut_{\mathbb{Z}}(S)$, and it does not act trivially on $B$, then either $B$ has genus $1$ and all fibres are of type $\mathrm{I}_{1}$ or $\mathrm{II}$, or $B$ has genus 00 and there are at most two fibres not of type $\mathrm{I}_{1}$ or $\mathrm{II}$ (nodal, resp. cuspidal cubic).
2.5. Singular fibres and Mordell–Weil group
If $\chi(S)>0$, as assumed throughout this paper, then necessarily there is a singular fibre $F=mF^{\prime}$ with non-smooth support by inspection of the Euler number formula (2.6).
Note that the rank of the Dynkin diagram (assumed to be negative-definite by convention) equals one less than the number of fibre components (which we denote by $m_{v}$ if $F=f^{*}v$, slightly abusing notation).
If the fibration (2.3) admits a section, we call the elliptic surface $S$ jacobian (note that this excludes multiple fibres). Then the sections correspond bijectively to the rational points on the generic fibre $\mathcal{C}$; after choosing a zero section denoted by $O$ or $O_{S}$, the sections thus form an abelian group, called the Mordell–Weil group and denoted by $\mathrm{MW}(S)$, which is finitely generated if $\chi(S)>0$. We refer to [SS19] for more information on the theory of Mordell–Weil groups.
The Picard number of $S$ can be expressed by the Shioda–Tate formula:
(2.7) $\rho(S)=2+\mathrm{rank\,}\mathrm{MW}(S)+\sum_{v\in B}(m_{v}-1).$
Here the last summands correspond exactly to the Dynkin diagrams of the reducible fibres; usually one omits the fibre component intersecting $O$, often called identity component and denoted by $\Theta_{0}$. This paves the way towards the representation of the trivial lattice $\mathrm{Triv}(S)$ generated by fibre components and sections as an orthogonal sum
(2.8) $\mathrm{Triv}(S)=\langle O,F\rangle\oplus\bigoplus_{v\in B}(\text{Dynkin type associated to }F_{v}).$
Note that the leftmost summand is unimodular of parity determined by $O^{2}=-\chi(S)$.
The choice of zero section $O$ fixes a group structure on the smooth fibres which carries over to the smooth locus $F^{\#}$ of any singular fibre. We note the non-canonical isomorphism
(2.9) $F^{\#}\cong\begin{cases}\mathbb{G}_{m}\times A_{F},&\text{ if $F$ is multiplicative (type $\mathrm{I}_{n}$),}\\ \mathbb{G}_{a}\times A_{F},&\text{ if $F$ is additive (all other types)}\end{cases}$
Here $A_{F}$ denotes the discriminant group of the Dynkin type $D$ corresponding to the fibre, i.e. $A_{F}=D^{\vee}/D$. Note that $A_{F}$ is an abelian group which corresponds bijectively to the simple fibre components; in fact, $A_{F}$ is cyclic unless $F$ has type $\mathrm{I}_{n}^{*}$ for $n\in 2\mathbb{N}_{0}$ (whence $A_{F}\cong(\mathbb{Z}/2\mathbb{Z})^{2}$). In addition, $A_{F}$ is also endowed with the structure of a quadratic form, called discriminant form, taking values in $\mathbb{Q}/2\mathbb{Z}$.
Throughout this paper, it will be instrumental to consider torsion sections, and the automorphisms of $S$ given by translations (see (4.1)). Note that torsion sections are always disjoint from the zero section (generally true unless the characteristic of the ground field divides the order of the section). In terms of lattices, torsion sections can be encoded in the intersection pattern with the singular fibres as
(2.10) $\mathrm{MW}(S)_{\mathrm{tor}}\cong\mathrm{Triv}(S)^{\prime}/\mathrm{Triv}(S)\subset\bigoplus_{v\in B}A_{F_{v}}$
where $\mathrm{Triv}(S)^{\prime}=(\mathrm{Triv}(S)\otimes\mathbb{Q})\cap\mathrm{Num}(S)$ denotes the primitive closure of $\mathrm{Triv}(S)$ inside $\mathrm{Num}(S)$. In practice, this sees us searching for a dual vector of square $-2\chi(S)$ in order to find a torsion section.
The jacobian elliptic fibration is called extremal if the Picard number attains the maximum value $\rho(S)=h^{1,1}(S)$ while $\mathrm{MW}(S)$ is finite. Equivalently, $\mathrm{rank\,}\mathrm{Triv}(S)=h^{1,1}(S)$. For classifications, see [MP86] and [SZ01].
With respect to the geometric group structure, it follows that there is an injection
(2.11) $\mathrm{MW}(S)_{\mathrm{tor}}\hookrightarrow F^{\#}$
at any fibre. In general, this implies that $\mathrm{MW}(S)$ can be generated by at most 2 elements. At an additive fibre, (2.9) forces (2.11) to specialize to
(2.12) $\mathrm{MW}(S)_{\mathrm{tor}}\hookrightarrow A_{F}.$
Note that this implies that $\mathrm{MW}(S)_{\mathrm{tor}}=\{O\}$ if there is a fibre of type $\mathrm{II}$ or $\mathrm{II}^{*}$, and that $\mathrm{MW}(S)_{\mathrm{tor}}$ is cyclic if there is an additive fibre of type different from $\mathrm{I}_{n}^{*}$ for $n\in 2\mathbb{N}_{0}$.
With a view towards numerically trivial automorphisms, we record the following essential property:
Fact 2.3.
At an additive fibre, any two torsion sections meet different components.
Turning back to an arbitrary elliptic surface $S$ with $\chi(S)>0$, the relative Jacobian over $B$ associates to $S$ a jacobian elliptic surface which we denote by $J(S)$, or to keep track of the induced fibration,
$J(f)\colon\;\;J(S)\rightarrow B.$
Intriguingly, $S$ and $J(S)$ share many invariants, such as Betti numbers, Euler number, geometric genus, Euler characteristic, types of singular fibres (up to multiplicity, i.e. $J(S)$ has fibres of type $F^{\prime}$), and also the Picard number (cf. [CD89, §5.3]). It follows that (2.7) holds for $S$ with $\mathrm{MW}(S)$ replaced by $\mathrm{MW}(J(S))$. Another role of $\mathrm{MW}(J(S))$, this time for the automorphism group of $S$, will be explored in Section 4.
2.6. Local behavior of an automorphism around an $m\mathrm{I}_{k}$ fibre
In studying a fibration preserving automorphism, it is important to control the action around a fixed fibre. This kind of detailed local analysis has been used rather effectively by [Cai09] in order to give a global bound on the number of automorphisms acting trivially on the second cohomology group of an elliptic surface.
However, as our examples show, the crucial [Cai09, Lemma 1.5 (iii)] is not correct; the technical flaw in its proof seems to occur in the third paragraph on page 4232 of that paper: the automorphism $\bar{\alpha}$ of $\bar{S}_{\Delta}$, which was constructed by Kodaira, does not necessarily descend to $S_{\Delta}$.
As the reader can see, in Theorem 1.1, (iii), we show that $\mathrm{Aut}_{\mathbb{Q}}(S)$ is trivial if $p_{g}(S)>0$ and all multiple fibres of $S\to B$ have smooth support (in particular, if there are no multiple fibres, e.g. if $S$ is a jacobian elliptic surface).
In a sense, this result rescues Cai’s theorem up to some extent, showing that the only trouble occurs from the multiple fibres of type ${}_{m}I_{k}$. It is therefore interesting, also for future applications, to look which results hold true in the local analysis for these fibres.
Hence we provide in this subsection two nontrivial results in the flavour of [Cai09, Lemma 1.5].
Lemma 2.4.
Let $f\colon S\rightarrow\Delta$ be a relatively minimal elliptic fibration over the unit disk such that the central fibre $F_{0}=f^{*}0$ is the only singular fibre. Let $\sigma\in\mathrm{Aut}(S)$ be an automorphism of finite order preserving the fibration structure, so that it maps a fibre to another fibre and induces an automorphism $\sigma_{\Delta}\in\mathrm{Aut}(\Delta)$. Suppose that the following holds:
  1. (1)
    $\sigma$ and $\sigma_{\Delta}$ have the same order.
  2. (2)
    $F_{0}$ is a multiple fibre of type $m\mathrm{I}_{0}$.
Then $\sigma$ acts on $(F_{0})_{\mathrm{red}}$ by translations. In particular, we have $e(F_{0}^{\sigma})=0$.
Proof.
Let $k$ be the order of $\sigma$ and $\sigma_{\Delta}$. Then we can form the following diagram
[…]
where $\phi(u)=u^{m}$, $\psi(z)=z^{k}$, and $\nu$ is the normalization. When restricted to $\Delta\setminus\{0\}$, the horizontal maps are covering maps:
${\widetilde{S}\setminus\tilde{f}^{-1}(0)}$${S\setminus f^{-1}(0)}$${\bar{S}\setminus\bar{f}^{*}(0)}$${\Delta\setminus\{0\}}$${\Delta\setminus\{0\}}$${\Delta\setminus\{0\}}$$\scriptstyle{\tilde{f}}$$\scriptstyle{\Phi}$$\scriptstyle{f}$$\scriptstyle{\Psi}$$\scriptstyle{\bar{f}}$$\scriptstyle{\phi}$$\scriptstyle{\psi}$
Using the universal property of fibre products, one has an induced morphism
$\mu\colon\widetilde{S}\setminus\tilde{f}^{-1}(0)\rightarrow\bar{S}\setminus\bar{f}^{*}(0)\times_{\Delta,\psi\circ\phi}\Delta.$
Since $\deg\Phi=|\sigma|=|\sigma_{\Delta}|=\deg\phi$, one sees that $\deg\mu=1$ and hence $\mu$ is an isomorphism.
Since $\pi_{1}(\Delta\setminus 0)\cong\mathbb{Z}$, any of its covering space is regular with cyclic Galois group. In our case, the covering maps $\psi$, $\phi$, and $\psi\circ\phi$ are Galois with Galois groups:
$\mathrm{Gal}(\psi)=\langle\sigma_{\Delta}\rangle\cong\mathbb{Z}/k\mathbb{Z},\,\,\mathrm{Gal}(\phi)\cong\mathbb{Z}/m\mathbb{Z},\,\,\text{and }\mathrm{Gal}(\psi\circ\phi)\cong\mathbb{Z}/mk\mathbb{Z}$
It follows that the corresponding covering maps $\Psi$, $\Phi$, $\Psi\circ\Phi$ are also Galois with Galois groups:
$\mathrm{Gal}(\Psi)=\langle\sigma\rangle\cong\mathbb{Z}/k\mathbb{Z},\,\,\mathrm{Gal}(\Phi)\cong\mathbb{Z}/m\mathbb{Z},\,\,\text{and }\mathrm{Gal}(\Psi\circ\Phi)\cong\mathbb{Z}/mk\mathbb{Z}.$
Consequently, the branched covering $\widetilde{S}\rightarrow\bar{S}$ enjoys the same property.
Note that the central fibre $\tilde{F}_{0}$ of $\tilde{f}$ is a smooth elliptic curve. Since $F_{0}$ is of type $m\mathrm{I}_{0}$, the group $\mathrm{Gal}(\Phi)\cong\mathbb{Z}/m\mathbb{Z}$ acts on $\tilde{F}_{0}$ by translations. It is now evident that the group $\mathrm{Gal}(\Psi\circ\Phi)\cong\mathbb{Z}/mk\mathbb{Z}$ necessarily acts on $\tilde{F}_{0}$ by translations. It follows that $\mathrm{Gal}(\Psi)=\langle\sigma\rangle$ acts on the quotient $F_{0}$ by translations. ∎
Remark 2.5.
Assume that we have a singular fibre of type $mI_{k}$ of an elliptic fibration, and with $k\geq 1$. Then the automorphism group of the reduced fibre equals $(\mathbb{C}^{*})^{k}\rtimes D_{k}$, where $D_{k}$ is the dihedral group (of order $2k$), sitting in an exact sequence
$1\rightarrow\mathbb{Z}/k\rightarrow D_{k}\rightarrow\mathbb{Z}/2\rightarrow 1$
Lemma 2.6.
Let $f\colon S\rightarrow\Delta$ be a relatively minimal elliptic fibration over the unit disk such that the central fibre $F_{0}=f^{*}0$ is the only singular fibre. Let $\sigma\in\mathrm{Aut}(S)$ be an automorphism of finite order preserving the fibration structure, so it maps a fibre to another fibre and induces an automorphism $\sigma_{\Delta}\in\mathrm{Aut}(\Delta)$. Suppose that the following holds:
  1. (1)
    $\sigma$ and $\sigma_{\Delta}$ have the same order.
  2. (2)
    $F_{0}$ is a multiple fibre of type $m\mathrm{I}_{k}$, with $k\geq 0$.
Then $\sigma$ acts in such a way that the Euler characteristic of the fixed point set of $\sigma$ on $(F_{0})_{\mathrm{red}}$ equals $k$ or 00, and the first alternative holds if $\sigma$ preserves the components of $(F_{0})_{\mathrm{red}}$.
Proof.
Let $r$ be the order of $\sigma$ and $\sigma_{\Delta}$. Then again we can form the following diagram
[…]
where, after choosing a coordinate $z$ in which the action of $\sigma_{\Delta}$ is linear, $\phi(u)=u^{m}$, $\psi(z)=z^{r}$, and $\nu$ is the normalization. When restricted to $\Delta\setminus\{0\}$, the horizontal maps are covering maps:
${\widetilde{S}\setminus\tilde{f}^{-1}(0)}$${S\setminus f^{-1}(0)}$${\bar{S}\setminus\bar{f}^{*}(0)}$${\Delta\setminus\{0\}}$${\Delta\setminus\{0\}}$${\Delta\setminus\{0\}}$$\scriptstyle{\tilde{f}}$$\scriptstyle{\Phi}$$\scriptstyle{f}$$\scriptstyle{\Psi}$$\scriptstyle{\bar{f}}$$\scriptstyle{\phi}$$\scriptstyle{\psi}$
Using the universal property of fibre products, one has an induced morphism
$\mu\colon\widetilde{S}\setminus\tilde{f}^{-1}(0)\rightarrow\bar{S}\setminus\bar{f}^{*}(0)\times_{\Delta,\psi\circ\phi}\Delta.$
Since $\deg\Phi=|\sigma|=|\sigma_{\Delta}|=\deg\phi$, one sees that $\deg\mu=1$ and hence $\mu$ is an isomorphism.
Since $\pi_{1}(\Delta\setminus 0)\cong\mathbb{Z}$, any of its covering space is regular with cyclic Galois group. In our case, the covering maps $\psi$, $\phi$, and $\psi\circ\phi$ are Galois with Galois groups:
$\mathrm{Gal}(\psi)=\langle\sigma_{\Delta}\rangle\cong\mathbb{Z}/r\mathbb{Z},\,\quad\mathrm{Gal}(\phi)\cong\mathbb{Z}/m\mathbb{Z}\cong r\mathbb{Z}/mr\mathbb{Z}<\mathrm{Gal}(\psi\circ\phi)\cong\mathbb{Z}/mr\mathbb{Z}$
It follows that the corresponding covering maps $\Psi$, $\Phi$, $\Psi\circ\Phi$ are also Galois with with Galois groups:
$\mathrm{Gal}(\Psi)=\langle\sigma\rangle\cong\mathbb{Z}/r\mathbb{Z},\,\quad\mathrm{Gal}(\Phi)\cong\mathbb{Z}/m\mathbb{Z}<\mathrm{Gal}(\Psi\circ\Phi)\cong\mathbb{Z}/mr\mathbb{Z}.$
Consequently, the branched covering $\widetilde{S}\rightarrow\bar{S}$ enjoys the same property.
Denote by $\tau$ the generator of $\mathrm{Gal}(\Psi\circ\Phi)\cong\mathbb{Z}/mr\mathbb{Z}$, ($\tau$ induces $\sigma$).
Note that the central fibre $\tilde{F}_{0}$ of $\tilde{f}$ is a reduced fibre of type $\mathrm{I}_{mk}$.
Since $F_{0}$ is of type $m\mathrm{I}_{k}$, the group $\mathrm{Gal}(\Phi)\cong\mathbb{Z}/m\mathbb{Z}$ acts on $\tilde{F}_{0}$ mapping to $(\mathbb{C}^{*})^{mk}\rtimes D_{mk}$ in such a way that its projection in $D_{mk}$ is $k\mathbb{Z}/mk\mathbb{Z}<\mathbb{Z}/mk<D_{mk}$.
The group the group $\mathrm{Gal}(\Psi\circ\Phi)\cong\mathbb{Z}/mr\mathbb{Z}$, being a cyclic group, maps to $\mathbb{Z}/mk<D_{mk}$ (else, since $\mathbb{Z}/m\mathbb{Z}<\mathbb{Z}/mr\mathbb{Z}$, then $\mathbb{Z}/m\mathbb{Z}$ and a reflection would not generate a cyclic group, contradicting that every subgroup of a cyclic group is cyclic).
Hence we can write $\tau=\rho\tau^{\prime}$, where $\tau^{\prime}$ is a cyclical permutation; and since $\tau$ lies over the transformation $u\mapsto\epsilon u$, where $\epsilon$ is a primitive $mr$-th root of unity, it follows that $(\rho_{i+1}\rho_{i}^{-1})=\epsilon$. Since we have a cycle of curves, the previous equation implies $\epsilon^{mk}=1$, this means that $r|k$.
$\sigma$ acts on the set of components of the singular fibre $F_{0}$ via a rotation. If this rotation is nontrivial, then there are no fixed points on $F_{0}$.
If instead $\sigma$ preserves all the components, we know that it lies in $(\mathbb{C}^{*})^{k}$, hence it fixes all the nodes.
If $\sigma$ fixes another point, then it would act as the identity on some component of $(F_{0})_{\mathrm{red}}$, and there the Euler number of the fixed point set equals $e(\mathbb{P}^{1})=2$.
It suffices now to show that $\sigma$ does not act as the identity on two neighbouring components.
If in fact $\sigma$ acts on the $i$th component via $\beta_{i}$, then, since $\sigma(z)=\tau(z)=\epsilon^{m}z$, in view of the fact that in local coordinates at a node of $(F_{0})_{\mathrm{red}}$ it acts via multiplication by $(\beta_{i+1},\beta_{i}^{-1})$ we reach a contradiction by assuming that $\beta_{i+1}=\beta_{i}=1$.
First because we should have $(\beta_{i+1}\beta_{i}^{-1})^{m}=\epsilon^{m}$, absurd; second because $\sigma$, acting as the identity on the tangent space a fixed point, must be the identity. ∎
Remark 2.7.
Contrary to the statement in [Cai09, Lemma 1.5 (iii)], the fixed point set may contain not only points, but also fibre components, as will be inherent in Construction 7.5. In fact, we shall exploit this extensively when exhibiting cohomologically trivial automorphisms in Sections 10, 11.
3. Isotrivial fibrations and their cohomologically and numerically trivial automorphisms
From now on, unless otherwise noted, $S$ denotes a minimal properly elliptic surface with $\chi(S)>0$ while $B,C$ will be curves and $X$ can be any relatively minimal elliptic surface (often a jacobian surface, which in many cases may be rational).
In this section we assume that the pluricanonical fibration $f\colon S\rightarrow B$ is isotrivial.
This means that all the smooth fibres are isomorphic to each other and implies that there exists a $G$-Galois covering $p:C\rightarrow B=C/G$ such that the pull-back $C\times_{B}S$ is birational to a product $C\times E$, compatibly with the respective projections onto $C$ (see [Ser96], Section 2, and also the arguments in 1.5.3.1 and 1.5.3.2 of [BCP23]).
Remark 3.1.
$G$ acts on $C$ and on $E$, since $f$ is a holomorphic bundle with fibre $E$ on a Zariski open set $B^{*}$ of $B$, and the covering $p:C\rightarrow B=C/G$ of the base corresponds to the kernel of the monodromy with values in $\mathrm{Aut}(E)$.
$G$ acts faithfully on $C$ by construction, while we may assume that $G$ also acts faithfully on $E$, otherwise we can take $G^{\prime}:=\mathrm{Ker}(G\rightarrow\mathrm{Aut}(E))$ and replace $C$ by $C/G^{\prime}$.
Observe that $C$ has genus $g\geq 2$ since $S$ has Kodaira dimension $1$.
$G$ does not act freely on $C\times E$, otherwise we would have $\chi(S)=0$ and be in the situation considered in Part I, as referenced under [CFGLS24].
Since $G$ acts faithfully on the elliptic curve $E$, and not just via translations (since the action of $G$ on the product is not free), as in Part I we write
$G=T\rtimes\mu_{r},\ r\in\{2,3,4,6\},$
where $T$ is a finite group of translations of $E$ normalized by the group $\mu_{r}$ of $r$-th roots of unity.
We shall henceforth identify $G$ with the corresponding group of automorphisms of $E$, and in these terms we shall describe the monodromy homomorphism of the covering $C\rightarrow C/B=B$.
We write
$S\sim_{\mathrm{bir}}(C\times E)/\Delta_{G},$
and since by minimality of $S$ and since $Kod(S)=1$ the group of birational automorphisms of $S$, respectively $C\times E$, equals the group of biregular automorphisms, as in 3.7 of Part I we find that
(3.1) $\mathrm{Aut}(S)=N_{\Delta_{G}}/\Delta_{G},$
where $N_{\Delta_{G}}$ is the normalizer of the diagonal subgroup $\Delta_{G}<G\times G$ inside $\mathrm{Aut}(C\times E)$.
We continue to follow the notation of Part I: for $\Psi\in N_{\Delta_{G}}$ we write
$\Psi(x,z)=(\psi_{1}(x),\lambda z+\phi(x)).$
Remark 3.2.
Assume now that $\Psi$ yields a numerically trivial automorphism: then (as in section 4 of Part I), both fibrations, onto $B$, respectively onto $E/G\cong\mathbb{P}^{1}$ are preserved hence in particular $\phi(x)=c$, a constant.
Before focusing on the group $\mathrm{Aut}_{\mathbb{Z}}(S)$, we make some steps towards the classification of the possible groups $\mathrm{Aut}_{\mathbb{Q}}(S)$ in the isotrivial case.
First of all, some easy group theoretical consideration.
Assume that $\Psi$, represented by $(\psi_{1}(x),\psi_{2}(z))$, yields an element $\Psi_{S}\in\mathrm{Aut}_{\mathbb{Q}}(S)$.
Observe that if $g(z)=\epsilon z+t,\epsilon\in\mu_{r},\ t\in T$, then conjugation by $\psi_{2}$ sends $g$ to
$g^{\prime}:=\psi_{2}\circ g\circ\psi_{2}^{-1}=c+\lambda(g(\lambda^{-1}(z-c)))=\epsilon z+(1-\epsilon)c+\lambda t.$
Then $\psi_{2}(z)=\lambda z+c$ normalizes the action of $G$ on $E$ if and only if
(3.2) $\lambda T=T,\ \ \;(1-\epsilon)c\in T\;\;\forall\epsilon\in\mu_{r}.$
It follows, doing the same calculations performed in the proof of Lemma 4.3 of Part I, that the second part of (3.2) is equivalent
  • for $r=2$, to the condition $2c\in T$,
  • for a harmonic (Gauss) elliptic curve $E$, $r=4$, to $2c\in T$,
  • for an equianharmonic (Fermat) elliptic curve, $r=3$, to $3c\in T$,
  • for an equianharmonic (Fermat) elliptic curve, $r=6$, to $c\in T$, hence $\psi_{2}\in G$.
In fact, $\lambda$ has multiplicative order dividing respectively $2,4,6$, according to the case of a general elliptic curve, a harmonic one, an equianharmonic one: and then the condition $\lambda T=T$ holds automatically for $r\geq 3$ since by assumption $\mu_{r}(T)=T$.
3.1. Conditions for the numerical triviality of $\Psi_{S}$.
The singularities of $Y:=(C\times E)/G$ are cyclic quotient singularities with group cyclic of order $r^{\prime}$ dividing $r$.
For $r^{\prime}=2$, we get $A_{1}$ singularities, for $r^{\prime}=3$, a quotient singularity of type $A_{2}$ or one of type $\frac{1}{3}(1,1)$, for $r^{\prime}=4$, $A_{3}$ or $\frac{1}{4}(1,1)$, for $r^{\prime}=6$, $A_{5}$ or $\frac{1}{6}(1,1)$.
Take now the minimal resolution $Z$ of the singular quotient $Y:=(C\times E)/G$: then $Z\rightarrow B$ is an elliptic fibration, but not necessarily relatively minimal, and one has to successively contract the exceptional $(-1)$ curves to reach $S$, the relative minimal model.
Let us call for simplicity a fibre $F$ of $f\colon S\rightarrow B$ ‘singular’ if $F_{\mathrm{red}}$ is singular.
The ‘singular’ fibres are of type $\mathrm{I}_{0}^{*}$ for $r^{\prime}=2$, $\mathrm{IV}^{*}$ or $\mathrm{IV}$ if $r^{\prime}=3$, $\mathrm{III}^{*}$ or $\mathrm{III}$ if $r^{\prime}=4$, $\mathrm{II}^{*}$ or $\mathrm{II}$ if $r^{\prime}=6$ (see Table 5.2 on p.105 of [SS19]). If $\Psi_{S}$ is numerically trivial, then these singular fibres are left fixed, except possibly for the irreducible fibres, which are then of type $\mathrm{II}$. One may observe that $Z$ is obtained from $S$ via a canonical blow up procedure, making the fibre a normal crossing divisor. Hence, if the fibre is preserved, then the numerically trivial automorphism preserves all the exceptional components of $Z\rightarrow Y$.
In particular we conclude that
  1. (1)
    $\Psi:=(\psi_{1},\psi_{2})$ must act trivially on $H^{*}(C\times E,\mathbb{Q})^{G}$.
  2. (2)
    $\Psi_{Y}$ fixes each singular point of $Y$, except possibly for the points whose corresponding fibres are of type $\mathrm{II}$, which may be permuted,
  3. (3)
    for each singular point $y$ of $Y$ which is fixed, $\Psi_{Z}$ must preserve the irreducible components of the exceptional divisor in $Z$ over $y$.
  4. (4)
    Conversely, by the theorem of Mayer–Vietoris, if these conditions are satisfied, then $\Psi_{S}$ is numerically trivial.
The following Lemma will be quite useful concerning property (1) above.
Lemma 3.3.
Consider the factorization
$C\rightarrow D:=C/T\rightarrow B=C/G=D/\mu_{r},$
and set $E^{\prime}:=E/T$, $S^{\prime}:=(C\times E)/T$.
Then
$H^{*}(C\times E,\mathbb{Q})^{G}=[H^{*}(C\times E,\mathbb{Q})^{T}]^{\mu_{r}}=H^{*}(S^{\prime},\mathbb{Q})^{\mu_{r}}=[H^{*}(D,\mathbb{Q})\otimes H^{*}(E^{\prime},\mathbb{Q})]^{\mu_{r}}.$
Hence condition (1): $\Psi:=(\psi_{1},\psi_{2})$ acts trivially on $H^{*}(C\times E,\mathbb{Q})^{G}$ is automatically verified if $D$ has genus 00.
The case where $D,B$ have both genus $1$ cannot occur.
While, if $D$ has genus $\geq 2$, or $D$ has genus $=1$ (and $B$ has genus 00) no automorphism with $\psi_{1}=1$, and with $\psi_{2}$ not a translation, is numerically trivial, except for the case where $D$ has genus $\geq 2$, $r=6$ and $H^{1}(D,\mathbb{C})$ contains all nontrivial character spaces except those which correspond to a primitive character.
Finally, $D$ has genus 00 if and only if $D\rightarrow B=\mathbb{P}^{1}$ is branched only in two points with local monodromies $\epsilon,\epsilon^{-1}\in\mu_{r}$.
Proof.
The first statement is obvious since $T$ acts trivially on $H^{*}(E,\mathbb{Q})=H^{*}(E^{\prime},\mathbb{Q})$, and $H^{*}(D,\mathbb{Q})=H^{*}(C,\mathbb{Q})^{T}$.
For the second statement we observe that any product action is the identity on $H^{i}(D,\mathbb{Q})\otimes H^{2-i}(E^{\prime},\mathbb{Q})$ for $i=0,2$. Then for
${\mathcal{H}}:=[H^{1}(D,\mathbb{Q})\otimes H^{1}(E^{\prime},\mathbb{Q})]^{\mu_{r}}$
the second factor $H^{1}(E^{\prime},\mathbb{C})$ splits as $V_{1}\oplus V_{-1}$ according to the two characters $1,-1$ of $\mu_{r}$, considered as elements of $\mathbb{Z}/r\mathbb{Z}$.
If $D=\mathbb{P}^{1}$, then ${\mathcal{H}}=0$ and there is nothing to verify.
If $D$ has genus $\geq 2$, then since $\mu_{r}$ (by Lefschetz) acts faithfully on $H^{0}(\Omega^{1}_{D})$, this representation contains, if $r=2,3,4$ either $V_{1}$ or $V_{-1}$, hence $H^{1}(D,\mathbb{C})$ contains $V_{1}\oplus V_{-1}$. If $\psi_{1}=1$, then the action on $V_{1}\otimes V_{-1}$ is nontrivial if $\psi_{2}$ acts nontrivially on $H^{0}(\Omega^{1}_{E})$: hence $\psi_{2}$ must be a translation.
If instead $r=6$, the same argument applies if $H^{0}(\Omega^{1}_{D})$ contains a primitive character. There remains the case where $H^{1}(D,\mathbb{C})$ contains both $V_{2}\oplus V_{-2}$ and $V_{3}$ but not $V_{1}\oplus V_{-1}$.
A similar argument applies if $D$ has genus $1$ and $B=\mathbb{P}^{1}$. Since then $\mu_{r}$ acts on $D$ by a non translation, faithfully, hence in suitable coordinates $z\mapsto\lambda z$ with $\lambda$ a primitive r-th root of $1$; then we observe that, for $r=2,3,4,6$, $\lambda$ can only be $\epsilon,\epsilon^{-1}$.
If $D,B$ have genus $1$, then $\mu_{r}$ acts on $D$ via a translation, hence the action of $\mu_{r}$ on $D\times E^{\prime}$ is free, hence $C\times E\rightarrow S$ is unramified, a contradiction since $\chi(S)>0$.
Finally, if $D=\mathbb{P}^{1}$, an automorphism of order $r$ has necessarily the form $(x_{0},x_{1})\mapsto(x_{0},\lambda x_{1})$ in suitable coordinates, with $\lambda$ an r-th root of $1$.
Remark 3.4.
We observe (see Lemma 2.5 of Part I) that, if none of the local monodromies of $C\rightarrow C/G=B$ are translations, then $S$ has trivial torsion subgroup and $\mathrm{Aut}_{\mathbb{Q}}(S)=\mathrm{Aut}_{\mathbb{Z}}(S)$.
3.2. The case where $\psi_{1}$ is the identity
This is the case when the genus of $B$ is at least $2$, or if the genus of $B$ is $1$ and not all the ‘singular fibres’ are of type II (for instance, if not all the quotient singularities have cyclic group of order $r^{\prime}=6$); because then $\psi_{1}$ induces an isomorphism of $B$ acting trivially on the cohomology of $B$, and with a fixed point if $B$ has genus $1$: hence $\psi_{1}$ induces the identity of $B$ and we can assume that $\psi_{1}$ is the identity, by replacing $\Psi$ with another element in the coset $\Psi\Delta_{G}$.
If $r\neq 6$, then there are no fibres of type $\mathrm{II}$, hence all the ‘singular fibres’ are kept fixed. If moreover $B\cong\mathbb{P}^{1}$, then the covering $C\rightarrow B$ is branched in at least three points and if at least three local monodromies are not translations in $G$, then we have at least three ‘singular’ fibres and $\psi_{1}$ can be made to be the identity.
Observe that at least two local monodromies are not translations, hence there remains to consider the case of $2$ ‘singular’ fibres and $s\geq 2$ multiple fibres. Now, two multiple fibres cannot be permuted by a cohomologically trivial automorphism (see lemma 2.1 of Part I). Hence if $\Psi$ is cohomologically trivial, then $\psi_{1}$ maps to the identity of $B$ and $\psi_{1}$ can be made to be the identity.
Observe finally that if $\psi_{2}$ is a translation, and $\psi_{1}$ is the identity, then (1) is automatically verified. The case where $\psi_{2}$ is not a translation was analysed in Lemma 3.3.
An easy but crucial fact is the following.
Remark 3.5.
Assume that $\Psi_{S}\in\mathrm{Aut}_{\mathbb{Q}}(S)$ is induced by $\Psi:=(\mathrm{id}_{C},\psi_{2})$, and that we have a local monodromy $\sigma$ over the point $Gx$, and that $\sigma(y)=y$.
Then $(x,y)$ maps to a singular point $P$ of $Y$ and the fibre of the projection $Y\to E/G$ containing this singular point is in bijection with $E/\sigma$. Hence condition (2) that $\Psi_{Y}$ fixes $P$ is equivalent to the condition that $\psi_{2}(y)=y$.
If $\sigma$ is the stabilizer of $y$ under the $G$-action, this means that $\psi_{2}$ must belong to the subgroup generated by $\sigma$.
3.3. Cohomologically trivial automorphisms
The main result of this section proves and sharpens Theorem 1.5.
Theorem 3.6.
If $S$ has $\chi(S)>0$, the pluricanonical fibration
$f\colon S\rightarrow B$
is isotrivial, then $|\mathrm{Aut}_{\mathbb{Z}}(S)|\leq 3$.
More precisely, if the group $G\cong T\rtimes\mu_{r}$, then first of all $|\mathrm{Aut}_{\mathbb{Z}}(S)|\leq r$, but indeed, if $r=6$, then $\mathrm{Aut}_{\mathbb{Z}}(S)$ has order at most $3$, if $r=4$, then $\mathrm{Aut}_{\mathbb{Z}}(S)$ has order at most $2$, and if $r=3$ then $\mathrm{Aut}_{\mathbb{Z}}(S)$ is trivial.
For $r=4$ there are infinitely many families with an automorphism $\Psi$ of order $4$ which is numerically trivial, but not cohomologically trivial, but such that $\Psi^{2}$ is cohomologically trivial.
For $r=6$, there are infinitely many families with $|\mathrm{Aut}_{\mathbb{Z}}(S)|=3$.
Proof.
We have just seen that $\mathrm{Aut}_{\mathbb{Z}}(S)$ acts trivially on $C/G=B$ for $r\neq 6$.
We shall prove in 3.4 that this also happens for $r=6$, so let us assume this property from now on.
Consider the fibration $S\rightarrow E/G=\mathbb{P}^{1}$. The bona fide singular fibres (they can be also multiple) occur over the branch points of $E\rightarrow E/G$, but these are $4$ for $r=2$, $3$ for $r=3$, $3$ for $r=4$, but with different branching indices $(4,4,2)$ , three for $r=6$ with branching indices are $(2,3,6)$.
For $r=2$ the corresponding fibres correspond to different components of a fibre of type $\mathrm{I}_{0}^{*}$, hence they cannot be permuted by a numerically trivial automorphism. A similar picture holds also for $r=3$ and $r=4$ since we have observed that the singular points of $Y$ on a fibre (with $r^{\prime}=r$) of type different from $\mathrm{II}$ must be kept fixed.
For $r=6$ we use simply that the branching indices are different.
The conclusion is that $\mathrm{Aut}_{\mathbb{Q}}(S)$ acts trivially on $E/G$.
Hence
$\mathrm{Aut}_{\mathbb{Z}}(S)\subset N_{G\times G}(\Delta_{G})/\Delta_{G}\cong Z(G),$
where $Z(G)$ is the centre of $G$.
We apply now Lemma 4.3 on page 19 of Part I, classifying the possible groups $Z(G)$: these have cardinality at most $9$.
A basic observation is moreover that, since there is an additive fibre for $f$, whose components cannot be permuted, then if $\Psi=(\mathrm{id},\psi_{2})\in\mathrm{Aut}_{\mathbb{Q}}(S)$, then $\psi_{2}$ cannot be a translation (see Lemma 4.3).
We apply now Lemma 3.3 and conclude that $D$ has genus 00 if $r\neq 6$.
For $r=6$, we use that $\mu_{6}$ centralizes $G$ only if $G=\mu_{6}$.
Assume now that $G=\mu_{6}$: in this case we have $C=D$, and $S$ has no multiple fibres.
Assume now that $\psi_{2}$ is a generator of $\mu_{6}$: then all the local monodromies are generators of $\mu_{6}$ by Remark 3.5.
We need next to verify property (2), that $\Psi$ fixes the singular points of $Y$.
Consider a fibre $F$ of $f$ image of $\{x\}\times E$ where the stabilizer of $x$ is $G$: then $\Psi$ is the identity on $E/G$ and property (2) is verified.
We may assume without loss of generality that $G$ acts by the character $1$ on the tangent space at $x$, that is, there is a local coordinate $w$ such that $w\mapsto\epsilon w$.
Take a nonzero element $\omega\in H^{0}(\Omega^{1}_{C})$ which belongs to a character space $V_{j}$. We have already found that $j\neq 1,-1$. In the local coordinate $w$, we can write $\omega=\phi(w)dw$, hence the generator of $G$ sends
$\omega=\phi(w)dw\mapsto\phi(\epsilon w)\epsilon dw=\epsilon^{j}\phi(w)dw,$
therefore $\phi$ and $\omega$ vanish at the point $x$.
The conclusion is that all holomorphic 1-forms $\omega$ vanish on $x$, which is impossible since $C$ has positive genus.
Hence for $r=6$ we must necessarily have that $\mathrm{Aut}_{\mathbb{Z}}(S)$ has order at most $3$.
For $r=2,3,4$ we conclude immediately that the group $\mathrm{Aut}_{\mathbb{Z}}(S)$ has order at most $r$.
3.4. Proof that the action on the base $B$ is trivial also in the case $r=6$.
If the action on $B$ is nontrivial, $B$ has genus $\leq 1$.
As in the proof of Proposition 6.7, since $\chi(S)>0\Leftrightarrow e(S)>0$, the automorphism induced by $\Psi$ has some fixed points. If the fixed set $\Sigma$ projects onto $B$, then clearly $\Psi$ acts as the identity on $B$. If $\Sigma$ maps onto a finite set, the action on $B$ is a translation admitting fixed points, hence it is the identity.
Proposition 6.7 shows that also if the base $B$ is $\mathbb{P}^{1}$ and $\chi(S)\geq 2$ the action on the base is the identity.
There remains the case where $q(S)=p_{g}(S)=0\Rightarrow e(S)=12$, in this case $B$ is $\mathbb{P}^{1}$.
By Kodaira’s canonical bundle formula (2.4), $K_{S}$ is given as the pull back of a divisor of degree $-1$ on $\mathbb{P}^{1}=B$ plus the contribution $\sum_{j}\frac{m_{j}-1}{m_{j}}F_{j}$ over the multiple fibres, hence in our case of Kodaira dimension $Kod(S)=1$ there are at least two multiple fibres.
By the cited Lemma 2.1 of Part I local monodromies which are translations (hence correspond to smooth multiple fibres) yield fixed points for the action on $B$ of the automorphism $\sigma$ induced by $\psi_{1}$, and we have here seen that the same holds true for other local monodromies which are not of the form
(3.3) $z\mapsto\epsilon z+t,\ \epsilon=\exp\left(\frac{1}{3}\pi i\right),\,t\in E;$
in fact, the local monodromy transformation is acting around a fixed point in $C$ via the root of unity with minimal argument. Therefore, if the monodromy value in $G$ is of the form $z\mapsto\epsilon z+t$, then at the fixed points the 2 eigenvalues are the same and we get a singularity of type $\frac{1}{r}(1,1)$.
We can therefore assume that the number of local monodromies which are not of the above form (3.3) are at most two, hence the number of monodromies which are translations equals exactly $2$ and the other ones are of type (3.3).
[…]
Hence we have $6$ points with monodromy of type (3.3), and $2$ points whose monodromies are translations.
The set ${\mathcal{D}}$ of $6$ points with monodromy (3.3) is permuted by $\sigma$, but two points can be exchanged only if they have the same monodromy.
Now $\sigma$ fixes the two points with local monodromy equal to a translation, hence $\sigma$ belongs to a cyclic group $\subset\mathbb{C}^{*}$, permuting the set ${\mathcal{D}}$, and moreover all the orbits different from the two fixed points of $\sigma$ have the same cardinality.
Hence $\sigma$ belongs to a subgroup of the cyclic group $\mathbb{Z}/6$, and let us assume that $\sigma$ generates the subgroup $H$ induced from the first components $\psi_{1}$ of cohomologically trivial automorphisms of $S$; we denote by $d$ the order of $\sigma$, hence $d\in\{2,3,6\}$ if $\sigma$ is nontrivial.
We can choose coordinates such that $H$ acts on $\mathbb{P}^{1}$ via $x\mapsto\zeta x$, with $\zeta$ the primitive $d$-th root of $1$ with smallest argument.
Let us exclude first the case where $d=6$. For $d=6$, then all the monodromies of points of ${\mathcal{D}}$ are equal (and changing the origin in the elliptic curve $E$ we can make them to be of the form $\epsilon z$) hence the two monodromies which are translations are equal, and they must be of the form $t,-t$, where $t=\frac{1-\epsilon^{2}}{3}$.
Consider again the factorization used in Lemma 3.3
$C\rightarrow D:=C/T\rightarrow D/\mu_{6}=C/G=B=\mathbb{P}^{1}.$
Since all the $\mu_{6}$ valued monodromies are equal, $D$ is the familiar Fermat sextic, whose equation in affine coordinates equals
$\{(x,z)\mid F(x,z):=z^{6}-x^{6}+1=0\}.$
The space of holomorphic of differentials $H^{0}(D,\Omega^{1}_{D})\cong H^{0}(\mathcal{O}_{D}(3))$ has as basis
$\mathrm{Res}\left(\frac{1}{F}\ dx\wedge dz\right)\ (z^{j}x^{i}),\ i+j\leq 3.$
There are four character spaces as a representation of $\mu_{6}$, $V_{1},V_{2},V_{3},V_{4}$ and a basis element as above belongs to $V_{1+j}$. While the character of the group $H$ on the same basis element equals $1+i$.
Hence
$[(H^{0}(D,\Omega^{1}_{D})\oplus\overline{H^{0}(D,\Omega^{1}_{D})})\otimes H^{1}(E,\mathbb{C})]^{\mu_{6}}=[V_{1}\otimes\overline{H^{0}(E,\Omega^{1}_{E})}]\oplus[\overline{V_{1}}\otimes H^{0}(E,\Omega^{1}_{E})].$
Whatever the action of $\psi_{2}$ on $H^{0}(E,\Omega^{1}_{E})$, the automorphism $\Psi$ shall never act trivially on this subspace of the $G$-invariant cohomology, since $V_{1}$ contains $4$ different character spaces for the action of $\sigma$.
Hence this action is not numerically trivial and the case $d=6$ is excluded.
We proceed quite similarly with the cases $d=2,3$: then again, since we have $6$ monodromies of type (3.3), we have a similar factorization, and we look again at the subspace of the cohomology of $C\times E$ given by the pull back of the cohomology of $D\times E$. The character spaces as a representation of $\mu_{6}$ are the same as before (just changes the equation $F$), and we must see how $H=<\sigma>$ acts.
For $d=3$, on $V_{1}$ we have the characters $1,2,0$ of $H$, and again we have a contradiction; for $d=2$, we have the characters $1,0$ of $H$, hence this case is also excluded. ∎
3.5. Proof of Theorem 3.6 for $r=3$
We show now that the case of $\mathrm{Aut}_{\mathbb{Z}}(S)$ of order $r=3$ does not exist.
Theorem 3.7.
For each isotrivially fibred properly elliptic surface $S$ with $q(S)=p_{g}(S)=0$ and $r=3$, the group $\mathrm{Aut}_{\mathbb{Z}}(S)$ is trivial.
Proof.
We saw in Theorem 3.6 that, for $r=3$, $\mathrm{Aut}_{\mathbb{Z}}(S)$ has order $\leq 3$ and, if equality holds, then $\mu_{3}$ centralizes $G$, hence $G=T\times\mu_{3}$ where $T=\mathbb{Z}/3$.
We take generators $\sigma\in\mu_{3}$ such that $\sigma(z)=\epsilon z$, and $\tau$ such that $\tau(z)=z+\eta$, where $\eta=\frac{1-\epsilon}{3}$.
Without loss of generality, if $\mathrm{Aut}_{\mathbb{Z}}(S)$ has order three, we can assume that it is generated by $\Psi$, the automorphism of $S$ induced by $\mathrm{id}\times\sigma$ on $C\times E$.
By Remark 3.5 we may take $C$ to be the $G$-covering of $\mathbb{P}^{1}$ branched in $2+n_{1}+n_{2}=2+n$ points, $x_{1},x_{2},\xi_{1},\dots,\xi_{n_{1}},\xi_{n_{1}+1},\dots,\xi_{n}$, with local monodromies
$(\sigma,\sigma^{-1},\tau,\dots,\tau,\tau^{-1},\dots,\tau^{-1}).$
$S$ has $q(S)=p_{g}(S)=0$.
On $E$, the elements with fixed points are $\sigma,\sigma^{\prime}:=\sigma\tau,\sigma^{\prime\prime}:=\sigma\tau^{2}$ (the last is the inverse of $\sigma^{-1}\tau$) and their inverses: each has a triple of fixed points, mapping to three points $y_{1},y_{2},y_{3}$ of $E/G=\mathbb{P}^{1}$, which are the three branch points, with local monodromies $(\sigma,\sigma^{\prime},\sigma^{\prime\prime})$.
$Y=(C\times E)/G$ has three singular points of type $\frac{1}{3}(1,1)$ lying over $(x_{1},y_{1})$ and three singular points of type $A_{2}$ lying over $(x_{2},y_{1})$.
Let $\hat{S}$ be the minimal resolution of singularities of $Y$: then the fibre of $\hat{S}$ over $x_{2}$ contains three $A_{2}$ configurations, hence it is a fibre of type $\mathrm{IV}^{*}$ ($\tilde{E_{6}}$).
The fibre over $x_{1}$ is not minimal, since the proper transform of the fibre of $Y$ over $x_{1}$ is a $(-1)$-curve $C_{0}$; contracting it, we get a fibre of type $\mathrm{IV}$, with three components $\Delta_{0},\Delta_{1},\Delta_{2}$ meeting in a point $P\in S$.
Through this point $P$ pass all the fibres of the second projection (the rational map $\pi:S\dasharrow E/G$). We denote by $F^{\prime}_{1},\dots,F^{\prime}_{n}$ the divisors such that $3F^{\prime}_{j}$ is the fibre of $f$ over $\xi_{j}$, while $\Phi_{2}$ is such that $3\Phi_{2}$ is the fibre of $\pi$ over $y_{2}$.
Since $|G|=9$, it follows that $F^{\prime}:=F^{\prime}_{1},\Phi:=\Phi_{2}$ generate a unimodular odd sublattice $U^{\prime}\subset\mathrm{Num}(S)$; moreover $e(S)=12\Rightarrow b_{2}(S)=10$, and we choose the divisors $\Delta^{\prime}_{1}:=\Delta_{1}-F^{\prime},\Delta^{\prime}_{2}:=\Delta_{2}-F^{\prime}$ to generate an $A_{2}$ lattice, and since the fibre of type $\mathrm{IV}^{*}$ consists of a central curve ${\mathcal{X}}$ meeting $A^{\prime}_{1},A^{\prime}_{2},A^{\prime}_{3}$, with tails $A^{\prime\prime}_{1},A^{\prime\prime}_{2},A^{\prime\prime}_{3}$ we omit $A^{\prime\prime}_{2}$ in order to obtain an $E_{6}$ lattice. To make it orthogonal to $U^{\prime}$ we replace ${\mathcal{X}}$ by ${\mathcal{X}}^{\prime}:={\mathcal{X}}-F^{\prime}$.
Both $A_{2}$ and $E_{6}$ have discriminant group $\mathbb{Z}/3$, hence the orthogonal direct sum
$\Lambda^{\prime}=U^{\prime}\oplus A_{2}\oplus E_{6}\subset\Lambda=\mathrm{Num}(S)=H^{2}(S,\mathbb{Z})/\operatorname{Tors}$
has index three.
The choice of $\Lambda$ amounts to the choice of an isotropic subgroup $\cong\mathbb{Z}/3$ in the discriminant group of the non unimodular lattice $U^{\prime}\oplus A_{2}\oplus E_{6}$, which is the orthogonal direct sum $\mathbb{Z}/3\oplus\mathbb{Z}/3$.
We take as generator of the discriminant group $(A_{2})^{\vee}/A_{2}$ the linear form $v_{1}$ dual to $\Delta^{\prime}_{2}$, which is represented by the divisor $-\frac{1}{3}(\Delta^{\prime}_{1}+2\Delta^{\prime}_{2})$.
Whereas we take as generator of the discriminant group $(E_{6})^{\vee}/E_{6}$ the linear form $v_{2}$ dual to $A^{\prime\prime}_{1}$, which is represented by the divisor
$-\frac{1}{3}(4A^{\prime\prime}_{1}+5A^{\prime}_{1}+6{\mathcal{X}}^{\prime}+3A^{\prime}_{2}+4A^{\prime}_{3}+2A^{\prime\prime}_{3}).$
An easy calculation shows that $v_{1}^{2}=-\frac{2}{3},v_{2}^{2}=-\frac{4}{3}$, hence all such isotropic subspaces are given by the multiples of $(v_{1},v_{2})$ or of $(-v_{1},v_{2})$.
One sees however that the second choice amount to exchanging $\Delta_{1}$ with $\Delta_{2}$. Hence we go for the first choice, and let $v$ denote the cohomology class in $\Lambda$ corresponding to the pair $(v_{1},v_{2})$ (hence $v^{2}=-2$).
Note that $\operatorname{Tors}(S)\cong(\mathbb{Z}/3)^{n-1}$, generated by $F^{\prime}_{1}-F^{\prime}_{2},F^{\prime}_{1}-F^{\prime}_{3},\dots,F^{\prime}_{1}-F^{\prime}_{n}$.
Lemma 3.8.
There is an $N\in\mathbb{Z}$ with $N\leq n-1$ such that
$D:=D_{N}:=v+\Phi+NF^{\prime}$
satisfies $D\geq 0$ and $|D|=\{D\}$.
Proof.
We first prove that for $N=n-1$ we get $D_{n-1}\geq 0$. To see this, we observe some easy formulae, the first following from Kodaira’s canonical bundle formula (2.4),
$K_{S}=2(F^{\prime}_{1}+\dots+F^{\prime}_{n})-F\sim(2n-3)F^{\prime},\;\ v^{2}=-2,\;v_{1}\Phi=v_{2}\Phi=v\Phi=0,\;F^{\prime}v=0,$
hence
$D_{n-1}^{2}=(2n-3)=K_{S}D_{n-1}.$
By Riemann-Roch (since $\chi(S)=1$) we obtain $h^{0}(D_{n-1})+h^{0}(K_{S}-D_{n-1})\geq 1$.
Since $(K_{S}-D_{n-1})F=-3$ and $F$ is nef, we conclude that $h^{0}(K_{S}-D_{n-1})=0$, hence $D_{n-1}$ is an effective divisor.
To arrange for the second claim, it suffices to find some $N\leq n-1$ such that $h^{0}(D_{N})=1$. For this purpose, consider now the following short exact sequence:
$0\rightarrow\mathcal{O}_{S}(D_{k-1})\rightarrow\mathcal{O}_{S}(D_{k})\rightarrow\mathcal{O}_{F^{\prime}}(D_{k})\rightarrow 0$
Since the cokernel is a line bundle of degree $1$ on the elliptic curve $F^{\prime}$, Riemann–Roch gives $h^{0}(\mathcal{O}_{F^{\prime}}(D_{k}))=1,$ hence
$h^{0}(\mathcal{O}_{S}(D_{k}))-h^{0}(\mathcal{O}_{S}(D_{k-1}))\leq 1.$
Therefore there exists a minimal integer $N\in\mathbb{Z}$ such that
$h^{0}(\mathcal{O}_{S}(D_{N}))=1,\ h^{0}(\mathcal{O}_{S}(D_{N-1}))=0.$
as claimed.
Remark 3.9.
We emphasize that $N$ is the minimal integer such that $D_{N}\geq 0$, as we shall use again in 3.6 and then in Sections 10, 11.
Note that $D$ is not irreducible, since we have
(3.4) $D\Delta_{j}=j,j=0,1,2,\ D{\mathcal{X}}=1,DA^{\prime\prime}_{1}=1,DA^{\prime\prime}_{2}=-1,$
while for the other components of the fibre of type $\mathrm{IV}^{*}$ the intersection number is 00.
Hence $D$ contains $A^{\prime\prime}_{2}+A^{\prime}_{2}$ as subdivisor, and $D^{\prime}:=D-A^{\prime\prime}_{2}-A^{\prime}_{2}$ has intersection numbers all equal to zero with the components of the fibre of type $\mathrm{IV}^{*}$, except $D^{\prime}A^{\prime}_{2}=D^{\prime}A^{\prime\prime}_{1}=1.$
By our choice of the minimality of $N$ the vertical part of $D^{\prime}$ does not contain $F^{\prime}$, and a fortiori no full fibre.
Lemma 3.10.
Either $D^{\prime}$ is irreducible, or it is a sum of effective divisors $D^{\prime}=D_{0}+{\mathcal{E}}+\Delta+F^{\prime\prime}$, where $D_{0}$ is irreducible, $F^{\prime\prime}$ is a sum of submultiple fibres $F^{\prime}_{j}\,(j>1)$, ${\mathcal{E}}$ is supported on the fibre of type $\mathrm{IV}^{*}$, but is not equal to a multiple of this fibre, and similarly $\Delta$ is supported on the fibre of type $\mathrm{IV}$, but is not equal to a multiple of this other fibre.
Proof of the Lemma.
Since the surface $S$ has multiple fibres, hence does not have a section, the horizontal component $D_{0}$ of $D^{\prime}$ (which is a trisection different from $\Phi$ (otherwise $v\in\Lambda^{\prime}$)) is irreducible. Each other component is vertical, hence it is either a submultiple fibre different from $F^{\prime}_{1}$, or is contained in the two fibres with singular support. Here the components (with multiplicities) never sum up to a full fibre by the minimal choice of $N$ in defining $D$.
Claim 3.11.
$\Psi(D_{0})\neq D_{0}$.
Conclusion of proof of Theorem 3.7
Given this, since $|D_{0}|=\{D_{0}\}$, we conclude that $\Psi(D_{0})$ is not linearly equivalent to $D_{0}$: since $q(S)=0$, we can conclude that $\Psi$ is not cohomologically trivial, thus proving Theorem 3.7. ∎
Proof of Claim 3.11.
Assume in fact that $\Psi(D_{0})=D_{0}$: then the sets $D_{0}\cap\Delta_{j},D_{0}\cap A^{\prime}_{j}$ and $D_{0}\cap A^{\prime\prime}_{j}$ are left invariant by $\Psi$. It will be sufficient to argue at the fibre of type $\mathrm{IV}$, but a similar argument works at the $\mathrm{IV}^{*}$ fibre.
By Lemma 3.10, there are 3 alternatives for the support of $D^{\prime}$ on the $\mathrm{IV}$ fibre:
$\Delta=0,\;\;\;\Delta=\Delta_{0},\;\;\;\Delta=\Delta_{0}+\Delta_{1}.$
Because, in the cases $\Delta_{1},\Delta_{2},\Delta_{1}+\Delta_{2}$ the intersection number of $D_{0}$ with $\Delta_{0}$ would be negative, a contradiction. Similarly we can exclude the case where the support equals $\Delta_{0}+\Delta_{2}$, since then the intersection number of $D_{0}$ and $\Delta_{1}$ would be negative.
In each case, we read off that one component is disjoint from $D_{0}$ (which therefore does not meet the node of the fibre) and the other two fibre components are met with multiplicity one resp. two. Up to symmetry, it thus suffices to study the case from (3.4) with $D^{\prime}$ and $D_{0}$ having the same intersection numbers $D^{\prime}\Delta_{j}=D_{0}\Delta_{j}=j$.
It remains to use that $\Psi$ acts on each component $\Delta_{j}$ with exactly two fixed points, namely the node and one other point $P_{j}$, say. To see this, note that the curve $C_{0}$ contracted to a point in $S$ is fixed pointwise under $\Psi$ (since it meets the strict transforms of the $\Delta_{j}$ in three distinct points), hence the $\Delta_{j}$ cannot be fixed pointwise (this follows also from an easy local calculation).
From $\Psi(D_{0})=D_{0}$, we thus infer that $D_{0}\cap\Delta_{j}=\{P_{j}\}$ for $j=1,2$, with tangential intersection at $P_{2}$ which we shall now use to establish a contradiction.
Take local coordinates $(x,t)$ such that $t$ is the lift of a local coordinate on the base curve $B$, where $\Psi$ acts as the identity.
Hence $\Psi^{*}(t)=t$, while $\Psi^{*}(x)=\epsilon x$, for $\epsilon$ a generator of $\mu_{3}$.
If $D_{0}$ is smooth and has the same tangent as $\Delta_{2}=\{t=0\}$. then $D_{0}=\{(x,t)|t=F(x)\}$. By $\Psi$-invariance, $F(x)$ contains only powers of $x$ with exponent divisible by $3$, hence $(D_{0}\cdot\Delta_{2})_{P_{2}}\geq 3$, a contradiction.
If instead $D_{0}$ has a double point, its quadratic part is an eigenvector, hence either $t^{2}$, or $tx$, or $x^{2}$. In the first case the equation of $D_{0}$, modulo the ideal generated by $t$, contains only powers of $x$ with exponent divisible by $3$, hence $D_{0}\cdot\Delta_{2}\geq 3$, the same contradiction.
In the second case the other terms, modulo $t$, contain $x$ with exponent congruent to $1$ modulo $3$, and at least $3$, hence $D_{0}\cdot\Delta_{2}\geq 4$, contradiction.
In the third case the equation of $D_{0}$ contains only monomials where $x$ has exponent congruent to $2$ modulo $3$, hence the equation is divisible by $x^{2}$ and $D_{0}$ is not irreducible and reduced.
Obviously $D_{0}$ cannot have multiplicity $\geq 3$ at $P_{2}$, hence, all in all, we have derived the required contradiction to the assumption $\Psi(D_{0})=D_{0}$; this completes the proof of Claim 3.11.
3.6. Proof of Theorem 3.6 for $r=4$
We devote ourselves now to the case $r=4$. If $\mathrm{Aut}_{\mathbb{Z}}(S)$ has order $4$, then $\mu_{4}$ centralizes $G$, hence $T\cong\mathbb{Z}/2$ and it is generated by $t_{0}:=\frac{1+i}{2}$ ($G=T\times\mu_{4}$).
It follows that $D=\mathbb{P}^{1}$, and that $C$ is a double covering of $D$ branched in $4m$ points, with $m\geq 2$.
By looking at the two reducible fibres we see that $b_{2}(S)=10$, hence $q(S)=p_{g}(S)=0$.
In view of Remark 3.5 there is only one possibility for $\mathrm{Aut}_{\mathbb{Z}}(S)$ to be of order $4$ (hence $r=4$), if, after a change of coordinates in the elliptic curve $E$, we let $\Psi$ be induced by the identity on $C$ and by $\psi_{2}(z)=\sigma$ where $\sigma(z):=iz$.
The monodromy with values in $G=T\times\mu_{4}$ (up to a different order of the branch points) must be: $(\sigma,\sigma^{-1},\tau_{0},,\dots,\tau_{0})$ (where $\tau_{0}$, denoting translation by $t_{0}$, appears $(m=2s)$ times).
To understand the elliptic fibration $f:S\to\mathbb{P}^{1}$, we first consider the singular quotient
$S_{0}=(C\times E)/G\stackrel{{\scriptstyle f_{0}}}{{\longrightarrow}}C/G\cong\mathbb{P}^{1}.$
Here $f_{0}$ has double fibres $F_{i}=2F_{i}^{\prime}$ with smooth support at the $2s$ points with monodromy $\tau_{0}$, and a 4-fold rational curve $C_{0}$ at the other two branch points. The fibre above each of these two points contains 3 surface singularities (coming from the three orbits of 2-torsion points of $E$ under $\mu_{4}$), but the types are different for $\sigma$ and $\sigma^{-1}$:
  • $2\times A_{3}+1\times A_{1}$, resulting in a fibre of type $\mathrm{III}^{*}$ on the minimal resolution $\tilde{S}_{0}$;
  • $2\times\frac{1}{4}(1,1)+1\times A_{1}$, resulting in the following configuration on $\tilde{S}_{0}$:
[…]
Here $C_{2}$ has multiplicity two as fibre component and the self-intersection numbers are indicated to the left. Contracting first $C_{0}$ and then the image of $C_{2}$, we obtain the desired elliptic surface $S$ with a fibre of type $\mathrm{III}$ (and $S$ is Enriques, if $s=1$, and properly elliptic for $s>1$). This admits two natural bisections $O_{S},R_{S}$ induced by the 2-torsion of $E$. On the singular model $S_{0}$, $R_{S}$ meets exactly the singularities of type $A_{1}$ (induced by the 2-torsion points $\frac{1}{2},\frac{i}{2}$), and $O_{S}$ the other two, induced by 00 and $t_{0}$. On $S$, it follows that $O_{S}$ meets the $\mathrm{III}^{*}$ fibre transversally in the simple components $\Theta_{0},\Theta_{6}$, and $\mathrm{III}$ in one point (different from the node) on each component $C_{1},C_{3}$, while $R_{S}$ intersects $\mathrm{III}$ in the node (which $C_{2}$ was contracted to) and $\mathrm{III}^{*}$ in the central double component $\Theta_{7}$, as depicted in the following dual graph:
[…]
Then properties (1) and (2) are verified and $\Psi$ induced by $(\mathrm{id},\psi_{2})$ generates a subgroup isomorphic to $\mu_{4}$ which acts numerically trivial, and trivially on the first homology group since $q(S)=0$ and by Lemma 2.5 of Part I [CFGLS24].
The submultiple fibres of both fibrations, and the other fibre components (exceptional over $Y$), on which $\Psi$ acts trivially, do not induce a system of generators of $\mathrm{Num}(S)=H^{2}(S,\mathbb{Z})/\operatorname{Tors}$, but generate an index two subgroup.
To see this, just consider the unimodular lattice $U=\langle F_{1}^{\prime},O_{S}\rangle$ and realize that root lattices $A_{1}\oplus E_{7}$ supported naturally on the singular fibres by omitting $\Theta_{0},C_{3}$ as in (2.8) embed into $U^{\perp}$ once we subtract $F_{1}^{\prime}$ from the components $\Theta_{6},C_{1}$ met by $O_{S}$:
$\mathrm{Num}(S)\supset U\oplus A_{1}\oplus E_{7}.$
One can complement this sublattice to $\mathrm{Num}(S)$ uniquely by adding a vector $u$ which is the sum of the generators of the discriminant groups $A_{1}^{\vee}/A_{1}$ and $E_{7}^{\vee}/E_{7}$. Note that we may assume that $u^{2}=-2$. We are now in the position to prove the remaining statements of Theorem 3.6.
Claim 3.12.
$\Psi^{2}$ is cohomologically trivial.
Proof.
Since $\Psi$ is numerically trivial, we have that $\Psi(u)=u+\eta$ where $\eta\in\operatorname{Pic}(S)$ is a torsion element (of order $1$ or $2$). It follows then that $\Psi^{2}$ is cohomologically trivial, since $\Psi^{2}(u)=\Psi(u+\eta)=u+2\eta=u$.
Claim 3.13.
$\Psi$ is not cohomologically trivial.
Proof.
We have to prove to $\Psi(u)\neq u$ in $\operatorname{Pic}(S)$. To this end, we replace $u$ by the following auxiliary divisor
$D_{N}=u+O_{S}+\Theta_{6}+NF_{1}^{\prime}$
which is effective for $N\geq s$ by Riemann–Roch, using that $g(O_{S})=s-1$ by Riemann–Hurwitz, so $O_{S}^{2}=-2$ by adjunction and $D_{N}^{2}=2(N-1)$, while $D_{N}.K_{S}=O_{S}.K_{S}=2(s-1)$. Picking the minimal $N\leq s$ such that $D_{N}\geq 0$, and setting $D=D_{N}$, we can ensure that $|D|=\{D\}$ as in Lemma 3.8. Hence $\Psi(u)=u$ in $\operatorname{Pic}(S)$ is equivalent to $\Psi(D)=D$ as divisors.
Since $D.F_{1}^{\prime}=1$, the support of $D$ contains an irreducible bisection $C\neq O_{S},R_{S}$, but it may still include some fibre components (possibly multiple times), i.e.
$C=D-\sum_{i}\Theta_{i}.$
Regardless of the precise details, this implies that
$C^{2}\leq D^{2},\;\;\;C.O_{S}\leq D.O_{S}=N-1\;\;\;\text{ and }\;\;\;C.R_{S}\leq D.R_{S}.$
We shall now compare $C$ against the fixed loci of $\Psi$ and of $\Psi^{2}$. Starting with the latter, we have
$\mbox{Fix}({\Psi^{2}})=\Theta_{1}\cup\Theta_{3}\cup\Theta_{5}\cup O_{S}\cup R_{S}\cup U$
where $U$ comprises two isolated points on each elliptic curve $E_{i}\,(i=1,\ldots,2s)$. Since $C=\Psi^{2}(C)$ by Claim 3.12, $C$ intersects each $E_{i}$ in one of these two points or in the intersection point with $O_{S}$ or $R_{S}$.
Switching to $\Psi$, the fixed locus becomes
$\mbox{Fix}(\Psi)=\Theta_{3}\cup O_{S}\cup U_{1}\cup U_{2}$
where $U_{1}$ comprises the five remaining intersection points of components of the reducible fibres and $U_{2}=R_{S}\cap(\Theta_{7}\cup F_{1}^{\prime}\cup\ldots\cup F_{2s}^{\prime})$ consists of another $12s+$ 1 points.
Assume that $\Psi(C)=C$. Since $\Psi$ interchanges the points in $U$, the bisection $C$ can intersect each elliptic curve $F_{i}^{\prime}$ only in the intersection point with $O_{S}$ or $R_{S}$. It follows that
(3.5) $C.(O_{S}+R_{S})\geq 2s\geq 2N.$
On the other hand, expressing $R_{S}$, up to torsion, as a linear combination of our given $\mathbb{Q}$-basis of $\mathrm{Num}(S)$,
$R_{S}\sim 2F_{1}^{\prime}-\Theta_{1}-2\Theta_{2}-3\Theta_{3}-2\Theta_{4}-\Theta_{5}-2\Theta_{7}+O_{S},$
(obtained from the non-zero intersection numbers $R_{S}.F_{i}^{\prime}=R_{S}.\Theta_{7}=R_{S}.C_{1}=R_{S}.C_{3}=1$), we derive that $D.R_{S}=N$. As above, it follows that $C.R_{S}\leq N$. Together with $C.O_{S}\leq N-1$, this gives the required contradiction to (3.5). Hence $\Psi(C)\neq C$ as claimed. ∎
In view of the freedom in choosing the locations of the $2s$ multiple fibres, Claims 3.12 and 3.13 complete the proof of Theorem 3.6 for $r=4$.
3.7. Proof of Theorem 3.6 for $r=6$
The proof of Theorem 3.6 is completed by the following result covering the case $r=6$:
Theorem 3.14.
There are infinitely many families of isotrivial surfaces $S$ with group $G=\mathbb{Z}/3\rtimes\mu_{6}$ with $\mathrm{Aut}_{\mathbb{Z}}(S)$ of order $3$.
Proof.
We let $C\rightarrow\mathbb{P}^{1}$ the Galois $G$-covering with monodromies
$(\sigma,\sigma^{-1},\eta_{1},\dots\eta_{n}),\ \sigma(z):=\epsilon z,\ \eta_{j}\in T:=\mathbb{Z}/3\left(\frac{1+\epsilon}{3}\right),\ \sum_{j}\eta_{j}=0.$
$S$ is birational to $Y:=(C\times E)/G$, and the pluricanonical fibration $f:S\rightarrow C/G$ has the first singular fibre of type $\mathrm{II}$ (a cuspidal cubic) and the second one of type $\mathrm{II}^{*}$, that is, of type $\tilde{E_{8}}$.
The other singular fibres are $3F^{\prime}_{1},\dots,3F^{\prime}_{n}$, where each $F^{\prime}_{j}$ is again isomorphic to the Fermat elliptic curve.
The only singular points of $Y$ lie in the first two fibres of the fibration over $C/G$ (with monodromies $\sigma,\sigma^{-1}$) and in the first and third fibres of the projection onto $E/G$, i.e. those with branching indices $6$ and $2$, while there are no singular points over the point of $E/G$ with branching index $3$.
The first fibre of the rational fibration of $S\dasharrow E/G$ consists of the proper transform $\Phi_{0}$ of the first fibre of the fibration $Y\rightarrow E/G$, counted with multplicity $6$, and the last curves of the configurations of type $A_{5}$, respectively $A_{2}$, coming from the resolution of the singular points of $Y$ lying in the second fibre of $S\rightarrow C/G$ and in the first fibre of $S\dasharrow E/G$.
We remove in the configuration of type $A_{5}$ the $(-2)$ curve which intersects $\Phi_{0}$, and we obtain from the $\mathrm{II}^{*}$ fibre a configuration of type $E_{8}$, intersecting $\Phi_{0}$ only on $\Theta$, the last component of $A_{2}$ configuration (appearing with multiplicity $2$ on the fibre).
Now, $S$ has $p_{g}(S)=q(S)=0$, hence the second Betti number $b_{2}(S)=10$, and $\mathrm{Num}(S)$ contains the direct sum $E_{8}\oplus U^{\prime}$, where $U^{\prime}$ is the lattice generated by $F^{\prime}_{1}$ and $\Phi_{0}$, which is unimodular because
$F^{\prime}_{1}\cdot\Phi_{0}=1$
by inspection of their multiplicities in $Y$. Since $E_{8}$ is unimodular, and $F^{\prime}_{1}$ is orthogonal to $E_{8}$, we can subtract from $\Phi_{0}$ the dual vector $\Theta^{\vee}\in E_{8}^{\vee}=E_{8}$ so that $U^{\prime\prime}=\langle F^{\prime}_{1},\Phi_{0}-\Theta^{\vee}\rangle$ is still unimodular and now orthogonal to $E_{8}$. Hence we get an orthogonal direct sum:
$\mathrm{Num}(S)=E_{8}\oplus U^{\prime\prime}.$
Since $\Psi$ acts trivially on $H_{1}(S,\mathbb{Z})$ and also on $F^{\prime}_{1},\Phi_{0}$ and on $E_{8}$, we infer that $\Psi$ is cohomologically trivial.
Remark 3.15.
One could also construct the above surfaces from those considered in the proof of Theorem 3.7 by quotienting by the involution $(\mathrm{id},\epsilon^{3})$ (which is not numerically trivial).
We emphasize that, with the proof of Theorem 3.6 completed, we have also covered Theorem 1.5. In addition, inspecting the proof of the previous Theorem 3.6, we get a small sharpening of Theorem 1.3 (ii).
Theorem 3.16.
If $S$ has $\chi(S)>0$, and the pluricanonical fibration
$f\colon S\rightarrow B$
is isotrivial, then, if the group $G\cong T\rtimes\mu_{r}$, then $|\mathrm{Aut}_{\mathbb{Q}}(S)|\leq r$ for $r\leq 4$ and $\leq 3$ for $r=6$, unless $q(S)=p_{g}(S)=0$ ($B$ is rational and $\chi(S)=1$), in which case we have the estimate $|\mathrm{Aut}_{\mathbb{Q}}(S)|\leq rs\leq r(P_{2}(S)+1)$, where $s$ is the number of multiple fibres of the fibration and the inequality improves to $|\mathrm{Aut}_{\mathbb{Q}}(S)|\leq 3s\leq r(P_{2}(S)+1)$ in case $r=6$.
There are infinitely many examples with $|\mathrm{Aut}_{\mathbb{Q}}(S)|=4s=4(P_{2}(S)+1)$, for each value of $s\in 2\mathbb{N}$.
Proof.
To deduce the first claims from our previous arguments, we use the inequality
(3.6) $|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|\leq s=P_{2}(S)+1$
to be proved in Proposition 6.1 (3) (a). Here, for the equality to the right, we use $K_{S}=-F+\sum_{1}^{s}(m_{j}-1)F^{\prime}_{j}$ by the canonical bundle formula (2.4), hence $2K_{S}=(s-2)F+\sum_{1}^{s}(m_{j}-2)F^{\prime}_{j}$ and $P_{2}(S)=s-1.$
The given bounds for $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ then follow from Theorem 3.6 using (1.2).
In order to prove the last assertion, we construct the following examples, where $s=2m$ is strictly positive, $E$ is the Gaussian elliptic curve,
$G=(\mathbb{Z}/2)\eta\times\mu_{4},\ \eta:=\frac{1+i}{2},$
$\sigma(z):=iz$ is a generator of $\mu_{4}$, $t(z):=z+\eta$ is translation by $\eta$, and $C\rightarrow\mathbb{P}^{1}=C/G$ is the Galois covering branched in $2+2m$ points with local monodromies $(\sigma,\sigma^{-1},t,\dots,t)$.
We assume that the first two branch points are $0,\infty$, while the other $s$ points are $\{1,\zeta,\zeta^{2},\dots,\zeta^{s-1}\}$, where $\zeta\in\mu_{s}$ is a generator.
Since $\xi\mapsto\zeta\xi$ defines an automorphism $\phi$ of $\mathbb{P}^{1}$ which preserves the branch set and the monodromy, $\phi$ lifts to an automorphism $\phi^{\prime}$ of $C$ which centralizes $G$ (see for instance [Cat15], section 6.1), and which therefore has order $s^{\prime}=2^{j}s$ for some $0\leq j\leq 2$. In fact, the ramification indices of the composite $C\rightarrow\mathbb{P}^{1}\rightarrow\mathbb{P}^{1}/\langle\phi\rangle$ over the images of 00 and $\infty$ are both $4s$, and hence $s^{\prime}=4s$.
We let $\mu_{s^{\prime}}\times\mu_{4}$ we have similar actions for other values of $r$, for instance for $r=3,6,$ we let $s=3m$ and take all local monodromies to be equal. be the group of automorphisms of $C\times E$ such that
$\Phi(x,z):=(\phi^{\prime}x,z),\ \ \epsilon(x,z)=(x,iz).$
These automorphisms centralize $\Delta_{G}$, hence they induce a group of automorphisms of $Y=(C\times E)/\Delta_{G}$, and of $S$, with cardinality at least $4s=s^{\prime}=\mathrm{ord}(\Phi)=\mathrm{ord}(\Phi_{S})$. (since no power of $\Phi$ different from the Identity is in $\Delta_{G}$). It suffices to show that both $\Phi_{S}$, $\epsilon_{S}\in\mathrm{Aut}_{\mathbb{Q}}(S)$.
This was already seen for $\epsilon_{S}$, since $\epsilon_{Y}$ acts as the identity on the first two singular fibres of $Y$, hence $\epsilon_{S}$ acts as the identity on the set of components of the first two singular fibres of $S$, of type $\mathrm{III},\mathrm{III}^{*}$ respectively, and preserves the fibres of the second fibration (onto $E/G$). The same argument applies to $\Phi_{S}$, because $\phi$ fixes the first two branch points, hence again $\Phi_{Y}$ acts as the identity on the two first singular fibres of $Y$.
Therefore $\Phi_{S}$, $\epsilon_{S}$ are numerically trivial, and by the given upper bound $|\mathrm{Aut}_{\mathbb{Q}}(S)|\leq rs=4s$, they generate the full group $\mathrm{Aut}_{\mathbb{Q}}(S)$ of order $4s$. Note that actually $\Phi_{S}$ suffices as a generator by inspection of its order, so $\epsilon_{S}=\Phi^{s}$. Since $\Phi_{S}$ permutes the $s$ multiple fibres (of multiplicity $2$) of the pluricanonical fibration cyclically of order $s$, it generates $\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}$ by (3.6). ∎
4. Jacobian elliptic fibrations and their Mordell–Weil groups
Let $f\colon S\rightarrow B$ be a relatively minimal elliptic surface with $\chi(S)>0$ and let $J(S)$ be its Jacobian.
As mentioned in the introduction, the function field extension $\mathbb{C}(S)\supset\mathbb{C}(B)$ provides a curve ${\mathcal{C}}$ of genus $1$ (the fibre of $f$ over the generic point of $B$) defined over the non algebraically closed field ${\mathcal{K}}:=\mathbb{C}(B)$. The subgroup of $\mathrm{Aut}(S)$ of automorphisms preserving the fibration and inducing the identity on $B$ is isomorphic to the subgroup $\mathrm{Aut}_{{\mathcal{K}}}({\mathcal{C}})$ of the automorphisms of ${\mathcal{C}}$ which act as the identity on the base field ${\mathcal{K}}$.
These embed in $\mathrm{Aut}_{\bar{{\mathcal{K}}}}({\mathcal{C}}\otimes\bar{{\mathcal{K}}})$, where ${\mathcal{E}}:={\mathcal{C}}\otimes\bar{{\mathcal{K}}}$ is an elliptic curve over an algebraically closed field of characteristic zero hence
$\mathrm{Aut}_{\bar{{\mathcal{K}}}}({\mathcal{E}})={\mathcal{E}}\rtimes\mu_{r},\quad r\in\{2,4,6\},$
where $\mu_{r}$ is the group of $r$th roots of $1$ in $\mathbb{C}$, and ${\mathcal{E}}$ acts by translations via the isomorphism ${\mathcal{E}}\cong\operatorname{Pic}^{1}({\mathcal{E}})\cong\operatorname{Pic}^{0}({\mathcal{E}})$ provided by the choice of a point $0\in{\mathcal{E}}$ ($\operatorname{Pic}^{d}({\mathcal{E}})$ being as usual the set of linear equivalence classes of divisors of degree $d$).
We denote as usual $\mathrm{Jac}({\mathcal{C}}):=\operatorname{Pic}^{0}({\mathcal{C}})$, and ${\mathcal{C}}$ is a principal homogeneous space over $\mathrm{Jac}({\mathcal{C}})$ ([Shaf65], Chapter 2 of [Dolg81]), in particular, the Mordell–Weil group, the group of ${\mathcal{K}}$-rational points of $\mathrm{Jac}({\mathcal{C}})$ acts on ${\mathcal{C}}$.
In fact, the automorphisms in $\mathrm{Aut}_{{\mathcal{K}}}({\mathcal{C}})$ are the elements of $\mathrm{Aut}_{\bar{{\mathcal{K}}}}({\mathcal{E}})$ which are left fixed by the action of the Galois group of the extension ${\mathcal{K}}\subset\bar{{\mathcal{K}}}$, hence the subgroup which has trivial image to $\mu_{r}$ consists of the Galois -invariant elements in ${\mathcal{E}}$, and these are the ${\mathcal{K}}$ rational points of $\mathrm{Jac}({\mathcal{C}})$.
The geometrical counterpart of $\mathrm{Jac}({\mathcal{C}})$ is the Jacobian surface $J(S)$.
By specializing to the fibre $F_{b}$ over a general point, we see that the condition of having trivial image to $\mu_{r}$ means that the restricted automorphism consists of a translation on $F_{b}$.
Hence we have:
Lemma 4.1.
Let $f\colon S\rightarrow B$ be a relatively minimal elliptic surface with $\chi(S)>0$. Then there is an injective homomorphism
(4.1) $t=t_{S}\colon\mathrm{MW}(J(f))\rightarrow\mathrm{Aut}_{B}(S),$
such that
$\mathrm{im}(t)=\{\sigma\in\mathrm{Aut}_{B}(S)\mid\text{$\sigma|_{F}$ is a translation for a general fibre $F$}\}.$
Lemma 4.2 ([DM22, Theorem 3.3]).
Let $f\colon S\rightarrow B$ be a relatively minimal elliptic surface with $\chi(S)>0$. Then there is a homomorphism
(4.2) $\Phi\colon\mathrm{Aut}_{f}(S)\rightarrow\mathrm{Aut}_{J(f)}(J(S))$
such that $\mathrm{Ker}(\Phi)=t(\mathrm{MW}(J(f)))$ and, for any $\sigma\in\mathrm{Aut}_{f}(S)$, the following holds:
  1. (1)
    Both $\sigma$ and $\Phi(\sigma)$ induce the same automorphism of $B$.
  2. (2)
    $\Phi(\sigma)$ preserves the zero section of $J(f)\colon J(S)\rightarrow B$.
  3. (3)
    If $\sigma\in\mathrm{Aut}_{\mathbb{Q}}(S)$, then $\Phi(\sigma)\in\mathrm{Aut}_{\mathbb{Q}}(J(S))$.
The automorphism $\Phi$ in the previous Lemma is just the natural action of $\mathrm{Aut}({\mathcal{C}})$ on $Jac({\mathcal{C}})=Pic^{0}({\mathcal{C}})$.
Lemma 4.3.
Let $f\colon S\rightarrow B$ be a relatively minimal elliptic surface with $\chi(S)>0$. Suppose that $f$ has a fibre of additive type. Then a nontrivial numerically trivial automorphism of $S$ does not induce fibrewise translations, that is,
$\mathrm{Aut}_{\mathbb{Q}}(S)\cap t(\mathrm{MW}(J(f))_{\mathrm{tor}})=\{\mathrm{id}_{S}\}$
Proof.
Let $f^{*}b$ be a fibre of additive type. Since $\mathrm{MW}(J(S))_{\mathrm{tor}}$ injects into the group of reduced components $A_{f^{*}b}$ by (2.12), it follows that an automorphism $t(P)\in t(\mathrm{MW}(J(S))_{\mathrm{tor}})$ is either trivial (e.g. if there is a fibre of type $\mathrm{II}$ or $\mathrm{II}^{*}$) or it permutes the simple components of $f^{*}b$ nontrivially, cf. Fact 2.3. In the latter case, $t(P)$ is obviously not numerically trivial. Therefore, we obtain the equality of the lemma.
Lemma 4.4.
Let $f\colon S\rightarrow B$ be a relatively minimal elliptic surface with $\chi(S)>0$. If $p_{g}(S)>0$, then we have
$\mathrm{Aut}_{B,{\mathcal{O}}}(S)=\mathrm{im}(t),\quad\mathrm{Aut}_{B,\mathbb{Q}}(S)\subset t(\mathrm{MW}(J(f))_{\mathrm{tor}}).$
Proof.
By [DL23, Lemma 2.9], $t(P)$ acts trivially on $H^{0}(S,K_{S})=H^{2}(S,{\mathcal{O}}_{S})^{\vee}$, and hence also trivially on the transcendental part of $H^{2}(S,\mathbb{Q})$.§§§One may also argue like this: The quotient map $S\rightarrow S/t_{P}|_{S}$ is an isogeny of elliptic surfaces, so $p_{g}(S)=p_{g}(S/t_{P}|_{S})$ and hence $t_{P}|_{S}$ acts trivially on $H^{0}(S,K_{S})$.
Since $\chi(S)>0$, the pullback $f^{*}\colon H^{1}(B,{\mathcal{O}}_{B})\rightarrow H^{1}(S,{\mathcal{O}}_{S})$ is an isomorphism. Since $t(P)\in\mathrm{Aut}_{B}(S)$ induces the trivial action on $B$, one sees that $t(P)$ also acts trivially on $H^{1}(S,{\mathcal{O}}_{S})$. It follows that $t(P)\in\mathrm{Aut}_{B,{\mathcal{O}}}(S)$ and hence $\mathrm{im}(t)\subset\mathrm{Aut}_{B,{\mathcal{O}}}(S)$.
For the other inclusion, take any $\sigma\in\mathrm{Aut}_{B,{\mathcal{O}}}(S)$. Assuming that $\sigma$ is not a translation on a general fibre $F$, it necessarily fixes some points on $F$ and is thus of finite order. Hence we can argue with the quotient surface $S/\langle\sigma\rangle$, which is a $\mathbb{P}^{1}$-fibration over $B$ whence $p_{g}(S/\langle\sigma\rangle)$ vanishes. On the other hand, by the choice of $\sigma$, we have $p_{g}(S/\langle\sigma\rangle)=p_{g}(S)>0$. This gives the required contradiction.
Therefore, $\sigma$ induces a translation on the general fibres of $f$. By Lemma 4.1, $\sigma$ comes from $\mathrm{MW}(J(S))$.
Finally, since $\mathrm{Aut}_{B,\mathbb{Q}}(S)$ is a finite subgroup of $\mathrm{Aut}_{B,{\mathcal{O}}}(S)$ and $t$ is injective, $\mathrm{Aut}_{B,\mathbb{Q}}(S)$ lies in the image of the torsion subgroup $\mathrm{MW}(J(f))_{\mathrm{tor}}$ of $\mathrm{MW}(J(f))$. ∎
Combining Lemmas 4.3 and 4.4, we obtain
Corollary 4.5.
Let $f\colon S\rightarrow B$ be a relatively minimal elliptic surface with $\chi(S)>0$. If $p_{g}(S)>0$ and if $f$ admits a fibre of additive type, then $\mathrm{Aut}_{B,\mathbb{Q}}(S)$ is trivial.
We conclude this section by recording the following consequences for jacobian fibrations.
Lemma 4.6.
Let $h\colon X\rightarrow B$ be a relatively minimal elliptic surface admitting a section and with $\chi(X)>0$. If $\mathrm{Aut}_{B,\mathbb{Q}}(X)$ is nontrivial, then
  1. (1)
    $\mathrm{MW}(X)$ is a finite group, that is, $\mathrm{MW}(X)=\mathrm{MW}(X)_{\mathrm{tor}}$;
  2. (2)
    $p_{g}(X)=0$, i.e. $X$ is rational.
Proof.
(1) Note that the numerical classes of two distinct sections of $h$ are different, since $P^{2}=-\chi(X)<0$ for any section $P$.
It follows that $\mathrm{Aut}_{B,\mathbb{Q}}(X)$ fixes (pointwise) each section of $h$. Since $\mathrm{Aut}_{B,\mathbb{Q}}(X)$ is a nontrivial group, the Mordell–Weil group $\mathrm{MW}(X)$ of sections is a finite group.
(2) If $p_{g}(X)>0$, then $\mathrm{Aut}_{\mathbb{Q}}(X)\subseteq t(\mathrm{MW}(X)_{\mathrm{tor}})=t(\mathrm{MW}(X))$ by Lemma 4.4 and (1). But then any torsion section $P$ meets some fibre in a non-identity component; this is a consequence of (2.10). It follows that $t(P)$ acts as a non-trivial permutation on the (simple) fibre components, hence $t(P)\not\in\mathrm{Aut}_{\mathbb{Q}}(X)$.
5. Bounding the torsion part of the Mordell–Weil groups
In view of Lemmas 4.4 and 4.6, it is crucial to bound $\mathrm{MW}(J(f))_{\mathrm{tor}}$, if one wants to bound $\mathrm{Aut}_{B,\mathbb{Q}}(S)$ and $\mathrm{Aut}_{\mathbb{Q}}(S)$. In this respect, the elliptic modular surfaces, introduced in [Shi72], play a key role, since $\mathrm{MW}(J(f))_{\mathrm{tor}}$ is essentially pulled back from the Mordell–Weil group of an elliptic modular surface.
Note that any abelian group generated by two elements can be realized as a subgroup of the Mordell–Weil group of an elliptic modular surface by [Shi72, Example 5.4]. The aim of this subsection is to give a linear bound of $|\mathrm{MW}(X)_{\mathrm{tor}}|$ for a jacobian elliptic surface $h\colon X\rightarrow B$ in terms of $g(B)$; see Proposition 5.5.
We recall, for each positive integer $N$, the following congruence modular groups ([Miy06, Section 4.2]):
(5.1) $\begin{split}&\Gamma(N)=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{SL}_{2}(\mathbb{Z})\,\Bigg{|}\,b\equiv c\equiv 0,a\equiv d\equiv 1\mod N\right\}.\\ &\Gamma_{1}(N)=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{SL}_{2}(\mathbb{Z})\,\Bigg{|}\,c\equiv 0,a\equiv d\equiv 1\mod N\right\}.\end{split}$
Let $B(N)$ and $B_{1}(N)$ be the modular curves associated to $\Gamma(N)$ and $\Gamma_{1}(N)$ respectively, and let $g(N)$ and $g_{1}(N)$ be their respective genera. For $N\geq 5$, we have, by [Shi72, (5.3)] and [Mar05, page 314, line 2],
(5.2) $\begin{split}&g(N)=\frac{1}{24}s(N)N^{3}-\frac{1}{4}s(N)N^{2}+1,\\ &g_{1}(N)=\frac{1}{24}s(N)N^{2}-\frac{1}{4}u(N)+1.\end{split}$
where $s,u\colon\mathbb{N}\rightarrow\mathbb{Z}$ are the multiplicative functions satisfying, for a prime number $p$ and an integer $n\geq 1$,
(5.3) $\begin{split}&s(p^{n})=1-\frac{1}{p^{2}},\\ &u(p^{n})=p^{n-2}(p-1)((n+1)p-n+1).\end{split}$
The formula (5.2) gives $g_{1}(N)=0$ for $5\leq N\leq 10$, and we may continue computing $g_{1}(N)$ for $11\leq N\leq 25$ as follows:
$N$ 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25
$g_{1}(N)$ 1 0 2 1 1 2 5 2 7 3 5 6 12 5 12
Based on this, we go on to bound $g_{1}(N)$ from below in terms of $N$:
Proposition 5.1.
Let $N\geq 5$ be an integer. Then
(5.4) $g_{1}(N)>\frac{N^{2}}{12\pi^{2}}-2.$
Later on, we will give in Corollary 5.4 a better lower bound when $N$ is sufficiently large.
Proof.
The inequality holds for $N\leq 25$ by looking at the above table. In the following, we may assume that $N\geq 26$. We rewrite (5.2) as
(5.5) $g_{1}(N)=\frac{1}{24}s(N)N^{2}\left(1-\frac{6u(N)}{s(N)N^{2}}\right)+1$
By the definition of the multiplicative functions $s$ and $u$ in (5.3), we have
$s(N)=\prod_{p\mid N}\left(1-\frac{1}{p^{2}}\right),\quad u(N)=\prod_{p^{n}||N}\left((n+1)p^{n}-2np^{n-1}+(n-1)p^{n-2}\right)$
where the product ranges over primes $p$ dividing $N$, and $p^{n}||N$ means that $p^{n}$ is the largest power of $p$ dividing $N$.
Since the infinite product $s(\infty):=\prod_{p\text{ prime}}(1-\frac{1}{p^{2}})$ converges to
$\left(\sum_{n\in\mathbb{N}}\frac{1}{n^{2}}\right)^{-1}=\frac{6}{\pi^{2}},$
we have
(5.6) $s(N)>s(N)\cdot\prod_{p{\nmid}N}\left(1-\frac{1}{p^{2}}\right)=s(\infty)=\frac{6}{\pi^{2}}.$
Plugging the lower bound (5.6) of $s(N)$ into the expression (5.5), we obtain
$g_{1}(N)\geq\frac{N^{2}}{4\pi^{2}}\left(1-\frac{6u(N)}{s(N)N^{2}}\right)+1.$
Thus, to show the inequality (5.4) of the proposition, it suffices to show the following claim. ∎
Claim 5.2.
For $N\geq 26$, we have
(5.7) $1-\frac{6u(N)}{s(N)N^{2}}\geq\frac{1}{3},\quad\text{or equivalently,}\quad t(N):=\frac{u(N)}{s(N)N^{2}}\leq\frac{1}{9}.$
Proof of the claim..
Note that the function $t(N)$ is multiplicative. We need to study the values $t(p^{n})$ at the prime powers $p^{n}$. By the definition of the functions $s(N)$ and $u(N)$ in (5.3), we have for a prime $p$ and an integer $n\geq 1$:
(5.8) $t(p^{n})=\frac{p^{n-2}\left((n+1)p^{2}-2np+n-1\right)}{\left(1-\frac{1}{p^{2}}\right)p^{2n}}=\frac{(n+1)p-(n-1)}{p^{n}(p+1)}$
Replacing $p$ and $n$ in (5.8) with two real variables $x,y\in\mathbb{R}_{\geq 2}$, and analyzing the function $t(x^{y})$, one sees that $t(p^{n})$ is a decreasing function of $p$ and $n$. For the small $p$ and $n$, we compute directly the value $t(p^{n})$ as follows:
$n$ $p$ $2$ 3 5 7 11 13 17 19 23
$1$ 2/3 $1/2$ $1/3$ $1/4$ $1/6$ $1/7$ $1/9$ $1/10$ $1/12$
$2$ $5/12$ $2/9$ $7/75$ $5/98$ $4/363$ 19/1183
$3$ $1/4$ $5/54$
$4$ $7/48$
$5$ $1/12$
We stop the computation along a column as soon as $t(p^{n})\leq\frac{1}{9}$, which is then sufficient for the proof of Claim 5.2.
Now suppose that $N\geq 26$, and it has a prime decomposition
$N=p_{1}^{n_{1}}\cdot\dots\cdot p_{r}^{n_{r}}.$
We proceed according to the value of $r$. If $r\geq 3$, then, by the mononicity of $t(p^{n})$ for $p$ and $n$, we have
$t(N)\geq t(2\cdot 3\cdot 5)=t(2)\cdot t(3)\cdot t(5)=\frac{2}{3}\cdot\frac{1}{2}\cdot\frac{1}{3}=\frac{1}{9}.$
If $r=2$, then, by the mononicity of $t(p^{n})$ for $p$ and $n$, we may assume that $p_{i}^{n_{i}}<26$ for both $i=1,2$, and the value of $t(N)$ is at most one of the following:
$t(2^{4}\cdot 3),\,t(2^{2}\cdot 3^{2}),\,t(2^{3}\cdot 5),\,t(2\cdot 5^{2}),\,t(2^{2}\cdot 7),\,t(2^{2}\cdot 11),\,t(2\cdot 13),$
$t(3^{2}\cdot 5),\,t(3\cdot 5^{2}),\,t(3^{2}\cdot 7),\,t(3\cdot 11),\,t(5\cdot 7)$
all of which can be directly checked to be less than $\frac{1}{9}$. For example,
$t(2^{4}\cdot 3)=t(2^{4})\cdot t(3)=\frac{7}{48}\cdot\frac{1}{2}=\frac{7}{96}<\frac{1}{9}.$
If $r=1$, then $N=p^{n}$ for some prime $p$, and by the mononicity of $t(p^{n})$ for $p$ and $n$ again, we have
$t(N)\leq\min\{t(2^{5}),\,t(3^{3}),\,t(7^{2}),\,t(29)\}=\min\left\{\frac{1}{12},\,\frac{5}{54},\,\frac{5}{98},\,\frac{1}{15}\right\}=\frac{5}{98}<\frac{1}{9}.$
This finishes the proof of the claim. ∎
As a side remark, we observe an asymptotic behavior of the ratio $g_{1}(N)/N^{2}$, which seems to be known ([Mar05, Theorem 7]), but not in the following form:
Lemma 5.3.
We have
[…]
Proof.
As is known (see for example [Mar05, Lemma 30]), we have
$\lim_{d\to\infty}\frac{u(N)}{d^{2}}=0,$
and hence by (5.2), $\frac{g_{1}(N)}{N^{2}}$ and $\frac{s(N)}{24}$ share the same upper and lower limits.
To compute the upper limit, we note that $s(N)<1$ in any case and, taking $d$ to be primes $p$, we have
$\lim_{p\to\infty}\frac{s(p)}{24}=\frac{1}{24}\lim_{p\to\infty}\left(1-\frac{1}{p^{2}}\right)=\frac{1}{24}.$
For the lower limit, we note that $s(N)>s(\infty)=\frac{6}{\pi^{2}}$ as in (5.6). On the other hand, taking the sequence of prime numbers $\{p_{r}\}_{r}$ and setting $N_{r}:=\prod_{1\leq i\leq r}p_{i}$, we obtain
$\lim_{r\to\infty}\frac{s(N_{r})}{24}=\frac{1}{24}\lim_{r\to\infty}\prod_{1\leq i\leq r}\left(1-\frac{1}{p_{i}^{2}}\right)=\frac{1}{24}\cdot\frac{6}{\pi^{2}}=\frac{1}{4\pi^{2}}.$
Corollary 5.4.
For any positive real number $\epsilon$, there is a positive integer $N_{\epsilon}$, depending on $\epsilon$, such that, for any $N\geq N_{\epsilon}$, we have
$g_{1}(N)\geq\frac{N^{2}}{4\pi^{2}+\epsilon}.$
Now we may use Proposition 5.1 to bound $|\mathrm{MW}(X)|$ for jacobian elliptic surfaces.
Proposition 5.5.
Let $h\colon X\rightarrow B$ be a jacobian elliptic surface with $\chi({\mathcal{O}}_{X})>0$. Let $g=g(B)$ be the genus of the base curve. Then
(5.10) $|\mathrm{MW}(X)_{\mathrm{tor}}|<12\pi^{2}(g+2).$
Moreover, if the fibration is isotrivial, then there is the uniform bound $|\mathrm{MW}(X)_{\mathrm{tor}}|\leq 4$.
Proof.
First suppose that $h$ is not isotrivial. Let $P\in\mathrm{MW}(X)_{\mathrm{tor}}$ be a torsion section of maximal order $N$. Then $|\mathrm{MW}(X)_{\mathrm{tor}}|\leq N^{2}$. Let $h_{N}\colon X_{1}(N)\rightarrow B_{1}(N)$ be the modular elliptic surface parametrizing elliptic curves with a $N$-torsion point $P_{N}$ for $N\geq 4$ (which thus features as a section of $X_{1}(N)$). Then there is a surjective morphism $\pi\colon B\rightarrow B_{1}(N)$ such that $h\colon X\rightarrow B$ and $P$ is obtained by pulling back $h_{N}$ and $P_{N}$ via the base change $\pi$. Since $N^{2}\leq 12\pi^{2}(g_{1}(N)+2)$ by Proposition 5.1 and $g=g(B)\geq g(B_{1}(N))=g_{1}(N)$, we obtain the claimed inequality (5.10).
If $h$ is isotrivial, then there are no $\mathrm{I}_{n}$ fibres with $n\geq 1$, and the order of $\mathrm{MW}(X)_{\mathrm{tor}}$ is at most $4$ by (2.12) (cf. [SS19, 5.6.1]), which is of course smaller than $12\pi^{2}(g+2)$. ∎
Remark 5.6.
The bound $|\mathrm{MW}_{\mathrm{tor}}|\leq N^{2}$ in the proof of Proposition 5.5 is not optimal; for instance, if $|\mathrm{MW}_{\mathrm{tor}}|=N^{2}$, then the genus $g$ of the base curve $B$ should rather be bounded from below by $g(N)$. We omit the details for brevity.
A computation of the genera of the modular curves for all intermediate groups between $\mathbb{Z}/N\mathbb{Z}$ and $(\mathbb{Z}/N\mathbb{Z})^{2}$ such that the obvious condition $g_{1}(N)\leq 4$ holds, e.g. using Magma [BCP97], yields the following sharp upper bounds for $|\mathrm{MW}_{\mathrm{tor}}|$ in terms of the genus $g$:
$\begin{array}[]{c||c|c|c|c|c}g&0&1&2&3&4\\ \hline\cr|\mathrm{MW}_{\mathrm{tor}}|\leq&25&36&36&49&50\end{array}$
6. General upper bounds for $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ and $|\mathrm{Aut}_{\mathbb{Z}}(S)|$
Let $f\colon S\rightarrow B$ be a minimal properly elliptic surface with $\chi(S)>0$. Our aim in this section is to give an upper bound for $|\mathrm{Aut}_{\mathbb{Q}}(S)|$ in terms of the irregularity $q(S)$. If the relative Jacobian $J(S)$ is not rational, see Corollary 6.13. If $J(S)$ is rational, or, equivalently, $p_{g}(S)=0$, the situation is quite different, and we treat the latter case first.
6.1. The case where $J(S)$ is rational
Proposition 6.1.
Let $f\colon S\rightarrow B$ be a minimal properly elliptic surface such that $J(f)\colon J(S)\rightarrow B$ is a rational elliptic surface. Then the following holds.
  1. (1)
    If $f$ does not admit any singular fibre of additive type, or, equivalently, if $J(f)$ is semistable, then we have
    $|\mathrm{Aut}_{\mathbb{Q}}(S)|=|\mathrm{Aut}_{B,\mathbb{Q}}(S)|\leq 9$
    and the upper bound can be attained. Moreover, in the equality case, $J(S)$ must be the extremal rational elliptic surface $X_{3333}$ in the list of [MP86], [Mir89, VIII.1.4].
  2. (2)
    If $f$ admits a singular fibre of additive type, then we have
    $|\mathrm{Aut}_{B,\mathbb{Q}}(S)|\leq 2,\quad|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|\leq 3,\quad|\mathrm{Aut}_{\mathbb{Q}}(S)|\leq 4.$
    unless $J(S)$ is one of the surfaces $X_{22}$, $X_{33}$, $X_{44}$, $X_{11}(j)$ in [MP86], [Mir89, VIII.1.4].
  3. (3)
    If $J(S)$ is one of the isotrivial elliptic surfaces $X_{22},X_{33}$, $X_{44}$, $X_{11}(j)$, then the following holds.
    1. (a)
      $\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}$ is finite cyclic, and its order divides $P_{2}(S)+1$, where the $P_{2}(S):=h^{0}(2K_{S})$ is called the second plurigenus of $S$;
    2. (b)
      $\mathrm{Aut}_{B,\mathbb{Q}}(S)$ is a subgroup of $\mathrm{Aut}_{B,\mathbb{Q}}(J(S))=\mu_{r}$ with $r=6,4,3,2$ in the given order of the isotrivial surfaces; moreover, it is a proper subgroup if $r=6$.
    3. (c)
      Conversely, for $X=X_{22},$ $X_{33}$, $X_{44}$, $X_{11}(j)$, given $s\in\mathbb{N}$, there is a minimal properly elliptic surface $S$ such that $J(S)=X$ and $|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|=P_{2}(S)+1=s$.
Observe that part (3) gives a different (not completely self-contained) proof of some assertions in Theorem 3.16. Explicit equations of $X_{33},X_{44}$, and $X_{11}(j)$ can be found in Example 9.3.
Proof of Proposition 6.1.
Consider the homomorphism
$\Phi\colon\mathrm{Aut}_{f}(S)\rightarrow\mathrm{Aut}_{J(f)}(J(S))$
of Lemma 4.2. Since $S$ has Kodaira dimension $1$, and $f$ is the pluricanonical fibration, while, for $J(S)$, $J(f)$ is the plurianticanonical fibration in view of the canonical bundle formula (2.4), and of the fact that there are no multiple fibres, we have
$\mathrm{Aut}_{f}(S)=\mathrm{Aut}(S),\quad\mathrm{Aut}_{J(f)}(J(S))=\mathrm{Aut}(J(S))$
hence $\Phi$ is indeed a homomorphism between the full automorphism groups. Restricting $\Phi$ to $\mathrm{Aut}_{\mathbb{Q}}(S)$, we obtain, by Lemma 4.2, the exact sequences
(6.1) $\begin{split}&1\rightarrow t(\mathrm{MW}(J(S)))\cap\mathrm{Aut}_{\mathbb{Q}}(S)\rightarrow\mathrm{Aut}_{\mathbb{Q}}(S)\rightarrow\mathrm{Aut}_{\mathbb{Q}}(J(S)),\\ &1\rightarrow t(\mathrm{MW}(J(S)))\cap\mathrm{Aut}_{B,\mathbb{Q}}(S)\rightarrow\mathrm{Aut}_{B,\mathbb{Q}}(S)\rightarrow\mathrm{Aut}_{B,\mathbb{Q}}(J(S)),\\ &1\rightarrow\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}\rightarrow\mathrm{Aut}_{\mathbb{Q}}(J(S))|_{B}.\end{split}$
(1) We have $|\mathrm{MW}(J(S))_{\mathrm{tor}}|\leq 9$ in any case ([SS19, Corollary 8.21]), and equality holds exactly if $J(S)=X_{3333}$. If $J(f)$ is semi-stable, then it is not in [DM22, Table 3] and hence $\mathrm{Aut}_{\mathbb{Q}}(J(f))$ is trivial. By the first exact sequence of (6.1), we obtain
$\mathrm{Aut}_{\mathbb{Q}}(S)\subset t(\mathrm{MW}(J(S))),$
and (1) follows.
(2) If $f$ admits an additive fibre, then by Lemma 4.3, the homomorphism $\mathrm{Aut}_{\mathbb{Q}}(S)\rightarrow\mathrm{Aut}_{\mathbb{Q}}(J(S))$ of Lemma 4.2 is injective, and same is true for the restrictions to $\mathrm{Aut}_{B,\mathbb{Q}}$ and $\mathrm{Aut}_{\mathbb{Q}}|_{B}$.
If $J(S)$ is not one of the surfaces $X_{22}$, $X_{33}$, $X_{44}$, $X_{11}(j)$, then by [DM22, Theorem 6.4], we have
$|\mathrm{Aut}_{B,\mathbb{Q}}(J(S))|\leq 2,\quad|\mathrm{Aut}_{\mathbb{Q}}(J(S))|_{B}|\leq 3,\quad|\mathrm{Aut}_{\mathbb{Q}}(J(S))|\leq 4.$
Now (2) follows from these since by injectivity $|\mathrm{Aut}_{\mathbb{Q}}(S)|\leq|\mathrm{Aut}_{\mathbb{Q}}(J(S))|$ etc.
(3) If $J(S)$ is one of the surfaces $X_{22},X_{33}$, $X_{44}$, $X_{11}(j)$, the pair of fibres of $f$ with singular support, i.e. with $F_{red}$ singular, which we may assume to be $f^{*}0$ and $f^{*}\infty$, are preserved by $\mathrm{Aut}_{\mathbb{Q}}(S)$. It follows that $\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}$ fixes the two points 00 and $\infty$ of $B=\mathbb{P}^{1}$, and hence is finite cyclic. Since $S$ is properly elliptic, there exist multiple fibres, say $f^{*}b_{i}=m_{i}F^{\prime}_{i}$ with $1\leq i\leq s$. Then $\{b_{i}\mid 1\leq i\leq s\}$ consists of free orbits of the cyclic group $\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}$, and it follows that the order of $\mathrm{Aut}_{\mathbb{Q}}(S)$ divides $s$.
On the other hand, by the canonical bundle formula (2.4), we have
(6.2) $K_{S}=f^{*}(K_{B}+L)+\sum_{1\leq i\leq s}(m_{i}-1)F^{\prime}_{i}$
where $L$ is a divisor on $B$ with $\deg L=\chi(S)=1$. Therefore,
$P_{2}(S)$ $=$ $h^{0}\left(S,f^{*}(2K_{B}+2L)+\sum_{1\leq i\leq s}\underbrace{(2m_{i}-2)F^{\prime}_{i}}_{=f^{*}b_{i}+(m_{i}-2)F^{\prime}_{i}}\right)$
$=$ $h^{0}\left(S,f^{*}\left(2K_{B}+2L+\sum_{1\leq i\leq s}b_{i}\right)\right)$
$=$ $h^{0}\left(B,2(K_{B}+L)+\sum_{i=1}^{s}b_{i}\right)=s-1.$
It follows that the order of $\mathrm{Aut}_{\mathbb{Q}}(S)$ divides $s=P_{2}(S)+1$ and (3a) is proved. (This also provides the remaining input for (3.6) in the proof of Theorem 3.16, implying Theorem 1.3 (ii).)
(3b) Since $\mathrm{Aut}_{B,\mathbb{Q}}(S)\cap t(\mathrm{MW}(J(S)))$ is trivial by Lemma 4.3, the second exact sequence of (6.1) shows that $\mathrm{Aut}_{B,\mathbb{Q}}(S)\rightarrow\mathrm{Aut}_{B,\mathbb{Q}}(J(S))$ is an injection. But then, by [DM22] (or by direct computation), $\mathrm{Aut}_{B,\mathbb{Q}}(J(S))=\mu_{r}$ where $r=6,4,3,2$ depends on $J(S)=X_{22},X_{33},X_{44},X_{11}(j)$ respectively.
Assume that $r=6$. Then $J(S)=X_{22}$. By the discussion of Section 3, there is a Galois base change $C\rightarrow B$ with Galois group $G$ such that
  • $G$ also acts on $E$ faithfully so that the image $G_{E}$ of the action homomorphism $G\rightarrow\mathrm{Aut}(E)$ is of the form $T\rtimes\mu_{6}$, where $T\subset G_{E}$ consists of translations;
  • $S$ is birationally equivalent to $(C\times E)/\Delta_{G}$ over $B=C/G$, where $\Delta_{G}\subset G\times G$ is the diagonal subgroup acting on $C\times E$;
  • $\mathrm{Aut}(S)=N_{\Delta_{G}}/\Delta_{G}$ by (3.1), where $N_{\Delta_{G}}$ is the normalizer of $\Delta_{G}$ in $\mathrm{Aut}(C\times E)$.
By Remark 3.2, we may lift any $\sigma\in\mathrm{Aut}_{B,\mathbb{Q}}(S)$ to $\Psi\in N_{\Delta_{G}}$ such that
$\Psi(x,\,z)=(x,\,\sigma_{E}(z))$
where $\sigma_{E}(z)=\lambda z+c$ for some $\lambda\in\mu_{6}$ and $c\in E$. Since $\Psi\in N_{\Delta_{G}}$, that is, $\Psi\Delta_{G}\Psi^{-1}=\Delta_{G}$, we infer that $\sigma_{E}$ lies in the centre of $G_{E}$. Since $S$ has Kodaira dimension 1 and $J(S)$ is rational, there must be multiple fibres of $f$. This implies that the translation subgroup $T$ of $G_{E}$ is nontrivial. Since $\sigma_{E}$ commutes with the elements of $T$, which are of the form $z\mapsto z+t$, it must hold $\lambda t=t$. It is then clear that $\lambda$ cannot generate $\mu_{6}$, since a generator of $\mu_{6}$ does not fix any non-neutral point of $E$. It follows that the image of $\mathrm{Aut}_{B,\mathbb{Q}}(S)\rightarrow\mathrm{Aut}_{B,\mathbb{Q}}(J(S))=\mu_{6}$ is a proper subgroup of $\mu_{6}$, and hence has order at most $3$.
(3c) has been proved for $X_{33}$ (i.e. $r=4$) in the proof of Theorem 3.16, with actual equality $|\mathrm{Aut}_{\mathbb{Q}}(S)|=rs$. Without requiring the respective equality, but only the equality $|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|=P_{2}(S)+1=s$, examples for $X_{22},X_{44}$ and $X_{11}(j)$ are easily exhibited along the same lines, compare the footnote in the proof of Theorem 3.16; see also Section 9 for an alternative construction.
Thus the proof of Proposition 6.1 is complete. ∎
6.2. Cohomologically trivial automorphisms
Next we study the cohomologically trivial automorphism group $\mathrm{Aut}_{\mathbb{Z}}(S)$, mainly in the case where $J(S)$ is rational. Then the bounds from Proposition 6.1 (1) carry over directly, so we will be mainly concerned with the isotrivial cases from Proposition 6.1 (2), but we keep the arguments as general as possible.
Lemma 6.2.
Let $f\colon S\rightarrow B$ be a relatively minimal elliptic surface with $\chi(S)>0$. If $g(B)=0$, then
  1. (1)
    $\mathrm{Aut}_{\mathbb{Z}}(S)$ preserves each multiple fibre of $f$.
  2. (2)
    $\mathrm{Aut}_{\mathbb{Z}}(S)|_{B}=\{\mathrm{id}_{B}\}$ unless $f$ has at most 2 multiple or reducible fibres in total.
Proof.
(1) this was already observed in section 2.3, and follows from [CFGLS24, Lemma 2.1].
(2) Let $\sigma\in\mathrm{Aut}_{\mathbb{Z}}(S)$ and $\sigma|_{B}\in\mathrm{Aut}_{\mathbb{Z}}(S)|_{B}$ its induced automorphism of $B$, that is, $f\circ\sigma=\sigma|_{B}\circ f$. Let $f^{*}{b_{i}}$, $1\leq i\leq s$ be the collection of all the multiple fibres together with the reducible fibres. Since $\sigma$ preserves each $f^{*}{b_{i}}$, the automorphism $\sigma|_{B}$ fixes each $b_{i}$ for $1\leq i\leq s$. If $s\geq 3$, then necessarily $\sigma|_{B}=\mathrm{id}_{B}$ since $B\cong\mathbb{P}^{1}$. ∎
Corollary 6.3.
If $f\colon S\rightarrow B$ is a minimal properly elliptic surface with $\chi(S)>0$ such that $J(S)$ is rational.
Then $\mathrm{Aut}_{\mathbb{Z}}(S)|_{B}$ is trivial.
Proof.
Suppose on the contrary that $\mathrm{Aut}_{\mathbb{Z}}(S)|_{B}$ is nontrivial. Then so is $\mathrm{Aut}_{\mathbb{Z}}(J(S))|_{B}$, and $J(S)$ lands in [DM22, Table 3]. Since $S$ has Kodaira dimension 1 while $J(S)$ is rational, there exist at least two multiple fibres of $f$. Then, by inspecting [DM22, Table 3], the total number of reducible fibres and multiple fibres is at least 3. It follows that $\mathrm{Aut}_{\mathbb{Z}}(S)|_{B}$ is trivial by Lemma 6.2.
Remark 6.4.
If $J(S)$ is not rational, then $\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}$ and hence also $\mathrm{Aut}_{\mathbb{Z}}(S)|_{B}$ is trivial by Proposition 6.7.
Corollary 6.5.
Let $f\colon S\rightarrow B$ be a minimal properly elliptic surface such that $J(S)$ is a rational elliptic surface. Then the following holds.
  1. (1)
    $|\mathrm{Aut}_{\mathbb{Z}}(S)|=|\mathrm{Aut}_{B,\mathbb{Z}}(S)|\leq 9$, and equality may hold only if $J(S)=X_{3333}$.
  2. (2)
    If $J(f)$ is not semistable (in particular, if $f$ is isotrivial), then $|\mathrm{Aut}_{\mathbb{Z}}(S)|\leq 4$, and equality may hold only if $J(S)=X_{33}$.
Proof.
(1) The first equality follows from Corollary 6.3 combined with (1.2) restricted to $\mathrm{Aut}_{\mathbb{Z}}(S)$, and the inequality together with the characterization of the case of potential equality follows from Proposition 6.1.
(2) If $J(f)$ is not semistable, then
$|\mathrm{Aut}_{B,\mathbb{Z}}(S)|\leq|\mathrm{Aut}_{B,\mathbb{Q}}(S)|\leq 4$
where the second inequality and the characterization of the case of potential equality is again by Proposition 6.1 (2) and (3b). ∎
Remark 6.6.
We don’t know whether the bounds in Corollary 6.5 are sharp.
Sections 10 and 11 exhibit non-isotrivial properly elliptic surfaces with $|\mathrm{Aut}_{\mathbb{Z}}(S)|=2$ and $3$.
6.3. The case where $J(S)$ is not rational
Proposition 6.7.
Let $f\colon S\rightarrow B$ be a minimal properly elliptic surface with $\chi(S)>0$ such that $J(S)$ is not rational. Then the group $\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}$ is trivial.
Proof.
We can divide the discussion into the following three cases:
  1. (1)
    $g(B)\geq 1$;
  2. (2)
    $\chi(S)\geq 3$;
  3. (3)
    $g(B)=0$ and $\chi(S)=2$.
Take $\sigma\in\mathrm{Aut}_{\mathbb{Q}}(S)$, and let $\sigma|_{B}$ denote its image in $\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}$. We need to show that $\sigma|_{B}=\mathrm{id}_{B}$ if one of the above conditions holds.
(1) First suppose that $g(B)\geq 1$. We have $f^{*}\colon H^{1}(B,\mathbb{Q})\hookrightarrow H^{1}(S,\mathbb{Q})$, and the inclusion is compatible with the actions of $\mathrm{Aut}_{\mathbb{Q}}(S)$ and $\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}$ on the source and the target respectively. Since $\sigma$ induces the trivial action on $H^{1}(S,\mathbb{Q})$, $\sigma|_{B}$ acts trivially on $H^{1}(B,\mathbb{Q})$. If $g(B)\geq 2$, it is well-known that the numerically trivial $\sigma|_{B}$ should be $\mathrm{id}_{B}$. If $g(B)=1$ then $B$ is an elliptic curve and $\sigma|_{B}$ acts as a translation on $B$. Since $e(S)=12\chi(S)>0$, by the topological Lefschetz fixed point theorem, $S^{\sigma}\neq\emptyset$. It follows that
$\emptyset\neq f(S^{\sigma})\subset B^{\sigma|_{B}}$
and $\sigma|_{B}$ must be the identity.
(2) Now assume that $\chi(S)\geq 3$. By the canonical bundle formula (2.4), we have
$|K_{S}|=f^{*}|K_{B}+L|+\sum_{1\leq i\leq r}(m_{i}-1)F_{i}$
where $L$ is a divisor on $B$ with $\deg L=\chi(S)$. Thus the canonical map $\varphi_{S}$ of $S$ factors through $B$ as follows:
[…]
where $p_{g}=h^{0}(S,K_{S})$ is the geometric genus of $S$ and $\varphi_{B}$ is the morphism associated to the linear system $|K_{B}+L|$. Since $\deg L=\chi(S)\geq 3$, the morphism $\varphi_{K_{B}+L}$ is an embedding.
Then we have a commutative diagram
[…]
where the last vertical equality is because $\sigma$ acts trivially on $H^{0}(S,K_{S})$, which is a direct summand of $H^{2}(S,\mathbb{C})$ in the Hodge decomposition. It follows that $\sigma|_{B}=\mathrm{id}_{B}$, and $\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}=\{\mathrm{id}_{B}\}$.
(3) Finally, assume that $g(B)=0$ and $\chi(S)=2$. Then the relative Jacobian $J(S)$ over $B$ is a K3 surface. By the proof of the Torelli theorem for K3 surfaces ([PS72], [BR75]) one knows that $\mathrm{Aut}_{\mathbb{Q}}(J(S))$ is trivial. Via the homomorphism (4.2), one sees that
$\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}=\mathrm{Aut}_{\mathbb{Q}}(J(S))|_{B}=\{\mathrm{id}_{B}\}.$
Remark 6.8.
Note that $\mathrm{Aut}_{\mathbb{Q}}(S)$ may be non-trivial, even if $J(S)$ is a K3 surface. This will follow from Construction 7.5 applied to suitable elliptic K3 surfaces admitting a torsion section (as in the proof of Theorem 1.1 (ii) at the end of Subsection 8.2).
Corollary 6.9.
Let $f\colon S\rightarrow B$ be a minimal properly elliptic surface with $\chi(S)>0$. If $p_{g}(S)>0$ and if $f$ admits a fibre of additive type or a section, then $\mathrm{Aut}_{\mathbb{Q}}(S)$ is trivial.
Proof.
We have
$\mathrm{Aut}_{\mathbb{Q}}(S)=\mathrm{Aut}_{B,\mathbb{Q}}(S)=\{\mathrm{id}_{S}\}$
where the first equality is by Proposition 6.7 and the second is by Corollary 4.5 resp. Lemma 4.6 (2). ∎
Corollary 6.10.
Let $f\colon S\rightarrow B$ be a minimal properly elliptic surface with $\chi(S)>0$. If $p_{g}(S)>0$ and if $f$ is isotrivial, then $\mathrm{Aut}_{\mathbb{Q}}(S)$ is trivial.
Proof.
Since $f$ is isotrivial and $\chi(S)>0$, there exists a fibre of additive type of $f$. Now apply Corollary 6.9. ∎
As a consequence of Corollaries 6.5 and 6.10, we obtain the following result which reproves most of Theorem 3.6:
Corollary 6.11.
Let $f\colon S\rightarrow B$ be a minimal properly elliptic surface with $\chi(S)>0$. If $f$ is isotrivial, then $|\mathrm{Aut}_{\mathbb{Z}}(S)|\leq 4$.
Now we turn to the kernel $\mathrm{Aut}_{B,\mathbb{Q}}(S)$ of the natural homomorphism $\mathrm{Aut}_{\mathbb{Q}}(S)\rightarrow\mathrm{Aut}(B)$. As the examples in Section 9 illustrate, there is no general uniform bound for $\mathrm{Aut}_{B,\mathbb{Q}}(S)$. Nevertheless, if $S$ is not isotrivial, we can bound $\mathrm{Aut}_{B,\mathbb{Q}}(S)$ by the genus $g(B)$ of the base curve $B$ (or in terms of other invariants of $S$); in contrast, the isotrivial case showcases a uniform bound independent of $g(B)$, see Theorem 3.16 , Proposition 6.1 and Corollary 6.10.
Proposition 6.12.
Let $f\colon S\rightarrow B$ be a minimal properly elliptic surface with $\chi(S)>0$. If $f$ is not isotrivial, then
(6.3) $|\mathrm{Aut}_{B,\mathbb{Q}}(S)|<12\pi^{2}(q(S)+2).$
Proof.
If $J(S)$ is rational, then the order of $\mathrm{Aut}_{B,\mathbb{Q}}(S)$ is at most $9$ by Proposition 6.1, and hence (6.3) holds in this case. If $J(S)$ is not rational, then $p_{g}(S)>0$, and we have $\mathrm{Aut}_{B,\mathbb{Q}}(S)\subset t(\mathrm{MW}(J(S))_{\mathrm{tor}})$ by Lemma 4.4. It follows by Proposition 5.5 that
$|\mathrm{Aut}_{B,\mathbb{Q}}(S)|\leq|\mathrm{MW}(J(S))_{\mathrm{tor}}|\leq 12\pi^{2}(g(B)+2)=12\pi^{2}(q(S)+2)$
where, for the last equality, we used the fact that $g(B)=q(S)$, due to the assumption that $\chi(S)>0$. ∎
Corollary 6.13.
Let $f\colon S\rightarrow B$ be a minimal properly elliptic surface with $\chi(S)>0$ such that $J(S)$ is not rational. Then
(6.4) $|\mathrm{Aut}_{\mathbb{Q}}(S)|<12\pi^{2}(q(S)+2).$
Proof.
By Propositions 6.7 and 6.12, we have
$|\mathrm{Aut}_{\mathbb{Q}}(S)|=|\mathrm{Aut}_{B,\mathbb{Q}}(S)|<12\pi^{2}(q(S)+2)$
using the exact sequence (1.2) restricted to $\mathrm{Aut}_{\mathbb{Q}}(S)$. ∎
Remark 6.14.
The bound (6.4) from Corollary 6.13 also holds if $J(S)$ is rational, but different from the four exceptional surfaces from Proposition 6.1 (3).
7. The main construction
In this section, we explain a method to construct elliptic surfaces
$f\colon S\rightarrow B$
with nontrivial $\mathrm{Aut}_{\mathbb{Q}}(S)$ by applying a suitable base change construction to a jacobian elliptic surface
$h\colon X\rightarrow B$
endowed with a (finite) subgroup $G\subset\mathrm{MW}(X)_{\mathrm{tor}}$. As a preparation, we recall how the group structure of the fibration $h$ extends over the smooth locus $F^{\#}$ of any the singular fibre $F$. In detail, if $F$ is multiplicative (or semi-stable), i.e. of Kodaira type $\mathrm{I}_{n}\;(n>0)$, then there is a short exact sequence
(7.1) $1\rightarrow\mathbb{C}^{*}\rightarrow F^{\#}\rightarrow\mathbb{Z}/n\mathbb{Z}\rightarrow 1$
which expresses the non-canonical isomorphism (2.9) in a canonical way. Here $\mathbb{C}^{*}\cong\mathbb{G}_{m}$ is identified with the identity component of $F^{\#}$ (intersecting the zero section $O$), and $1\in\mathbb{C}^{*}$ is identified with the intersection point $F^{\#}\cap O$, while $\mathbb{Z}/n\mathbb{Z}\cong A_{F}$ denotes the cyclic group of components of $F^{\#}$, with 00 corresponding to the identity component $\Theta_{0}$.
Definition 7.1.
Let $h\colon X\rightarrow B$ be a jacobian elliptic fibration, and $G\subset\mathrm{MW}(X)_{\mathrm{tor}}$ a subgroup. Suppose $h^{*}b$ is a fibre of type $\mathrm{I}_{n}$ for some $n>0$. We say $G$ splits at $F:=h^{*}b$ if, under the injection $\iota\colon G\rightarrow F^{\#}$ from (2.11), we have
$\iota(G)\;\cong\;G_{1}\times G_{2},$
where $G_{1}=\mathbb{C}^{*}\cap\iota(G)$, and $G_{2}$ is the image of $\iota(G)$ in $\mathbb{Z}/n\mathbb{Z}$ under (7.1).
Note that $G$ splits at $h^{*}b$ if and only if there are $P,\,P^{\prime}\in G$ satisfying the following conditions:
  1. (1)
    $G=\langle P,\,P^{\prime}\rangle$;
  2. (2)
    $t_{X}(P)$ acts freely on $h^{*}b$, or equivalently, $t_{X}(P)$ induces a permutation of order exactly $|P|$ on the components of $h^{*}b$;
  3. (3)
    $P^{\prime}$ and $O$ intersect $h^{*}b$ at the same component.
For immediate use, we give sufficient conditions that a subgroup $G\subset\mathrm{MW}(X)$ splits at every singular fibre of type $\mathrm{I}_{n}$. Note that these are independent of the precise shape of $X$.
Lemma 7.2.
Let $h\colon X\rightarrow B$ be a jacobian elliptic fibration with zero section $O$, and $G\subset\mathrm{MW}(X)_{\mathrm{tor}}$ a subgroup. If $G\cong\mathbb{Z}/N\mathbb{Z}\times\mathbb{Z}/N^{\prime}\mathbb{Z}$ is such that at every prime $p$, the $p$-adic valuations satisfy $|v_{p}(N)-v_{p}(N^{\prime})|\leq 1$, then $G$ splits at every singular fibre of type $\mathrm{I}_{n}$.
Proof.
Let $F$ be a fibre of type $\mathrm{I}_{n}\;(n>0)$. By the Chinese Remainder Theorem, we can argue at each prime divisor of $NN^{\prime}$ separately. If $v_{p}(N)=v_{p}(N^{\prime})=v_{p}$, then the $p$-Sylow group $S_{p}(G)$ satisfies
(7.2) $\iota(S_{p}(G))\cong\mu_{p^{v_{p}}}\times\mathbb{Z}/{p^{v_{p}}}\mathbb{Z}$
for size reasons. Meanwhile, if $\{v_{p}(N),v_{p}(N^{\prime})\}=\{v_{p},v_{p}+1\}$, then (7.2) becomes a proper inclusion $\iota(S_{p}(G))\supsetneq\mu_{p^{v_{p}}}\times\mathbb{Z}/{p^{v_{p}}}\mathbb{Z}$. Moreover, using the notation of Definition 7.1, either $\iota(S_{p}(G))\cap\mathbb{C}^{*}$ or $\mathrm{im}(\iota(S_{p}(G))\rightarrow\mathbb{Z}/n\mathbb{Z})$ has order $p^{v_{p}}$, and the other group has order $p^{v_{p}+1}$; in other words, we have
(7.3) $\iota(S_{p}(G))\cong\mu_{p^{v_{p}}}\times\mathbb{Z}/p^{v_{p}+1}\mathbb{Z},\quad\text{ or }\quad\iota(S_{p}(G))\cong\mu_{p^{v_{p}+1}}\times\mathbb{Z}/p^{v_{p}}\mathbb{Z},$
where the first alternative may only persist if $v_{p}(n)\geq v_{p}+1$. This completes the proof. ∎
Example 7.3.
The special case $G\cong(\mathbb{Z}/N\mathbb{Z})^{2}$ always splits.
Remark 7.4.
To illustrate that Lemma 7.2 is sharp, we highlight the extremal rational elliptic surface $X_{8211}$ from [MP86] (also described in [Bea82]). This has $\mathrm{MW}\cong\mathbb{Z}/4\mathbb{Z}$ not splitting at the fibre of type $\mathrm{I}_{2}$ as a generator $P$ meets $\mathrm{I}_{2}$ at the non-identity component.
Note, however, that the subgroup $\mathbb{Z}/2\mathbb{Z}\cong\langle 2P\rangle\subset\mathrm{MW}$ is splitting; applied to a deformation of $X_{8211}$ preserving the 2-torsion section, this will be used in Section 10.
We can now introduce our key construction. In practice, this results in an (algebraic!) logarithmic transformation which can be viewed as a generalization of Kondo’s construction of Enriques surfaces starting from a rational elliptic surface with a $2$-torsion section ([Kon86], later generalized in [HS11]).
Put differently, the construction realizes a homogenous space above $X$ explicitly while retaining control about the action of $\mathrm{Aut}(X)$.
Construction 7.5.
  1. (1)
[…]
  2. (2)
    By the Riemann Existence Theorem, we can define a $G$-cover
    $\widetilde{B}\stackrel{{\scriptstyle\bar{\pi}}}{{\longrightarrow}}B$
    with branch locus =${\mathcal{B}}=\{b_{1},\ldots,b_{s}\}$ containing all those points $b_{1},\ldots,b_{s_{0}}\in B$ underneath semi-stable fibres such that $P_{b_{i}}\neq O$, by choosing $t(P_{b_{i}})$ as local monodromy around $b_{i}\in{\mathcal{B}}_{0}=\{b_{1},\ldots,b_{s_{0}}\}$, complemented by suitable monodromies at the smooth fibres over $b_{s_{0}+1},\ldots,b_{s}$.
  3. (3)
    Consider the corresponding base change
[…]
    where $\widetilde{X}\rightarrow X\times_{B}\widetilde{B}$ is the minimal resolution and $\tilde{h}\colon\widetilde{X}\rightarrow\widetilde{B}$ is the induced elliptic fibration. Note that $X\times_{B}\widetilde{B}$ is normal, with an $A_{d-1}$-singularity over each node of a fibre of type $\mathrm{I}_{n}$ in the notation of (7.3), so that $\widetilde{X}$ acquires a fibre of type $\mathrm{I}_{nd}$ over each pre-image of the respective branch point. In addition, $\tilde{h}\colon\widetilde{X}\rightarrow\widetilde{B}$ is still relatively minimal.
  4. (4)
    For $P\in G$, we denote by $\sigma_{P}\in\mathrm{Aut}(\widetilde{B})$ the corresponding deck transformation of $\bar{\pi}$. At the same time, $P$ induces an automorphism $t_{X}(P)\in\mathrm{Aut}_{B}(X)$ as in (4.1). Hence the product group $G\times G$ acts on $X\times_{B}\widetilde{B}$ in a natural way: for $(P,\,P^{\prime})\in G\times G$ and $(x,\,\widetilde{B})\in X\times_{B}\widetilde{B}$,
    $(P,\,P^{\prime})(x,\,\widetilde{B}):=(t_{X}(P)(x),\,\sigma_{P^{\prime}}(\widetilde{B})).$
    By the uniqueness of the minimal resolution, the action of $G\times G$ lifts to $\widetilde{X}$, and $\tilde{h}$ is equivariant with respect to the $(G\times G)$-action on $\widetilde{X}$ and the $G$-action on $\widetilde{B}$.
The construction culminates in the following result which will form the key to constructing properly elliptic surfaces with non-trivial $\mathrm{Aut}_{\mathbb{Q}}$ or $\mathrm{Aut}_{\mathbb{Z}}$.
Theorem 7.6.
(1) The diagonal $\Delta_{G}\subset G\times G$ acts freely on $\widetilde{X}$.
(2) The quotient $S=\widetilde{X}/\Delta_{G}$ is a properly elliptic surface $S\stackrel{{\scriptstyle f}}{{\longrightarrow}}B$ such that $J(S)=X$ (or an Enriques surface if $p_{g}(X)=0,G=\mathbb{Z}/2\mathbb{Z},s_{0}=s=2$).
(3) Moreover, $S$ is endowed with a $G$-action which preserves each fibre of $f$ and acts by translations on the smooth fibres.
Proof.
(1) To confirm the free action of $\Delta_{G}$, it suffices to analyse the stabilizers $G_{b_{i}}$ at the semi-stable fibres which are the preimages of the $b\in{\mathcal{B}}_{0}$, since on the smooth ramified fibres, $\Delta_{G}$ acts by translations, hence without fixed points.
Letting $g=(P,P)\in\Delta_{G}$ for $O\neq P\in G_{b}$, we infer from the discussion preceding the theorem that $g$ rotates the fibre components of an $\mathrm{I}_{n}$ fibre of $X\times_{B}\widetilde{B}$ by $2\pi/d$ in the notation of (7.3) (because $t_{P}$ does so on $X$). It follows that same holds for the $\mathrm{I}_{nd}$ fibre of $\widetilde{X}$. Hence $\Delta_{G}$ acts freely everywhere.
(2) Since $\Delta_{G}$ respects the elliptic fibration on $\widetilde{X}$, $S$ is elliptic. More precisely, it inherits fibres of multiplicity $d_{i}$ at each branch point $b_{i}\in{\mathcal{B}}_{0}$, and similar at ${\mathcal{B}}\setminus{\mathcal{B}}_{0}$, depending on the chosen monodromy. It then follows from the canonical bundle formula (2.4) that $S$ is properly elliptic – or Enriques in the given special case. Due to the multiple fibres, the fibration $f$ admits no sections, but $P$ maps to a multisection of index $|G|$. This already shows that $J(S)=X$.
(3) The $G$-action then follows generally (regardless of the given construction) from the Jacobian since $G\subset\mathrm{MW}(J(S))$. Explicitly we can exhibit it using the group isomorphism
$G\cong(G\times G)/\Delta_{G},\quad P\mapsto(P,\,O)\mod\Delta G$
which endows $S$ with a $G$-action via the $(G\times G)$-action on $\widetilde{X}$, which commutes with the $\Delta_{G}$-action, thus descends to $S$, preserves each fibre of $\widetilde{X}$ (and thus of $f$) and acts by translations on the smooth fibres. ∎
Remark 7.7.
Elliptic surfaces with given Mordell–Weil groups can be obtained from classifications of rational elliptic surfaces (see [OS90], [SS19, Table 8.2]) or elliptic K3 surfaces (see [Shi00]). Of course, the Mordell–Weil group can also be read off from explicit equations, or inferred from universal constructions (cf. [Kub76]) as we shall exploit in Section 8.2.
In order to detect numerical trivial automorphisms of $S$, we first determine the generators of $\mathrm{Num}(S)_{\mathbb{Q}}$.
Lemma 7.8.
In the notation of Construction 7.5, let $\varpi:\widetilde{X}\to S$ denote the quotient map. Suppose that $Q_{1},\dots,Q_{m}\in\mathrm{MW}(X)$ generate the $\mathbb{Q}$-vector space $\mathrm{MW}(X)_{\mathbb{Q}}$. For $P\in\mathrm{MW}(X)$, let $\widetilde{P}:=\pi^{-1}(P)\subset\widetilde{X}$ be the inverse image of $P$ and $P_{S}:=\varpi(\widetilde{P})\subset S$ the image of $\widetilde{P}$. Then the $\mathbb{Q}$-vector space $\mathrm{Num}(X)_{\mathbb{Q}}$ is generated by the classes of $O_{S}$, $(Q_{j})_{S}$, $1\leq j\leq m$, together with the classes of the fibre components of $f\colon S\rightarrow B$. In particular, if $\mathrm{MW}(X)$ is finite, then $\mathrm{Num}(S)_{\mathbb{Q}}$ is generated by the class of $O_{S}$ together with classes of the fibre components.
Proof.
[…]
Proposition 7.9.
Suppose that the group $G$ is nontrivial in Construction 7.5. Then $G$ acts trivially on the cohomology ring $H^{*}(S,\mathbb{Q})$ if and only if there are no additive reducible fibres. In particular, this holds true if $|G|>4$.
Remark 7.10.
Construction 7.5 is quite a bit more flexible in the sense that the base change only requires stabilizer subgroups $H\subset G_{b}$ at the semi-stable fibres to yield a free $G$-action (cf. the pointer in Remark 7.4). However, the action will be numerically trivial if and only if $H=G_{b}$ at each $b\in{\mathcal{B}}_{0}$, for otherwise $t_{S}(P)$ does not preserve each component of the respective fibre for some $P\in G$ (as will be visible in the following proof).
Proof of Proposition 7.9.
Recall that the presence of an additive fibre implies that $|\mathrm{MW}_{\mathrm{tor}}|\leq 4$ by (2.12), hence the condition in the second statement. Indeed, for a section $P\in\mathrm{MW}_{\mathrm{tor}}$, $t_{X}(P)$ permutes the simple components of any additive fibre non-trivially by Fact 2.3. Since there is no ramification at the additive fibres, the same holds on $\widetilde{X}$ and on $S$.
Conversely, the construction, and the condition on the additive fibres, ensures that any $P\in G$ intersects the fibre $h^{*}b$ at the same component as $O$ for each $b\in B\setminus{\mathcal{B}}_{0}$. Hence $t_{S}(P)$ preserves each component of the fibre $f^{*}b$ for $b\in B\setminus{\mathcal{B}}_{0}$. For $b\in{\mathcal{B}}_{0}$, the same holds by construction. By Lemma 7.8, in order to show that $G$ acts trivially on $\mathrm{Num}(S)$, it suffices to check that, for any $P\in G$ and $Q\in\mathrm{MW}(X)$, one has
$t_{S}(P)(Q_{S})=Q_{S}$
where $Q_{S}:=\varpi(\pi^{-1}(Q))$. But this also follows from the construction: Using the notation of Lemma 7.8, we have
$Q_{S}=\varpi(\widetilde{Q})=\varpi\left(\bigcup_{R\in G}(\widetilde{Q}+\widetilde{R})\right)$
where $\widetilde{Q}+\widetilde{R}$ denotes the addition in $\mathrm{MW}(\widetilde{X})$, and hence
$t_{S}(P)(Q_{S})=t_{S}(P)\circ\varpi(\widetilde{Q})=t_{S}(P)\circ\varpi\left(\bigcup_{R\in G}(\widetilde{Q}+\widetilde{R})\right)\\ =\varpi\left(\bigcup_{R\in G}(\widetilde{Q}+\widetilde{R}+\widetilde{P})\right)=\varpi\left(\bigcup_{R\in G}(\widetilde{Q}+\widetilde{R})\right)=Q_{S}.$
Since $t_{S}(P)|_{F}$ is a translation for a smooth fibre $F$ of $f$, one sees that $t_{S}(P)$ acts trivially on the transcendental part of $H^{2}(S,\mathbb{Q})$. In conclusion, $G$ acts trivially on the whole $H^{2}(S,\mathbb{Q})$.
Since $\chi(S)>0$, the pullback $f^{*}\colon H^{1}(B,\mathbb{Q})\rightarrow H^{1}(S,\mathbb{Q})$ is an isomorphism. Now that $G$ induces the trivial action on $B$, it acts trivially on $H^{1}(S,\mathbb{Q})$. The proof of the numerical triviality of $G$ is thus complete. ∎
Remark 7.11.
Alternatively, in the last steps of the proof of Proposition 7.9, one could also argue directly with the quotient surface $S/G$ to deduce that $q(S/G)=q(S)$ and $p_{g}(S/G)=p_{g}(S)$, much like for an isogeny of a jacobian elliptic surface.
Remark 7.12.
We emphasize that Proposition 7.9 fits well with Theorem 1.1 (iii) as, by construction, the fibration $S\to B$ has multiple fibres of type $\phantom{}{}_{m}\mathrm{I}_{n}$ for suitable $m,n>1$.
8. Numerically trivial automorphisms – Proof of Theorems 1.1, 1.2
We are now in the position to prove the main results concerning the numerically trivial automorphism group $\mathrm{Aut}_{\mathbb{Q}}(S)$ when $p_{g}(S)>0$.
8.1. Proof of Theorem 1.1 (i)
If $p_{g}(S)>0$, then the proofs of Proposition 6.12 and Corollary 6.13 show that
(8.1) $\mathrm{Aut}_{\mathbb{Q}}(S)=\mathrm{Aut}_{B,\mathbb{Q}}(S)\hookrightarrow t(\mathrm{MW}(J(S))_{\mathrm{tor}}.$
Since the rightmost group is abelian with at most 2 generators, claim (i) follows. ∎
8.2. Proof of Theorem 1.1 (ii)
Let $M$ denote the maximal order of an element in $G$ or $M=4$ if $G\subset(\mathbb{Z}/2\mathbb{Z})^{2}$. Thanks to Theorem 7.6, proving the first statement basically amounts to applying Construction 7.5 to the universal elliptic curve $X=X(M)$ and $G=(\mathbb{Z}/M\mathbb{Z})^{2}$ (similar to what we did in the proof of Proposition 5.5; note that $G$ is splitting by Lemma 7.2, as recorded in Example 7.3). By Proposition 7.9, we get an abundance of properly elliptic surfaces $S$ with $\mathrm{Aut}_{\mathbb{Q}}(S)\supseteq(\mathbb{Z}/M\mathbb{Z})^{2}\supseteq G$.
For the second statement, we need to prove that actually equality persists in the above construction: $\mathrm{Aut}_{\mathbb{Q}}(S)=(\mathbb{Z}/M\mathbb{Z})^{2}$. More generally, by Lemma 7.2, we can also allow $G$ exactly as in Theorem 1.1 (ii). Then there is a universal elliptic curve $X(G)$ without additive fibres as soon as $|G|>4$; some explicit equations can be found in [Kub76]. By definition (or by inspection of the invariants from 8.6 which, of course, go up with the group growing), these surfaces have $\mathrm{MW}(X(G))=G$, so (8.1) implies that $\mathrm{Aut}_{\mathbb{Q}}(S)=\mathrm{MW}(X(G))=G$ as soon as $p_{g}(X(G))>0$; for instance, this holds whenever $|G|>9$. (Note that, $X(G)$ is extremal over $\mathbb{C}$, hence $\mathrm{MW}(X(G))$ is finite, by [Shi72, Thm 5.1].)
Whenever $|G|>4$ and $p_{g}(X(G))=0$, i.e. $X(G)$ exists and is rational, we apply a quadratic base change unramified at the singular fibres to obtain a semi-stable elliptic K3 surface $X^{\prime}$ with same $\mathrm{MW}_{\mathrm{tor}}$. Applying Construction 7.5 to $X^{\prime}$, we can use (8.1) to deduce the claim.
Alternatively, also for $|G|\leq 4$, we can pick a semi-stable elliptic K3 surface with the given Mordell–Weil group from [Shi00] and proceed as above. ∎
8.3. Proof of Theorem 1.1 (iii)
For a jacobian elliptic surface, or if there is an additive fibre, this follows from Corollary 6.9. Hence we may assume that all fibres are semi-stable or multiple of a smooth elliptic curve. Using Bertini’s theorem, we find a smooth irreducible multisection $D\subset S$ such that (quite generally)
  • $D$ does not pass through the singular points of $F^{\prime}$ for each fibre $F$;
  • $D$ is transversal to all the components of the singular fibres;
  • $D$ is either transversal to a smooth fibre $F$, or simply tangent at one point at most.
Using the assumption that no multiple fibre has singular support, the above implies that the map $D\to B$ is unramified at the singular fibres with singular support, and only ramified over points whose corresponding fibre is a smooth multiple fibre; it follows that the normalization $X$ of $S\times_{B}D$ is a smooth elliptic surface which, by construction, admits a section. Since the Jacobian commutes with base change (cf. [Mil86, §8], e.g.), we infer that $X$ is at the same time the base change of $J(S)$, i.e. $X=\mathrm{Jac}(S)\times_{B}D$.
Let $\varphi\in\mathrm{Aut}_{\mathbb{Q}}(S)$ be a non-trivial automorphism. By (8.1), there is $P\in\mathrm{MW}(J(S))_{\mathrm{tor}}$ such that $\varphi=t(P)$. We take the natural lift of $\varphi$ to $X$ given by the fibre product of $\varphi$ with the identity of $D$ and denote it by $\psi$, it is immediate that $\psi$ is the same as translation by $\tilde{P}$, the section of $X\to D$ induced by $P$. But $P$ and $\tilde{P}$ meet some reducible fibre (of type $\mathrm{I}_{n},n>1$) in a non-identity component by (2.10), so translation permutes fibre components and thus cannot be numerically trivial on $J(S)$ or $X$. Since the fibre is unramified, the same applies to $S$, giving the desired contradiction. ∎
8.4. Proof of Theorem 1.1 (iv)
This follows from Corollary 6.10.
8.5. Proof of Theorem 1.2 (i)
The inclusion (8.1) lets us directly infer (i) from Proposition 5.5 since the base curves of the elliptic fibrations on $S$ and on $J(S)$ are the same. ∎
8.6. Proof of Theorem 1.2 (ii) and (iii)
The bounds follow along the same lines as in the proof of Proposition 5.5. Namely, with $G\subset\mathrm{MW}(J(S))$ of size $|G|>4$, $J(S)$ is a base change of $X(G)$, and since $S$ and $J(S)$ share the same invariants as discussed in 2.5, we infer that
$q(X(G))\leq q(S),\;\;p_{g}(X(G))\leq p_{g}(S),\;\;\chi({\mathcal{O}}_{X(G)})\mid\chi(S)\;\;\text{ and }\;\;e(X(G))\mid e(S).$
Here the last two divisibilities form an easy topological consequence of the Euler number formula (2.6), since all singular fibres are semi-stable (cf. Proposition 7.9). The invariants for the given groups $G$ can be read off from [Miy06, §4.2] and [Shi72], for instance.
For $p=2,3,5$, we start again with a semi-stable rational elliptic surface $X_{0}$ with $\mathrm{MW}_{\mathrm{tor}}\cong\mathbb{Z}/p\mathbb{Z}$, namely #$\#44,\#63$ and #$\#67$ from [SS19, Table 8.2]. Applying a cyclic base change of degree $d>1$ ramified at the reducible fibre(s) and, in case there is only one reducible fibre, at a smooth fibre, we obtain an auxiliary elliptic surface $X$ with $\mathrm{MW}(X)\supset\mathbb{Z}/p\mathbb{Z}$ and with $\chi({\mathcal{O}})=p_{g}(S)+1=d$. Applying Construction 7.5 to $X$ gives the claim. ∎
9. Numerically trivial automorphisms – Proof of Theorem 1.3
Most of Theorem 1.3 has already been shown. Indeed, in the non-isotrivial case, the inequality of Theorem 1.3 (i) follows from Proposition 6.1 (1), and same for the conclusion that, in the case of equality, $J(S)=X_{3333}$. Conversely, by Lemma 7.2, we may apply Construction 7.5 to $X=X_{3333}$ and $G=\mathrm{MW}(X)\cong(\mathbb{Z}/3\mathbb{Z})^{2}$, to obtain a properly elliptic surface $S$ with $\mathrm{Aut}_{\mathbb{Q}}(S)=\mathrm{MW}(X)=(\mathbb{Z}/3\mathbb{Z})^{2}$, and hence the equality can actually be attained.
In the isotrivial case, the bound from Theorem 1.3 (ii) follows from Theorem 3.16 with the input from Proposition 6.1 (3). As a summary, one combines the statements of Proposition 6.1 (3) with (1.2) to conclude that
(9.1) $|\mathrm{Aut}_{\mathbb{Q}}(S)|=|\mathrm{Aut}_{B,\mathbb{Q}}(S)|\cdot|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|\leq 4\cdot|\mathrm{Aut}_{\mathbb{Q}}(S)|_{B}|.$
By Proposition 6.1 (3b) equality may only be attained if $J(S)=X_{33}$ as stated. The claimed examples were constructed in the proof of Theorem 3.16.
The remaining aim of this section is to give an alternative proof of Proposition 6.1 (3) (c) for $X_{33},X_{44}$ and $X_{11}(j)$; based on Construction 7.5 this will also provide surfaces attaining the maximum size of $\mathrm{Aut}_{\mathbb{Q}}(S)$ allowed by Theorem 3.16, see Lemma 9.4. For this purpose, we need a criterion for numerical triviality of an automorphism $\psi\in\mathrm{Aut}_{f}(S)$ that does not necessarily come from the group $G$ in Construction 7.5.
Lemma 9.1.
Let $S$ be a properly elliptic surface or an Enriques surface arising arising from Construction 7.5 for a jacobian elliptic surface $X$ with $\mathrm{MW}(X)$ finite.
Then an automorphism $\psi\in\mathrm{Aut}_{f}(S)$ is numerically trivial if and only if it satisfies the following two properties:
  1. (1)
    $\psi$ preserves every component of each reducible fibre;
  2. (2)
    in case $p_{g}(X)>0$, $\psi$ is contained in $\mathrm{Aut}_{B}(S)$, and the restriction $\psi|_{F}$ is a translation for a smooth fibre $F$ of $f\colon S\rightarrow B$.
Remark 9.2.
Recall that $\mathrm{Aut}(S)=\mathrm{Aut}_{f}(S)$ if $S$ is properly elliptic.
Indeed, in the set-up of Lemma 9.1, this only fails to hold in the special case where $S$ is an Enriques surface (cf. Remark 9.7).
Proof.
Suppose that $\psi\in\mathrm{Aut}_{f}(S)\setminus\{\mathrm{id}_{S}\}$ is numerically trivial. Then (1) holds by [CL21, Principle 1]. In case $p_{g}(S)>0$, the surface $S$ is properly elliptic, and we may apply Proposition 6.7 to have $\psi\in\mathrm{Aut}_{B}(S)$. By Lemma 4.4, $\psi|_{F}$ is a translation.
Next we show the ”if” part. Since $\mathrm{MW}(X)$ is finite, $\mathrm{Num}(S)_{\mathbb{Q}}$ is generated by the class of $O_{S}=\varpi(\pi^{-1}(O_{X}))$ together with the classes of the fibre components of $f$ by Lemma 7.8. Let $V\subset\mathrm{Num}(S)_{\mathbb{Q}}$ be the subspace generated by fibre components of $f$. Then, by (1) and the assumption that $\psi\in\mathrm{Aut}_{f}(S)$, we have
$\psi^{*}|_{V}=\mathrm{id}_{V}.$
By comparing the intersection numbers of $O_{S}$ and $\psi(O_{S})$ with a fibre, we infer that
$[\psi(O_{S})]=[O_{S}]+[\Theta],$
for some $[\Theta]\in V$. But then for any positive integer $n$, we have
$[\psi^{n}(O_{S})]=[O_{S}]+n[\Theta].$
Since $\psi$ has finite order, it must hold that $[\Theta]=0$. Hence $[\psi(O_{S})]=[O_{S}]$, and $\psi$ acts trivially on $\mathrm{Num}(S)_{\mathbb{Q}}$. In case $p_{g}(X)=0$, we have $p_{g}(S)=q(S)=0$, and hence $H^{1}(S,\mathbb{Q})=0$ and $\mathrm{Num}(X)_{\mathbb{Q}}\cong H^{2}(X,\mathbb{Q})$. It follows that $\psi$ is numerically trivial.
In case $p_{g}(X)>0$, since $\psi\in\mathrm{Aut}_{B}(S)$ and $\psi|_{F}$ is a translation for a smooth fibre $F$, $\psi$ acts trivially on the transcendental part of $H^{2}(X,\mathbb{Q})$ by Lemma 4.4. Of course, $\psi\in\mathrm{Aut}_{B}(S)$ also acts trivially on the base, hence of $H^{1}(S,\mathbb{Q})$. It follows that $\psi$ numerically trivial also in this case. ∎
We prove now rather explicitly Proposition 6.1 (3) (c). The surfaces are exhibited in the following example; for $r=4$, the construction can also be translated into an explicit equation for the surfaces confirming the existence part of Theorem 1.3 (ii), see Example 9.8.
Example 9.3.
Take $r\in\{2,3,4\}$. Depending on the value of $r$, we choose $h\colon X\rightarrow B=\mathbb{P}^{1}$ to be an isotrivial extremal rational surface, together with a cyclic subgroup $G\subset\mathrm{MW}(X)$ with a specified generator $P$, as follows (see [MP86] or [Mir89] for the notation):
  • If $r=2$, take $X$ to be one of the surfaces $X_{11}(\lambda)$, together with a 2-torsion section $P=(0,0)$. In this case, the Weierstrass equation of $X$ can be taken as
    $X_{11}(\lambda):\;\;y^{2}=x^{3}+tx^{2}+\lambda t^{2}x\;\;\;(\lambda\in\mathbb{C}^{*}),$
    so that the two fibres of type $\mathrm{I}_{0}^{*}$ lie over 00 and $\infty$ respectively.
  • If $r=3$, take $X$ to be $X_{44}$, together with a 3-torsion section $P=(0,0)$. In this case, the Weierstrass equation of $X$ can be taken as
    $X_{44}:\;\;y^{2}+ty=x^{3},$
    so that the two reducible fibres, of types $\mathrm{IV}$ and $\mathrm{IV}^{*}$ respectively, lie over 00 and $\infty$ respectively.
  • If $r=4$, take $X$ to be $X_{33}$, together with a 2-torsion section $P=(0,0)$. In this case, the Weierstrass equation of $X$ can be taken as
    $X_{33}:\;\;y^{2}=x^{3}+tx,$
    so that the two reducible fibres, of types $\mathrm{III}$ and $\mathrm{III}^{*}$ respectively, lie over 00 and $\infty$ respectively.
By [DM22], each of these surfaces has $\mathrm{Aut}_{\mathbb{Q}}(X)=\mathbb{C}^{*}$, with $\mu\in\mathbb{C}^{*}$ acting as
$\mu:\;(x,\,y,\,t)\mapsto(\mu^{2}x,\,\mu^{3}y,\,\mu^{r}t).$
Denote by $d$ the order of $P$, and take a degree $d$ cyclic base change $\bar{\pi}\colon\widetilde{B}\rightarrow B=\mathbb{P}^{1}$, given by the equation
(9.2) $\tilde{t}^{d}=t^{dm}-1$
where $m$ is a positive integer. Thus $\bar{\pi}$ is totally branched at the $dm$ points $b_{k}=\xi_{dm}^{k}$ with $10\leq k\leq md-$ 1 where $\xi_{n}=\exp(2\pi\sqrt{-1}/n)$ denotes a primitive $n$-th root of unity.
Now apply Construction 7.5 to $(X,\,G,\,\bar{\pi})$ as above so that the stabilizer over each $b_{k}$ is $G_{b_{k}}=\langle P\rangle$. Let $f\colon S\rightarrow B$ be the elliptic surface thus obtained. Note that, by Proposition 7.9, translation by $P$ does not induce a numerically trivial automorphism of $S$.
Let $\psi_{X,m}\in\mathrm{Aut}_{\mathbb{Q}}(X)=\mathbb{C}^{*}$ be the automorphism of order $rdm$ given by $\xi_{rdm}\in\mathbb{C}^{*}$, i.e.
(9.3) $\psi_{X,m}:\;\;(x,\,y,\,t)\mapsto(\xi_{rdm}^{2}x,\,\xi_{rdm}^{3}y,\,\xi_{rdm}^{r}t).$
This action lifts to an automorphism $\psi_{\widetilde{X},m}$ of $\widetilde{X}:=X\times_{B}\widetilde{B}$ given by
(9.4) $(x,\,y,\,t,\,\tilde{t})\mapsto(\xi_{rdm}^{2}x,\,\xi_{rdm}^{3}y,\,\xi_{rdm}^{r}t,\,\tilde{t}).$
which is smooth.
Being a numerically trivial automorphism, $\psi_{X,m}$ preserves each section $Q\in\mathrm{MW}(X)$ as well as each component of the two reducible fibres $h^{*}0,\,h^{*}\infty$. It follows that $\psi_{X,m}$ commutes with $t_{X}(P)\in\mathrm{Aut}_{B}(X)$, and hence $\psi_{\widetilde{X},m}$ commutes with $t_{X}(P)\times\mathrm{id}_{\widetilde{B}}$, which generates the action of $G\times\{O\}$ on $\widetilde{X}$. Also, $\psi_{\widetilde{X},m}$ commutes with the deck transformations of $\widetilde{X}\rightarrow X$, which is generated by $\mathrm{id}_{X}\times\sigma_{P}\in\mathrm{Aut}(\widetilde{X})$:
$(x,\,y,\,t,\,\tilde{t})\mapsto(x,\,y,\,t,\,\xi_{d}\tilde{t}).$
In conclusion, $\psi_{\widetilde{X},m}$ commutes with the $(G\times G)$-action on $\widetilde{X}$.
Thus, $\psi_{\widetilde{X},m}$ descends to an automorphism $\psi_{S,m}\in\mathrm{Aut}_{f}(S)$, which evidently preserves the components of the two reducible fibres $f^{*}0,\,f^{*}\infty$. Since $\mathrm{MW}(X)$ is finite, we may apply Lemma 9.1 to obtain the numerical triviality of $\psi_{S,m}$.
By construction, we see that the order of $\psi_{S,m}$ is $rdm$, because the induced automorphism $\psi_{S,m}|_{B}\in\mathrm{Aut}(B)$ has order $dm$, and the power $\psi:=\psi_{S,m}^{dm}$ lies in $\mathrm{Aut}_{B,\mathbb{Q}}(S)$ and has order $r$, as one can see by restricting to a general fibre of $S$.
On the other hand, since $S$ has $p_{g}(S)=p_{g}(J(S))=p_{g}(X)=0$ and exactly $dm$ multiple fibres, it was shown in the proof of Proposition 6.1 that $P_{2}(S)=dm-1$. Hence,
$\mathrm{Aut}(S)|_{B}=\langle\psi_{S,m}|_{B}\rangle\cong\mathbb{Z}/dm\mathbb{Z}.$
Lemma 9.4.
The automorphism $\psi_{S,m}$ generates the whole group $\mathrm{Aut}_{\mathbb{Q}}(S)$. Therefore, $|\mathrm{Aut}_{\mathbb{Q}}(S)|=r(P_{2}(S)+1)$.
Proof.
Any two elements $P_{1},P_{2}\in\mathrm{MW}(X)$ pass through distinct components of either reducible fibre $h^{*}0$ and $h^{*}\infty$. As noticed above, we infer that $t_{S}(\mathrm{MW}(X))\cap\mathrm{Aut}_{\mathbb{Q}}(S)=\{\mathrm{id}_{S}\}$, where $t_{S}\colon\mathrm{MW}(X)\rightarrow\mathrm{Aut}_{B}(S)$ is as in (4.1). Thus the homomorphism $\mathrm{Aut}_{B,\mathbb{Q}}(S)\rightarrow\mathrm{Aut}_{B,\mathbb{Q}}(J(S))$ induced by $\Phi$ from Lemma 4.2 is injective. We know that $\mathrm{Aut}_{B,\mathbb{Q}}(J(S))$ has order $r$ by [DM22, Table 3], which is the same as the order of $\psi_{S,m}^{dm}\in\mathrm{Aut}_{B,\mathbb{Q}}(S)$. Therefore, $\mathrm{Aut}_{B,\mathbb{Q}}(S)$ and $\mathrm{Aut}_{B,\mathbb{Q}}(J(S))$ have the same order, and the former is generated by $\psi_{S,m}^{dm}$. In conclusion, $\mathrm{Aut}_{\mathbb{Q}}(S)$ is generated by $\psi_{S,m}$. ∎
This completes the proof of Theorem 1.3. ∎
Remark 9.5.
For $l\mid md$, the elliptic surface $S$ with the numerically trivial automorphism $\psi_{S,m}^{l}$ deforms in an $(l-1)$-dimensional family, the polynomial on the right hand side of (9.2) admits $l-1$ additional monomials in the powers of $t^{dm/l}$.
Remark 9.6.
Since $\psi_{S,m}$ permutes multiple fibres of $f$, it cannot be cohomologically trivial by Lemma 6.2. However, $\psi=\psi_{S,m}^{dm}$ preserves every fibre, so it has a chance to be cohomologically trivial.
Indeed, we verified in 3.5 that, in the case $r=4$, $\psi^{2}$ is cohomologically trivial, but $\psi$ is not.
And we checked that $\psi$ is not cohomologically trivial in the case $r=3$.
Remark 9.7 (Enriques surfaces).
In the case $r=2$ of the above construction, if we take $m=1$, then $\psi_{X,1}$ has order 4 and $\psi_{X,1}^{2}=(-1)_{X}$. The constructed surface $S$ is an Enriques surface with an order $4$ numerically trivial automorphism $\psi_{S,1}$ such that $\psi_{S,1}^{2}=(-1)_{S}$, the involution induced by $(-1)_{X}$.
In fact, we may take the initial jacobian elliptic surface $X$ to be one of $X_{222}$, $X_{321}$, $X_{411}$ (in the notation of [MP86]) together with a $2$-torsion section $P$ as well, and then make an appropriate $(\mathbb{Z}/2\mathbb{Z})$-cover $\bar{\pi}\colon\widetilde{B}\rightarrow B=\mathbb{P}^{1}$. Applying Construction 7.5, we obtain another three families of elliptic Enriques surfaces $f\colon S\rightarrow B$ with a numerically trivial involution $(-1)_{S}\in\mathrm{Aut}_{B,{\mathbb{Q}}}(S)$.
In this way, as $X$ varies among the surfaces $X_{222}$, $X_{321}$, $X_{411}$ and $X_{11}(j)$, we obtain four families of Enriques surfaces $S$ with a numerically trivial involution $(-1)_{S}$, three of which turn out to be distinct (cf. [MN84], [Muk10]).
Example 9.8 ($r=4$).
Following [KS24], one can show that the surfaces $S$ originating from $X_{33}$ as above can be defined as minimal resolutions of
(9.5) $S_{0}:\;\;y^{2}=gx^{4}+g^{3}t$
where $g\in{\mathbb{C}}[t]$ generally can be any polynomial without multiple roots. In the above specific examples, $g=t^{dm}-1$.
10. Cohomologically trivial automorphisms – Proof of the order 2-statement of Theorem 1.4
The general statements (i) and (ii) of Theorem 1.4 follow from Corollaries 6.3 and 6.5. In this section, we construct properly elliptic surfaces $S$ with $\mathrm{Aut}_{\mathbb{Z}}(S)\cong\mathbb{Z}/2\mathbb{Z}$ in abundance, confirming this part of Theorem 1.4 (iii). That is, for any $s\in\mathbb{N}$, we construct a $2s$-dimensional family of non-isotrivial elliptic surfaces with $p_{g}=q=0$ admitting a cohomologically trivial involution.
Let $h\colon X\rightarrow B=\mathbb{P}^{1}$ be a rational elliptic surface with a 2-torsion section $P$ and a single type $\mathrm{I}_{8}$ reducible fibre $h^{*}\infty=\sum_{i=0}^{7}\Theta_{k}$ at $\infty$ such that $P$ intersects $\Theta_{4}$. Such $X$ depends on one parameter up to the standard action by ×${\mathbb{C}}^{*}\times\mathrm{GL}(2,{\mathbb{C}})$; to see this, just use the Weierstrass form
(10.1) $S\colon\quad y^{2}=x(x^{2}+2a_{2}x+1)$
where $a_{2}\in{\mathbb{C}}[t]$ has degree two and the $\mathrm{I}_{8}$ fibre has been located at $\infty$.
Let $\widetilde{B}\rightarrow B$ be a double cover branched at $\infty$ and above $12s-$ 1 points outside the zeros of the discriminant locus given by $a_{2}^{2}-1$.
Applying Construction 7.5, we obtain a $2s$-dimensional family of non-isotrivial elliptic surfaces $f\colon S\rightarrow B$ with an involution $\varphi=t_{S}(P)$ induced by $P\in\mathrm{MW}(X)$. By Proposition 7.9, $\varphi$ is numerically trivial.
There are $2s$ multiple fibres of $f$, say $2\mathrm{I}_{8}=2F^{\prime}=2F^{\prime}_{1},2F^{\prime}_{2},\ldots,2F^{\prime}_{2s}$ (each of the latter having smooth support, i.e. of type $2\mathrm{I}_{0}$) and
$K_{S}=-F+F^{\prime}_{1}+\ldots+F^{\prime}_{2s}.$
where $F$ is any fibre of $f$. The invariants of $S$ can be readily computed:
$p_{g}(S)=q(S)=0,\quad h^{0}(2K_{S})=2s-1$
If $s>1$, $S$ is a properly elliptic surface; if $s=1$, $S$ is the Enriques surface first studied by Barth–Peters [BP83] (as can be inferred from Figure 1).
Proposition 10.1.
The involution $\varphi$ is cohomologically trivial.
Proof.
The construction endows us with 10 curves on $S$ with square $(-2)$:
  • the eight components $\Theta_{0},\ldots,\Theta_{7}$ of the $2\mathrm{I}_{8}$ fibre, numbered cyclically such that
  • the bisection $O_{S}$ which is the image of $O$ and $P$ meets $\Theta_{0}$, and
  • another bisection $R_{S}$ which is the image of the missing generator $R$ of $\mathrm{MW}(X)$ (and of $R+P$) and which will be shown to meet $\Theta_{4}$.
The section $R$ on $X$ has height $1/2$ by [SS19, Table 8.2], and is thus of infinite order. In the degenerate case with an additional reducible fibre, of type $\mathrm{I}_{2}$, it becomes a torsion section of order $4$ such that $2R=P$. Explicitly, $R=(u,v)$ is determined, up to sign and translation by $P$, by the $x$-coordinate $u\in{\mathbb{C}}$ such that the right-hand side of (10.1) is a perfect square in ${\mathbb{C}}[t]$ (namely $v^{2}$). In particular, the construction verifies that $R$ meets $\Theta_{2}$ on $X$ (and $\Theta_{4}$ on $\widetilde{X}$, but there it’s the fourth component out of 16).
Note also that $O,\,P,\,R$ and $P+R$ are all disjoint on $X$, and so are $O_{S},R_{S}$ on $S$, but they are not smooth rational if $s>1$ (since they map $2:1$ to the base curve $\mathbb{P}^{1}$ with $2s$ ramification points), so we better not call them $(-2)$-curves. We depict these 10 curves in the following diagram:
[…]
Figure 1. Configurations of ten curves, forming a $\mathbb{Q}$-basis of $\mathrm{Num}(S)$
Together these 10 curves generate an index 4 sublattice $L$ of $\mathrm{Num}(S)$ (by inspection the determinant $-16$ of their Gram matrix since $\mathrm{Num}(S)$ is unimodular).Alternatively, consider the auxiliary lattice $L_{0}$ generated by the bisection $O_{S}$ and the fibre components $\Theta_{i}$. Then $L_{0}\cong U\oplus A_{7}$ has determinant $8$, and we can mimic the approach from the theory of Mordell–Weil lattices by applying the orthogonal projection $\pi:L\otimes\mathbb{Q}\to(L_{0}\otimes\mathbb{Q})^{\perp}$ to get $\pi(R_{S})=R_{S}-O_{S}-2E-a_{4}^{\vee}$ where $a_{4}^{\vee}$ denotes the dual vector of the middle vertex in the Dynkin diagram $A_{7}$. Then $\pi(R_{S})^{2}=-2$, yielding $\det(L)=\det(L_{0})\cdot\pi(R_{S})^{2}=-16$. For a $\mathbb{Z}$-basis of $\mathrm{Num}(S)$, we would like to complement them by fractions of two isotropic vectors which are motivated from the Enriques surface case $s=1$:
$D_{1}$ $=$ $2O_{S}+4\Theta_{0}+3(\Theta_{1}+\Theta_{7})+2(\Theta_{2}+\Theta_{6})+\Theta_{3}+\Theta_{5}\;\;(\text{type }\mathrm{III}^{*})$
$D_{2}$ $=$ $O_{S}+\Theta_{1}+2(\Theta_{0}+\Theta_{7}+\Theta_{6}+\Theta_{5}+\Theta_{4})+\Theta_{3}+R_{S}\;\;(\text{type }\mathrm{I}_{4}^{*})$
More precisely, in the Enriques surface case, each divisor induces an elliptic fibration which, a priori, features the divisor singular fibre or as half-pencil. The latter alternative, however, is excluded since multiple fibres are only supported on types $\mathrm{I}_{n}(n\geq 0)$; therefore, each divisor is 2-divisible, and we obtain $\mathrm{Num}(S)$ as desired.
Lemma 10.2.
For any $s\geq 1$, $\mathrm{Num}(S)$ is obtained from $L$ by adjoining the classes of $\frac{1}{2}D_{1}=\Theta_{4}^{\vee},\frac{1}{2}D_{2}=\Theta_{2}^{\vee}$.
The proof of the lemma will require a bit of work. In fact, along the way, we will deduce that $\varphi$ is cohomologically trivial!
Since $\mathrm{Num}(S)$ is a unimodular overlattice of $L$, it is encoded in an isotropic subgroup of the discriminant group $A_{L}=L^{\vee}/L$ of size $4$. We start by exhibiting generators for $A_{L}$. For this purposes, it suffices to note that the isotropic divisor
$D_{3}=O_{S}+2\Theta_{0}+3\Theta_{1}+4\Theta_{2}+5\Theta_{3}+6\Theta_{4}+4\Theta_{5}+2\Theta_{6}+3R_{S}=4\Theta_{7}^{\vee}.$
Note that the configuration is that of a divisor of Kodaira type $\mathrm{II}^{*}$, so by the same argument as above, it would be 2-divisible on an Enriques surface, but certainly not 4-divisible. Apparently, we get a vector $\frac{1}{4}D_{3}$ of order 4 in $L^{\vee}/L$, and another one by applying the horizontal symmetry of Figure 1:
$D_{3}^{\prime}=O_{S}+2\Theta_{0}+3\Theta_{7}+4\Theta_{6}+5\Theta_{5}+6\Theta_{4}+4\Theta_{3}+2\Theta_{2}+3R_{S}=4\Theta_{1}^{\vee}.$
By inspection of the coefficients of $\Theta_{3}$ and $\Theta_{5}$, the vectors $v=\frac{1}{4}D_{3},v^{\prime}=\frac{1}{4}D_{3}^{\prime}$ generate $A_{L}\cong(\mathbb{Z}/4\mathbb{Z})^{2}$ with intersection form $U(3/4)$, the hyperbolic plane scaled by $3/4$. It is then an easy exercise to check that the only isotropic vectors mod $2\mathbb{Z}$ (for $\mathrm{Num}(S)$ is even by the adjunction formula) are ±$\pm v,\pm v^{\prime}$ and the 2-torsion elements
[…]
For symmetry reasons, it thus suffices to inspect the following two cases:
(10.2) $\mathrm{Num}(S)=\left\langle L,\frac{1}{2}D_{1},\frac{1}{2}D_{2}\right\rangle\;\;\;\text{ or }\;\;\;\left\langle L,\frac{1}{4}D_{3}\right\rangle.$
We now focus on the isotropic vectors adjoined to $L$, to verify that $\varphi$ is cohomologically trivial and eventually rule out the second alternative. For this purpose, we let
$D=\frac{1}{2}D_{1},\frac{1}{2}D_{2}\;\;(\text{in the first case) \; or }\;\;\frac{1}{4}D_{3}\;\;(\text{in the second case}).$
We start by modifying $D$ to become effective. To this end, let $N\in{\mathbb{N}}_{0}$ and consider
$\hat{D}=\frac{1}{2}D_{1}+NF^{\prime}.$
Since $\hat{D}.F^{\prime}=D.F^{\prime}=1$, it follows from Riemann–Roch that
$\chi(\hat{D})=1+(N-(s-1))/2.$
Hence $\hat{D}\geq 0$ for all $N\geq s-1$ (otherwise $-\hat{D}\geq 0$, but then $(-\hat{D}.F^{\prime})=-1$ contradicting that $F^{\prime}$ is nef). Fixing the minimal such $N$, this satisfies $N\leq s-1$.
Claim 10.3.
$\hat{D}=C$ for some irreducible curve $C\subset S$.
Proof of Claim 10.3.
Writing $\hat{D}=C_{1}+\ldots+C_{r}$ for irreducible curves $C_{i}\subset S$, we have to show that $r=1$. Since $\hat{D}.F^{\prime}=1$, we may ssume without loss of generality that $C_{1}$ is a bisection while the other $C_{i}$ are fibre components. Since $C_{i}=F^{\prime}$ or $F$ are excluded by the minimality assumption for $N$, we can only have $C_{2}=\Theta_{j}$ for some component of the $\mathrm{I}_{8}$ fibre. Regardless of $j$, $\hat{D}-\Theta_{j}\geq 0$ implies that $(\Theta_{j-1}.(\hat{D}-\Theta_{j}))<0$ or $(\Theta_{j+1}.(\hat{D}-\Theta_{j}))<0$, since $D$ was defined as dual vector of some $\Theta_{l}$. Therefore $\Theta_{j-1}$ or $\Theta_{j+1}$ is contained in the support of $\hat{D}$. But this shows successively that $F^{\prime}\subset$ supp$(\hat{D})$ – which we excluded before. Hence $r=1$ as claimed. ∎
To see that $\varphi(C)=C$, we note that $C^{2}=2N$ by definition. Hence $C.\varphi(C)=2N$, since $\varphi$ is numerically trivial. Assuming that $\varphi(C)=C^{\prime}\neq C$, we can compare this intersection number against the number of intersection points forced upon $C$ and $C^{\prime}$ by the geometry of $S$.
To this end, consider the precise action of the involution on the fibre components. Indeed, the $12s-$ 1 ramified smooth fibres are fixed pointwise. It follows that $C$ meets $S^{\varphi}$ in at least $12s-$ 1 points, so
$C.C^{\prime}\geq 2s-1>2N=C.C^{\prime}.$
This gives the required contradiction and thus completes the argument that $\varphi(C)=C$. In consequence, $\varphi$ fixes a set of generators of $\operatorname{Pic}(S)$, so it is cohomologically trivial. This completes the proof of Proposition 10.1. ∎
Having analysed the action of $\varphi$, we come back to the proof of Lemma 10.2. For this purpose, we use the action of $\varphi$ on the ramified $\mathrm{I}_{8}$ fibre. One checks that the action alternates between
  • fixing a component pointwise, starting from the component $\Theta_{0}$ met by the bisection $O_{S}$, and continued with $\Theta_{2},\Theta_{4},\Theta_{6}$, and
  • acting as on involution on $\Theta_{1},\Theta_{3},\Theta_{5},\Theta_{7}$, only fixing the two nodes where the adjacent fibre components are intersected.
Explicitly, this can be seen by considering $\iota$ on $\widetilde{X}$, since both $\iota$ and $t_{P}$ induce the same involution on $S$. Now, for $D=\frac{1}{2}D_{1}$, the curve $C$ meets $\Theta_{4}$, and for $D=\frac{1}{2}D_{2}$, the curve $C$ meets $\Theta_{2}$; both components are fixed pointwise by $\varphi$ which is compatible with the property that $C=\varphi(C)$. For $D=\frac{1}{4}D_{3}$, however, the curve $C$ meets $\Theta_{7}$ which is not fixed pointwise by $\varphi$. Since the bisection $C$ cannot met the nodes, its intersection point with $\Theta_{7}$ is not fixed, contradicting the conclusion that $\varphi(C)=C$. Hence $\frac{1}{4}D_{3}\not\in\mathrm{Num}(S)$. This completes the proof of Lemma 10.2. ∎
Remark 10.4.
For an alternative infinite series of families of non-isotrivial properly elliptic surfaces $S$ with $|\mathrm{Aut}_{\mathbb{Z}}(S)|=2$, one could realize the isotrivial families for $r=4$ from 3.5 by way of Construction 7.5 applied to $X_{33}$, as in Section 9, and then deform (preserving $\Psi^{2}$) $X_{33}$ to $X_{321}$ (as hinted in Remark 9.7).
11. Cohomologically trivial automorphisms – Proof of the order 3-statement of Theorem 1.4
In this section, we construct a 1-dimensional family of non-isotrivial properly elliptic surfaces with $q=p_{g}=0$ admitting a cohomologically trivial automorphism of order 3, proving the remaining part of Theorem 1.4 (iii).
Let $X$ be the extremal rational elliptic surface with singular fibres of types $\mathrm{I}_{9}$ and three times $\mathrm{I}_{1}$ (No. 63 in [OS90] or [SS19, Table 8.3]). This has $\mathrm{MW}(X)=\langle P\rangle\cong\mathbb{Z}/3\mathbb{Z}$ and can be given in Weierstrass form
$X:\;\;y^{2}+txy+y=x^{3}.$
Upon numbering the components of the $\mathrm{I}_{9}$ fibre cyclically $\Theta_{0},\ldots,\Theta_{8}$ such that the zero section $O$ meets $\Theta_{0}$, we may assume that $P$ meets $\Theta_{3}$. This follows either from the abstract description
$\mathrm{MW}(X)\cong\operatorname{Pic}(X)/\mathrm{Triv}(X)=\mathrm{Triv}(S)^{\prime}/\mathrm{Triv}(S)$
from 2.5 where the trivial lattice Triv$(X)$ is generated by fibre components and zero section (in parallel to the argument around (11.2)), or from considering explicit equations as in [Bea82], [MP86] or [SS19, Table 8.3].
Let $\bar{\pi}\colon\widetilde{B}=\mathbb{P}^{1}\rightarrow B=\mathbb{P}^{1}$ be a cyclic cubic base change branched at $b_{1}$ and $b_{2}$ such that $h^{*}b_{1}$ is the $\mathrm{I}_{9}$ fibre and $h^{*}b_{2}$ is a smooth fibre $F$.
Applying Construction 7.5 to the triple $(X,\,\langle P\rangle,\,\bar{\pi})$, we obtain a properly elliptic surface $f\colon S\rightarrow B$ with $p_{g}=q=0$, together with an automorphism $\varphi\in\mathrm{Aut}_{B}(S)$ of order $3$. By Proposition 7.9, $\varphi=t_{S}(P)$ is numerically trivial.
Proposition 11.1.
The automorphism $\varphi$ is cohomologically trivial.
Proof.
First note that the elliptic fibration $f\colon S\rightarrow B$ has two triple fibres, of types $3\mathrm{I}_{0}$ and $3\mathrm{I}_{9}$ respectively, and a smooth rational trisection $B$ which is the image of $O$ (or $P$, or $-P$). By construction, we have $B^{2}=-3$.
The fixed locus $S^{\varphi}$ consists of
  • the smooth genus one curve $E$ supporting the $3\mathrm{I}_{0}$ fibre,
  • the components $\Theta_{0},\Theta_{3},\Theta_{6}$ of $\mathrm{I}_{9}$, numbered cyclically, as before, such that $B$ meets $\Theta_{0}$,
  • the 3 intersection points $\Theta_{i}\cap\Theta_{i+1}$ for $i=1,4,7$,
  • the 3 nodes of the (simple) $\mathrm{I}_{1}$ fibres.
From [CFGLS24, Lemma 2.5], we get that
$\mathbb{Z}/3\mathbb{Z}\cong H_{1}(S,\mathbb{Z})\cong H^{2}(S,\mathbb{Z})_{tor}=\operatorname{Pic}(S)_{tor}$
is generated by $E-D^{\prime}$ where $D^{\prime}=\Theta_{0}+\ldots+\Theta_{8}$ is the sum of the fibre components of the reduced $\mathrm{I}_{9}$ fibre. Note that $K_{S}=-F+2E+2D^{\prime}$ where $F$ denote any fibre.
To see that $\varphi$ acts cohomologically trivially, since it acts trivially on the first homology, we need generators of $\operatorname{Pic}(S)$. To this end, we switch to the Severi lattice $\mathrm{Num}(S)$ and consider the standard index 3 sublattice $L$,
(11.1) $\mathrm{Num}(S)\supset\langle E,B\rangle\oplus\langle\Theta_{1},\ldots\Theta_{8}\rangle\cong\begin{pmatrix}0&1\\ 1&-3\end{pmatrix}\oplus A_{8}=L$
which parallels the trivial lattice of a jacobian elliptic surface reviewed in 2.5. Note that integral overlattices of $L$ are in bijective correspondence with isotropic subgroups of the discriminant group $L^{\vee}/L$. Presently, this group is isomorphic to $A_{8}^{\vee}/A_{8}=\langle\Theta_{1}^{\vee}\rangle\cong\mathbb{Z}/9\mathbb{Z}$, with generator the dual vector $\Theta_{1}^{\vee}$ of square $-8/9$. Since $\mathrm{Num}(S)$ is unimodular, we infer that it is obtained from $L$ by adjoining the class of the dual vector
(11.2) $D_{0}=\Theta_{3}^{\vee}=-\frac{1}{3}(2\Theta_{1}+4\Theta_{2}+6\Theta_{3}+5\Theta_{4}+4\Theta_{5}+3\Theta_{6}+2\Theta_{7}+\Theta_{8});$
indeed, $3D_{0}\in L$ and $D_{0}^{2}=-2$, confirming the claim. Note that $D_{0}$ has three lifts in $\operatorname{Pic}(S)$ – which we can treat all alike (!), so we shall just write $D_{0}$ for some lift in $\operatorname{Pic}(S)$ or for the class in $\mathrm{Num}(S)$ without differentiating. Indeed:
Claim 11.2.
Independent of the chosen lift, $D=D_{0}+B+3E$ is effective.
Proof.
Since $D^{2}=D.K_{S}=D.E=1$, Riemann–Roch gives $\chi(D)=1$, so $h^{0}(D)+h^{0}(K_{S}-D)\geq 1$. But then $E.(K_{S}-D)=-1$, so $K_{S}-D$ cannot be effective, since $E$ is nef. Hence $D>0$ as claimed. ∎
As a consequence, we can write
$D=C_{1}+\ldots+C_{r}$
for irreducible curves $C_{i}\subset S$. Since $D.E=1$, we may assume that $C_{1}.E=1$ (so $C_{1}$ is another trisection of the elliptic fibration), while the $C_{i}$ are fibre components.
Claim 11.3.
$r=1$, i.e. $D=C_{1}$.
Proof.
If $D-E\geq 0$, then we use $(D-E).B=-1$ to infer that $B\subset$ supp$(D)$, but this is only possible if $C_{1}=B$. It follows that $D$, and thus also $D_{0}$, is supported on fibre components and on $B$, i.e. on the index 3 sublattice of $\mathrm{Num}(S)$ from (11.1). But this contradicts the choice of $D_{0}$; hence $E\not\subset$ supp$(D)$.
Assume that $\Theta_{i}\subset$ supp$(D)$ for some $i\in\{0,\ldots,8\}$. Then $\Theta_{i-1}.(D-\Theta_{i})<0$ or $\Theta_{i+1}.(D-\Theta_{i})<0$, and we can infer successively that each $\Theta_{j}$ is contained in supp$(D)$. But this gives $D-D^{\prime}\geq 0$, with the same contradiction as before.
Obviously this argument also covers the case $F\subset$ supp$(D)$ for any simple fibre $F$, so the claim follows. ∎
To complete the proof of Proposition 11.1, it remains to verify that $\varphi(C_{1})=C_{1}$, for then there is a set of generators of $\operatorname{Pic}(S)$ each of which is fixed by $\varphi$.
Assume to the contrary that $\varphi(C_{1})=C^{\prime}\neq C_{1}$. Then $C_{1}.C^{\prime}=1$ by the invariance of intersection numbers. On the other hand,
$C_{1}\cap\varphi(C_{1})\;\supset\;C_{1}\cap\mbox{Fix}(\varphi).$
By construction, $C_{1}$ not only intersects $E\subset\mbox{Fix}(\varphi)$ non-trivially, but also $\Theta_{3}\subset\mbox{Fix}(\varphi)$. Hence #$\#(C_{1}\cap\varphi(C_{1}))\geq 2$, contradicting the fact that $C_{1}.C^{\prime}=1$. It follows that $\varphi(C_{1})=C_{1}$ as required. This completes the proof of Proposition 11.1. ∎
Summary 11.4.
With Propositions 10.1, 11.1 in place, the proof of Theorem 1.4 is complete. ∎
Remark 11.5.
With more triple fibres with smooth support, $\varphi$ remains to be numerically trivial, but it is not obvious to decide whether it is cohomologically trivial: since multiple fibres $3E$ appear as $2E$ in the canonical divisor, the intersection numbers go up more quickly, and the argument from the proof of Proposition 11.1 using $C\cap S^{\varphi}$ is not decisive.
12. Proof of Theorem 1.6
If $S$ has $\chi(\mathcal{O}_{S})>0$, then the statement of Theorem 1.6 that $\mathrm{Aut}_{\mathbb{Z}}(S)$ is abelian follows from the results of this paper:
  • if $p_{g}>0$, then the statement follows from Theorem 1.1 (i);
  • if $p_{g}=0$, but the fibration is not isotrivial, then the statement can be derived from Corollary 6.3 since this implies that (8.1) also holds in this case, as in the proof of Theorem 1.1 (i);
  • if the fibration is isotrivial, then $|\mathrm{Aut}_{\mathbb{Z}}(S)|\leq 4$ by Theorem 3.6 which a posteriori implies that $\mathrm{Aut}_{\mathbb{Z}}(S)$ is abelian.
If $S$ has $\chi(\mathcal{O}_{S})=0$, we write $S$ as a free quotient of a product, $S=(C\times E)/G$. Then [CFGLS24, Thm 1.2] states explicitly that $\mathrm{Aut}_{\mathbb{Z}}(S)$ is abelian except for a special case where $G\cong\mathbb{Z}/2m\mathbb{Z}$ for odd $m$ acts by translations on $E$ and $\mathrm{Aut}_{\mathbb{Z}}(S)/\mathrm{Aut}^{0}(S)\cong\mathbb{Z}/2\mathbb{Z}=\langle\imath\rangle$. But here, the action of $\mathrm{Aut}^{0}(S)=E$ is induced by translations on the second factor of $C\times E$ while $\imath$ can only be induced by an involution on the first factor, possibly composed with translation by a 2-torsion point on $E$. In either case, these automorphisms commute on $C\times E$ and thus also on $S$. ∎
Remark 12.1.
For $\mathrm{Aut}_{\mathbb{Q}}(S)$, we only exhibited abelian examples in this paper (with $\chi>0$), but there are also non-abelian examples since one can show that any finite subgroup $H$ of $\mathrm{Aut}(\mathbb{P}^{1})$ can be realized as $\mathrm{Aut}_{\mathbb{Q}}(S)$ in case $\chi=0$.
It suffices to define $C$ as the fibre product of $B:=\mathbb{P}^{1}\rightarrow B^{\prime}:=\mathbb{P}^{1}/H$ and of a general ramified double cover $C^{\prime}\rightarrow B^{\prime}=(\mathbb{P}^{1}/H)$; letting $\iota$ be the involution of $C$ induced by the identity of $B$ and the covering involution of $C^{\prime}$, $H\times\langle\iota\rangle$ acts on $C$. We let $E=\mathbb{C}/\mathbb{Z}\oplus\tau\mathbb{Z}$ be an elliptic curve, and $S:=(C\times E)/(\mathbb{Z}/2)$, where $\mathbb{Z}/2$ acts via $(x,z)\mapsto(\iota(x),z+1/2)$. Then $H\times\{\mathrm{id}_{E}\}$ induces an inclusion $H<\mathrm{Aut}_{\mathbb{Q}}(S)$.
Acknowledgements
We thank John Voight for start-up aid regarding Magma. The second author would like to thank Professor Jin-Xing Cai for many helpful discussions on the topic of numerically trivial automorphisms.